### v3

Abstract
In this paper, we address a manifold constrained nonsmooth optimization problem involving the composition of a weakly convex function and a smooth mapping. To find a stationary point of the target problem, we propose a variable smoothing-type algorithm by combining the ideas of (i) translating the constrained problem into a Euclidean optimization problem with a smooth parametrization of the constraint set; (ii) exploiting a sequence of smoothed surrogate functions, of the cost function, given with the Moreau envelope of a weakly convex function. The proposed algorithm produces a vector sequence by the gradient descent update of a smoothed surrogate function at each iteration. In a case where the proximity operator of the weakly convex function is available, the proposed algorithm does not require any iterative solver for subproblems therein. By leveraging tools in the variational analysis, we show the so-called gradient consistency property, which is a key ingredient for smoothing-type algorithms, of the smoothed surrogate function used in this paper. Based on the gradient consistency property, we also establish an asymptotic convergence analysis for the proposed algorithm together with convergence rate analysis. Numerical experiments demonstrate the efficacy of the proposed algorithm.
keywords:
manifold constrained nonsmooth optimization, weakly convex composite, variable smoothing, gradient consistency, Moreau envelope, parametrization
pacs:
MSC Classification]49J52, 49K99
1 Introduction
In this paper, we consider the following nonsmooth optimization problem with a manifold constraint.
Problem 1.1.
Let $\mathcal{X}$ and $\mathcal{Z}$ be Euclidean spaces, i.e., finite-dimensional real Hilbert spaces, and $(\emptyset\neq)C\subset\mathcal{X}$ a connected closed embedded submanifold of $\mathcal{X}$ (see Example 2.3). Then,
$\mathrm{find}\ \bm{x}^{\star}\in\mathop{\mathrm{argmin}}\limits_{\bm{x}\in C}(\underbrace{h+g\circ\mathfrak{S}}_{\eqqcolon F})(\bm{x})=\mathop{\mathrm{argmin}}\limits_{\bm{x}\in\mathcal{X}}(h+g\circ\mathfrak{S}+\iota_{C})(\bm{x}),$ (1)
where $\iota_{C}$ is the indicator function of $C$ defined by
$\iota_{C}:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}:\bm{x}\mapsto\left\{\begin{array}[]{cl}0,&\mathrm{if}\ \bm{x}\in C;\\ +\infty,&\mathrm{if}\ \bm{x}\notin C,\end{array}\right$. (2)
and $h$, $\mathfrak{S}$, and $g$ satisfy the following:
  1. (i)
    $h:\mathcal{X}\to\mathbb{R}$ is continuously differentiable and its gradient $\nabla h:\mathcal{X}\to\mathcal{X}$ is Lipschitz continuous with a Lipschitz constant $L_{\nabla h}>0$ over $C$, i.e.,
    $(\bm{x}_{1},\bm{x}_{2}\in C)\quad\left\lVert\nabla h(\bm{x}_{1})-\nabla h(\bm{x}_{2})\right\rVert\leq L_{\nabla h}\left\lVert\bm{x}_{1}-\bm{x}_{2}\right\rVert;$ (3)
  2. (ii)
    $\mathfrak{S}:\mathcal{X}\to\mathcal{Z}$ is a continuously differentiable (possibly nonlinear) mapping;
  3. (iii)
    $g:\mathcal{Z}\to\mathbb{R}$ is a (possibly nonsmooth) Lipschitz continuous with a Lipschitz constant $L_{g}>0$, i.e.,
    $(\bm{z}_{1},\bm{z}_{2}\in\mathcal{Z})\quad\lvert g(\bm{z}_{1})-g(\bm{z}_{2})\rvert\leq L_{g}\left\lVert\bm{z}_{1}-\bm{z}_{2}\right\rVert,$ (4)
    and $\eta$-weakly convex function with $\eta>0$, i.e., $g+\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2}$ is convex over $\mathcal{Z}$. Moreover, $g$ is prox-friendly, i.e., the closed-form expression of the proximity operator (see (20)) of $g$ is available as a computable tool;
  4. (iv)
    $F$ is bounded below over $C$, i.e., $\inf\{F(\bm{x})\mid\bm{x}\in C\}>$-$\infty$.
Problem 1.1 has been serving as key models, e.g., for finding sparse solutions, in wide range of signal processing and machine learning. Such applications include, e.g., sparse Principal Component Analysis (PCA) [1], sparse variable PCA [2], robust sparse PCA [3], orthogonal dictionary learning/robust subspace recovering [4], and sparse spectral clustering [5].
One of simple cases of Problem 1.1 is a convex case where $h$ and $g$ are convex, $\mathfrak{S}$ is linear, and $C\coloneqq\mathcal{X}$. For the convex case, the so-called proximal splitting algorithms have been extensively applied as computationally reliable algorithms (see, e.g., [6, 7, 8]). The algorithms consist of simple steps involving the gradient of $h$ and the so-called proximity operator $\mathrm{prox}_{\mu g}:\mathcal{Z}\to\mathcal{Z}:\widebar{\bm{z}}\mapsto=\mathop{\mathrm{argmin}}\limits_{\bm{z}\in\mathcal{Z}}\left(g(\bm{z})+\frac{1}{2\mu}\left\lVert\bm{z}-\widebar{\bm{z}}\right\rVert^{2}\right)$ of $g$ with index $\mu>0$. Although the computation of $\mathrm{prox}_{\mu g}$ requires, in general, to solve a convex optimization problem, the proximity operators for many commonly-used functions, e.g., $\ell_{1}$-norm and nuclear norm, have closed-form expressions (see, e.g., [6, 9]).
In Problem 1.1, restricting choices of $g$ within convex functions are not always sufficient [10, 11, 12, 13, 14, 15] in order to enhance estimation performance especially for sparsity-aware applications. For example, $\ell_{1}$-norm has been widely used as a best convex approximation of the naive sparsity promoting function, i.e., $\ell_{0}$-pseudonorm, which counts the number of nonzero entries of a vector. However, the use of $\ell_{1}$-norm as $g$ leads to some underestimation of the desired sparse target because $\ell_{1}$-norm is a coercive function (see, e.g., [16]), i.e., $\ell_{1}$-norm excessively penalizes large magnitude of target. To overcome the underestimation effect caused mainly by the convex sparse promoting function, the utilization of weakly convex functions has been attracting great attention [10, 12]. Recent advances in proximal splitting techniques [11, 13, 14, 15] encourage us to find a global minimizer of Problem 1.1 under the convexity of $h+g\circ\mathfrak{S}$ with a linear operator and even with a weakly convex $g$.
Table 1: Summary of existing algorithms for Problem 1.1 and the proposed algorithm.
Method Ref. $g$ $\mathfrak{S}$ $C$ Single loop ?
Proximal linear [17] convex smooth unconstrained No
[18] weakly convex smooth unconstrained No
[19] convex smooth manifold No
Proximal gradient [20] convex linear manifold No
Subgradient [21] weakly convex identity manifold Yes
Variable smoothing [22] convex linear unconstrained Yes
[23] weakly convex linear unconstrained Yes
[24] convex linear manifold Yes
[25] weakly convex linear manifold Yes
Alg. 1 weakly convex smooth manifold Yes
(this paper)
Beyond the convex cases, nonconvex optimization problems with nonconvex constraint have a great potential for flexible formulation to align with real-world tasks. Indeed, the minimization of nonconvex $g\circ\mathfrak{S}$ with a nonlinear smooth mapping $\mathfrak{S}$ is known to have extremely wide applications including, e.g., robust phase retrieval [26], nonnegative matrix factorization [27], robust matrix recovery [28], and robust blind deconvolution [29]. In particular, manifold constrained optimization has been continuingly attracting great attention mainly by its remarkable expressive ability of the target conditions expected to be achieved by a solution (see, e.g., [30, 31, 32, 33]). Hence, Problem 1.1 is certainly of our great interest.
Literature review
We review some related works for finding a stationary point, of Problem 1.1, which satisfies a certain necessary condition called a first-order optimality condition (see (25)) to be a local minimizer. Table 1 summarizes distinctions of target settings among algorithms in the related works.
Proximal-type algorithms. The proximal gradient method (also known as an instance of the forward-backward splitting method, e.g., [6, 8]) has been extended by [20] from the unconstrained setting to the setting of Problem 1.1 with a special case where $C$ is an embedded submanifold of $\mathcal{X}$, $g$ is convex and $\mathfrak{S}$ is linear. Such an extended proximal gradient method [20] requires some iterative solvers for a certain proximal subproblem at every iteration, which is defined as convex optimization problems over the tangent (vector) space (see (89)). So far, closed-form solutions of such proximal subproblems have not yet been found even for simple cases, e.g., $g\coloneqq\left\lVert\cdot\right\rVert_{1}$. For Problem 1.1 where $\mathfrak{S}$ is smooth but $g$ remains to be convex, the prox-linear method [18, 17] for the unconstrained setting has also been extended by [19] to the manifold constrained setting. Under a certain condition, e.g., the Lipschitz continuity of $\mathrm{D}\mathfrak{S}(\cdot)$ (see (33)), the prox-linear methods [18, 17, 19] are guaranteed to converge to a stationary point by passing through a subsequence of the generated sequence. Moreover, for the prox-linear methods, the iteration complexity to obtain an $\epsilon$-approximate stationary point is given by $\mathcal{O}(\epsilon^{-2})$, where $\mathcal{O}$ stands for the Landau’s Big O notation. However, these prox-linear methods require some iterative solver for solving a convex optimization problem to update estimates $(\bm{x}_{n})_{n=1}^{\infty}\subset\mathcal{X}$.
In short, every update in the above proximal-type algorithms applicable for Problem 1.1 heavily relies on iterative solvers for solving subproblems. Convergence analyses of such proximal-type algorithms have been established basically with an exact solution to the subproblem, although such an exact solution is not available in general within a finite number of iterations of an iterative solver. Moreover, an iterative solver sometimes takes a lot of time to obtain even an approximate solution to a subproblem as will be shown in numerical experiments (see Section 5.2). These situations encourage us to develop a single-loop algorithm, such that any iterative solver is not required for solving subproblems at every iteration, for Problem 1.1.
Smoothing-type algorithms. For finding a stationary point of a nonsmooth function $J:\mathcal{X}\to\mathbb{R}$, smoothing methods have been extensively developed [34, 35, 36, 37, 38, 39, 40] with a smoothed surrogate function, say $J^{\langle\mu\rangle}:\mathcal{X}\to\mathbb{R}\ (\mu\in\mathbb{R}_{++})$, having the (sub-)gradient consistency property (see Remark 4.5) with $\lim_{\mathbb{R}_{++}\ni\mu\to 0}J^{\langle\mu\rangle}(\bm{x})=J(\bm{x})\ (\bm{x}\in\mathcal{X})$, where the gradient consistency property has been utilized as a key tool to express the first-order optimality condition in terms of a gradient sequence of $\nabla J^{\langle\mu\rangle}$ with $\mu\searrow 0$. The basic update of a smoothing gradient method [35] is given by $\bm{x}_{n+1}\coloneqq\bm{x}_{n}-\gamma_{n}\nabla J^{\langle\mu_{n}\rangle}(\bm{x}_{n})\ (n\in\mathbb{N})$ with a properly chosen stepsize $\gamma_{n}\in\mathbb{R}_{++}$, where smoothing parameters $\mu_{n}$ are conditionally updated with $\sigma\in(0,1)$ and $\delta\in\mathbb{R}_{++}$ by $\mu_{n+1}\coloneqq\sigma\mu_{n}$ if $\left\lVert\nabla J^{\langle\mu_{n}\rangle}(\bm{x}_{n})\right\rVert<\delta\mu_{n}$; and $\mu_{n+1}\coloneqq\mu_{n}$ otherwise. Thanks to the (sub-)gradient consistency property of $J^{\langle\mu\rangle}$, a cluster point of $(\bm{x}_{n})_{n=1}^{\infty}$ is guaranteed to be a (Clarke) stationary point of $J$ [35]. Based on the conditional update of smoothing parameters, many numerical algorithms, e.g., accelerated gradient methods [37, 40], projected gradient methods [34, 36] and manifold constrained smoothing methods [38, 39], for updating $(\bm{x}_{n})_{n=1}^{\infty}$ have been established with asymptotic convergence analyses in terms of a stationary point (or the optimal value of the cost function for convex optimization). In contrast, convergence rate analyses of smoothing methods have not been found except for a few papers, e.g., [41, 42] for an optimization problem involving $\ell_{p}$-quasinorm $(0<p\leq 1)$ (see Remark 4.10).
Another class of smoothing-type methods has been proposed with iteratively decreasing smoothing parameters $(\mu_{n})_{n=1}^{\infty}$ [22, 43, 44, 23, 45, 46, 24, 25, 47], where these methods are called variable/adaptive/dynamic smoothing methods. Thanks to the use of iteratively decreasing smoothing parameters, convergence rate analyses have been established for variable smoothing methods. For convex optimization problems, convergence rates in terms of the optimal value of the cost function have been shown for variable smoothing algorithms [22, 43, 44, 45, 46, 47], where an exceptional paper [46] guarantees that a generated sequence converges to a minimizer of the cost function.
Recently, for nonconvex nonsmooth optimization problems, a variable smoothing algorithm [22] has been extended by [23] to a special version of Problem 1.1 where $\mathfrak{S}$ is a linear operator and $C\coloneqq\mathcal{X}$. The variable smoothing algorithm in [23] uses a smoothed surrogate function $h+{}^{\mu}g\circ\mathfrak{S}$ with the Moreau envelope $^{\mu}g$ of $g$ (see (21)), and its convergence rate, i.e., iteration complexity, $\mathcal{O}(\epsilon^{-3})$ to obtain a certain $\epsilon$-approximate stationary point has been shown under the surjective linearity of $\mathfrak{S}$. This variable smoothing algorithm [23] with its convergence rate analysis has also been extended by [24, 25] to Problem 1.1 where $C$ is a compact embedded submanifold of $\mathcal{X}$, but $\mathfrak{S}$ remains to be linear. In contrast, for variable smoothing algorithms [23, 25] except [24], any asymptotic convergence analysis in terms of a stationary point has not been reported yet, where [24] presents an asymptotic convergence analysis by exploiting the compactness of $C$. Hence, it remains an open issue to establish an asymptotic convergence guarantee of the variable smoothing algorithms for Problem 1.1 even with an unconstrained case $C\coloneqq\mathcal{X}$ or case of a nonlinear differentiable mapping $\mathfrak{S}$.
Contributions of this paper
In this paper, we address the research question: Can we establish a single-loop algorithm for Problem 1.1 with an asymptotic convergence analysis and an iteration complexity of $\mathcal{O}(\epsilon^{-3})$ to obtain an $\epsilon$-approximate stationary point ? To answer affirmatively, we extend a variable smoothing algorithm [23] for Problem 1.1 even with a nonlinear differentiable mapping. Contributions of this paper are three-fold as follows.
  1. (i)
    Parametrization strategy for nonsmooth minimization with manifold constraint. Motivated by parametrization strategies, e.g., [48, 49, 50, 51, 52, 53, 54, 55], we introduce to translate Problem 1.1 into the following unconstrained nonsmooth optimization problem by assuming a manifold constraint set $C$ can be parameterized by a continuously differentiable mapping $\bm{\varphi}:\mathcal{Y}\to C(\subset\mathcal{X})$ defined over a Euclidean space $\mathcal{Y}$ such that $\bm{\varphi}(\mathcal{Y})\coloneqq\{\bm{\varphi}(\bm{y})\in\mathcal{X}\mid\bm{y}\in\mathcal{Y}\}=C$ (see Example 1.3 below for pairs of $C$ and $\bm{\varphi}$).
    Problem 1.2.
    Consider $F=h+g\circ\mathfrak{S}$ and $C\subset\mathcal{X}$ in Problem 1.1. Let $\bm{\varphi}:\mathcal{Y}\to\mathcal{X}$ be a continuously differentiable mapping such that $C=\bm{\varphi}(\mathcal{Y})$, where $\mathcal{Y}$ is a Euclidean space and $\bm{\varphi}$ is called a parametrization of $C$. Then,
    $\mathrm{find}\ \bm{y}^{\star}\in\mathop{\mathrm{argmin}}\limits_{\bm{y}\in\mathcal{Y}}(h+g\circ\mathfrak{S})\circ\bm{\varphi}(\bm{y})(=F\circ\bm{\varphi}(\bm{y})).$ (5)
    Example 1.3 (On parametrization; see Section 5 for other examples).
    1. (a)
      Let $C$ be a linear subspace of $\mathcal{X}$. Consider a linear operator $L:\mathcal{Y}\to\mathcal{X}$ satisfying $L(\mathcal{Y})\coloneqq\{L\bm{y}\in\mathcal{X}\mid\bm{y}\in\mathcal{Y}\}=C$. Then, $\bm{\varphi}\coloneqq L$ is a parametrization of $C$. In particular, $\bm{\varphi}\coloneqq\mathrm{Id}$ is a parametrization of $C$ if $C\coloneqq\mathcal{X}$.
    2. (b)
      For $\mathcal{X}\coloneqq\mathbb{R}^{2}$, let $C\coloneqq\{\bm{x}\in\mathbb{R}^{2}\mid\left\lVert\bm{x}\right\rVert=1\}$ be the unit circle. Consider $\bm{\varphi}:\mathbb{R}\to\mathbb{R}^{2}:\theta\mapsto[\cos(\theta)\ \sin(\theta)]^{\mathsf{T}}$. Then, $\bm{\varphi}$ is a parametrization of $C$.
    3. (c)
      Let $C\subset\mathcal{X}$ be an embedded submanifold of $\mathcal{X}$, endowed with the Riemannian metric induced from $\mathcal{X}$, such that $C$ is connected and closed in $\mathcal{X}$, and let $T_{C}(\bm{x})\subset\mathcal{X}$ be a tangent (vector) space to $C$ at $\bm{x}\in C$ (see (16)). In this case, for any $\bm{x}\in C$, $\mathrm{Exp}_{\bm{x}}:T_{C}(\bm{x})\to C$ with the exponential map [^1] By [32, Defs. 10.6 and 10.7, Thm. 10.8, Exe. 10.12] and [56, Thm. 4.27], the exponential map is uniquely defined over the tangent bundle $TC\coloneqq\{(\bm{x},\bm{v})\mid\bm{x}\in C,\bm{v}\in T_{C}(\bm{x})\}$ as $\mathrm{Exp}:TC\to C:(\bm{x},\bm{v})\mapsto\mathrm{Exp}_{\bm{x}}(\bm{v})\coloneqq\gamma_{\bm{v}}(1)$ with the geodesics $\gamma_{\bm{v}}:I(\subset\mathbb{R})\to C$ (see [32, Def. 5.38][56, (4.16)]) such that $\gamma_{\bm{v}}(0)=\bm{x}$, $\gamma_{\bm{v}}^{\prime}(0)=\bm{v}$ and the geodesic equation $\gamma_{\bm{v}}^{\prime\prime}(t)=0$ is satisfied for all $t\in I$, where $I\subset\mathbb{R}$ is an interval containing $[0,1]$. $\mathrm{Exp}$ is continuously differentiable by [56, Prop. 5.19], and surjective onto $C$ by [56, Cor. 6.21] together with the completeness (closedness in $\mathcal{X}$) of $C$. Hence, $\mathrm{Exp}_{\bm{x}}$ is a parametrization of $C$. See [56, p.136-p.142] for examples of the exponential map of some manifolds.
    In order to validate the idea of parametrization of $C$, we present sufficient conditions, in Theorem 3.2 and Corollary 3.4, for the equivalence between Problem 1.1 and Problem 1.2 in view of finding their stationary points.
  2. (ii)
    Proposal of a smoothed surrogate function of the nonsmooth function $F\circ\bm{\varphi}$ in Problem 1.2. We introduce a smoothed surrogate function $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ of $F\circ\bm{\varphi}=(h+g\circ\mathfrak{S})\circ\bm{\varphi}$ in Section 4.1 by using the Moreau envelope $^{\mu}g:\mathcal{Z}\to\mathbb{R}\ (\mu\in(0,\eta^{-1}))$ of $g$ (see (21)). In Theorem 4.4 (see also Remark 4.5 (a)), we show that $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ satisfies the gradient consistency property for the original function $F\circ\bm{\varphi}$, where the gradient consistency property will play a key role for establishing an asymptotic convergence property of the proposed algorithm.
    Remark 1.4 (Integral convolution type smoothed surrogate function).
    We can construct another smoothed surrogate function, with gradient consistency property, of $F\circ\bm{\varphi}$ via the integral convolution of $F\circ\bm{\varphi}$ and a kernel function, called mollifier (see [57, Exm. 7.19, Thm. 9.67] or [58, Thm. 4.7]), because $F\circ\bm{\varphi}$ is subdifferentially regular (see Lemma 2.4). However, it is questionable to obtain closed-form expressions of this type of smoothed surrogate functions for $F\circ\bm{\varphi}$ and its gradient, except for simple functions, e.g., the plus function $\mathbb{R}\ni x\mapsto\max\{x,0\}$ [59]. In contrast, the closed-form expressions of our smoothed surrogate function $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ and its gradient are available as computable tools in a case where $g$ is prox-friendly. Moreover, our smoothed surrogate function $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ with a linear $\mathfrak{S}$ and $\bm{\varphi}\coloneqq\mathrm{Id}$ has been employed in the existing variable smoothing algorithms [23, 24, 25], although the gradient consistency property has not been discussed. These situations motivate us to investigate the gradient consistency property of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ in Theorem 4.4.
  3. (iii)
    Extension of variable smoothing algorithms for Problem 1.2 with asymptotic convergence analysis. For finding a stationary point of Problem 1.1 via Problem 1.2, we extend a variable smoothing algorithm [23] (see Algorithm 1 in Section 4.2), where the proposed algorithm is designed as a single-loop algorithm. For the proposed algorithm, Theorem 4.9 (c) presents an asymptotic convergence analysis in terms of a stationary point, with the aid of the gradient consistency property of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$. Moreover, Theorem 4.9 (d) presents an iteration complexity of $\mathcal{O}(\epsilon^{-3})$ to obtain an $\epsilon$-approximate stationary point (see also Remark 4.10). Our asymptotic convergence analysis can also apply to the existing variable smoothing algorithm [23] for the unconstrained case because [23] is recovered by the proposed algorithm with a linear operator $\mathfrak{S}$ and $\bm{\varphi}\coloneqq\mathrm{Id}$. Finally, numerical experiments, in Section 5, demonstrate the efficacy of the proposed algorithm in two scenarios of the sparse spectral clustering and the sparse principal component analysis.
A preliminary short version of this paper was partially presented in [60].
Notation $\mathbb{N}$, $\mathbb{R}$, $\mathbb{R}_{+}$, and $\mathbb{R}_{++}$ denote respectively the sets of all positive integers, all real numbers, all nonnegative real numbers, and all positive real numbers. ${\left\langle\cdot,\cdot\right\rangle}_{\mathcal{X}}$ and $\left\lVert\cdot\right\rVert_{\mathcal{X}}$ stand respectively for the standard inner product and its induced norm defined in the Euclidean space $\mathcal{X}$, i.e., $\left\lVert\bm{x}\right\rVert_{\mathcal{X}}=\sqrt{{\left\langle\bm{x},\bm{x}\right\rangle}_{\mathcal{X}}}\ (\bm{x}\in\mathcal{X})$. We simply use $\left\langle\cdot,\cdot\right\rangle$} and $\left\lVert\cdot\right\rVert$ without any subscript if their domains are clear. The symbol $\mathrm{Id}$ stands for the identity operator. For a linear operator $A:\mathcal{X}\to\mathcal{Y}$, its adjoint operator $A^{*}:\mathcal{Y}\to\mathcal{X}$ is defined as the linear operator satisfying ${\left\langle\bm{x},A^{*}\bm{y}\right\rangle}_{\mathcal{X}}={\left\langle A\bm{x},\bm{y}\right\rangle}_{\mathcal{Y}}\ (\forall\bm{x}\in\mathcal{X},\bm{y}\in\mathcal{Y})$, $\mathrm{Ker}(A)\coloneqq\{\bm{x}\in\mathcal{X}\mid A\bm{x}=\bm{0}\}$ and $\mathrm{Ran}(A)\coloneqq\{A\bm{x}\in\mathcal{Y}\mid\bm{x}\in\mathcal{X}\}$ are respectively the kernel (null) space and the range space of $A$, and $\left\lVert A\right\rVert_{\rm op}\coloneqq\sup_{\left\lVert\bm{x}\right\rVert_{\mathcal{X}}\leq 1,\bm{x}\in\mathcal{X}}\left\lVert A\bm{x}\right\rVert_{\mathcal{Y}}$ denotes the operator norm of $A$. The symbol $(\cdot)^{\perp}$ denotes the orthogonal complement subspace of a given set. For a matrix $\bm{X}\in\mathbb{R}^{m\times n}$, $\bm{X}]_{i,j}$ denotes the $(i,j)$ entry of $\bm{X}$, and $\bm{X}^{\mathsf{T}}$ denotes the transpose of $\bm{X}$. For a given point $\widebar{\bm{x}}\in\mathcal{X}$ and a given set $E\subset\mathcal{X}$, $d:\mathcal{X}\times 2^{\mathcal{X}}\to\mathbb{R}_{+}\cup\{+\infty\}$ stands for the distance function, i.e.,
$d(\widebar{\bm{x}},E)\coloneqq\left\{\begin{array}[]{cl}\inf\{\left\lVert\bm{v}-\widebar{\bm{x}}\right\rVert\mid\bm{v}\in E\},&\mathrm{if}\ E\neq\emptyset;\\ +\infty,&\mathrm{if}\ E=\emptyset.\end{array}\right$. (6)
For a given $(a_{n})_{n=1}^{\infty}\subset\mathbb{R}$ and a given $\widebar{a}\in\mathbb{R}$, $a_{n}\searrow\widebar{a}$ and $(a_{n})_{n=1}^{\infty}\searrow\widebar{a}$ mean in this paper that $(a_{n})_{n=1}^{\infty}$ is monotonically nonincreasing together with $\lim_{n\to\infty}a_{n}=\widebar{a}$.
For a differentiable mapping $\mathcal{F}:\mathcal{Y}\to\mathcal{X}$, its Gâteaux derivative at $\bm{y}\in\mathcal{Y}$ is the linear operator $\mathrm{D}\mathcal{F}(\bm{y}):\mathcal{Y}\to\mathcal{X}$ defined by
$(\bm{v}\in\mathcal{Y})\quad\mathrm{D}\mathcal{F}(\bm{y})[\bm{v}]=\lim_{\mathbb{R}\setminus\{0\}\ni t\to 0}\frac{\mathcal{F}(\bm{y}+t\bm{v})-\mathcal{F}(\bm{y})}{t}.$ (7)
For a differentiable function $J:\mathcal{X}\to\mathbb{R}$, $\nabla J(\bm{x})\in\mathcal{X}$ is the gradient of $J$ at $\bm{x}\in\mathcal{X}$ if $\mathrm{D}J(\bm{x})[\bm{v}]={\left\langle\nabla J(\bm{x}),\bm{v}\right\rangle}$ for all $\bm{v}\in\mathcal{X}$. Note that for a differentiable mapping $\mathcal{F}:\mathcal{Y}\to\mathcal{X}$, we have the expression [^2] It can be checked by $(\forall\bm{v}\in\mathcal{Y})\ {\left\langle\nabla(J\circ\mathcal{F})(\bm{x}),\bm{v}\right\rangle}_{\mathcal{Y}}=\mathrm{D}(J\circ\mathcal{F})(\bm{x})[\bm{v}]=\mathrm{D}J(\mathcal{F}(\bm{x}))[\mathrm{D}\mathcal{F}(\bm{x})[\bm{v}]]={\left\langle\nabla J(\mathcal{F}(\bm{x})),\mathrm{D}\mathcal{F}(\bm{x})[\bm{v}]\right\rangle}_{\mathcal{X}}={\left\langle(\mathrm{D}\mathcal{F}(\bm{x}))^{*}[\nabla J(\mathcal{F}(\bm{x}))],\bm{v}\right\rangle}_{\mathcal{Y}}$. $\nabla(J\circ\mathcal{F})(\bm{y})=(\mathrm{D}\mathcal{F}(\bm{y}))^{*}[\nabla J(\mathcal{F}(\bm{y}))]\ (\bm{y}\in\mathcal{Y})$. For a function $J:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$, $\mathrm{dom}(J)\coloneqq\{\bm{x}\in\mathcal{X}\mid J(\bm{x})<+\infty\}\subset\mathcal{X}$ and $\mathrm{epi}(J):=\{(\bm{x},a)\in\mathcal{X}\times\mathbb{R}\mid J(\bm{x})\leq a\}\subset\mathcal{X}\times\mathbb{R}$ denote respectively the domain and the epigraph of $J$. Moreover, $J$ is said to be (i) proper if $\mathrm{dom}(J)\neq\emptyset$; (ii) lower semicontinuous if $\mathrm{epi}(J)$ is closed in $\mathcal{X}\times\mathbb{R}$; (iii) locally Lipschitz continuous at $\widebar{\bm{x}}\in\mathcal{X}$ if there exist an open neighborhood $\mathcal{N}_{\widebar{\bm{x}}}\subset\mathcal{X}$ of $\widebar{\bm{x}}$ and $L>0$ such that $\lvert J(\bm{x}_{1})-J(\bm{x}_{2})\rvert\leq L\left\lVert\bm{x}_{1}-\bm{x}_{2}\right\rVert\ (\bm{x}_{1},\bm{x}_{2}\in\mathcal{N}_{\widebar{\bm{x}}})$. For a function $J:\mathcal{X}\to\mathbb{R}\cup\{-\infty,+\infty\}$, the limit inferior [57, Definition 1.5] and the limit superior [57, p.13] of $J$ at $\widebar{\bm{x}}\in\mathcal{X}$ are defined respectively by
$\liminf_{\mathcal{X}\ni\bm{x}\to\widebar{\bm{x}}}J(\bm{x})$ $\coloneqq\sup_{\epsilon>0}\left(\inf_{\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert<\epsilon}J(\bm{x})\right)\coloneqq\sup\limits_{\epsilon>0}\left(\inf\left\{J(\bm{x})\in\mathbb{R}\cup\{-\infty,+\infty\}\mid\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert<\epsilon\right\}\right);$ (8)
$\limsup_{\mathcal{X}\ni\bm{x}\to\widebar{\bm{x}}}J(\bm{x})$ $\coloneqq\inf_{\epsilon>0}\left(\sup_{\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert<\epsilon}J(\bm{x})\right)\coloneqq\inf\limits_{\epsilon>0}\left(\sup\left\{J(\bm{x})\in\mathbb{R}\cup\{-\infty,+\infty\}\mid\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert<\epsilon\right\}\right).$ (9)
For a set sequence $(E_{n})_{n=1}^{\infty}$ with subsets $E_{n}\subset\mathcal{X}$, the outer limit [57, Definition 4.1] of $(E_{n})_{n=1}^{\infty}$ is defined by
$\mathop{\mathrm{Lim~sup}}_{n\to\infty}E_{n}\coloneqq\left\{\bm{v}\in\mathcal{X}\mid\exists\mathcal{N}\subset\mathbb{N},\exists\bm{v}_{n}\in E_{n}\ (n\in\mathcal{N})\ {\rm with}\ \lvert\mathcal{N}\rvert=\aleph_{0}\ {\rm and}\ \bm{v}=\lim_{\mathcal{N}\ni n\to\infty}\bm{v}_{n}\right\},$ (10)
where $\aleph_{0}$ stands for the cardinality of $\mathbb{N}$. With the outer limit of a set sequence, the outer limit [57, p.152] of a set-valued mapping $S:\mathcal{X}\rightrightarrows\mathcal{X}$ at $\widebar{\bm{x}}\in\mathcal{X}$ is defined as
$\mathop{\mathrm{Lim~sup}}_{\mathcal{X}\ni\bm{x}\to\widebar{\bm{x}}}S(\bm{x})\coloneqq\bigcup_{\mathcal{X}\ni\bm{x}_{n}\to\widebar{\bm{x}}}\mathop{\mathrm{Lim~sup}}_{n\to\infty}S(\bm{x}_{n})$ (11)
$=\left\{\bm{v}\in\mathcal{X}\mid\exists(\bm{x}_{n})_{n=1}^{\infty}\subset\mathcal{X},\ \exists\bm{v}_{n}\in S(\bm{x}_{n})\ \mathrm{with}\ \widebar{\bm{x}}=\lim_{n\to\infty}\bm{x}_{n}\ \mathrm{and}\ \bm{v}=\lim_{n\to\infty}\bm{v}_{n}\right\},$ (12)
where we follow a clear notation $\mathop{\mathrm{Lim~sup}}$” (used, e.g., in [61]). In this paper, even for the outer limit of a single-valued mapping $S:\mathcal{X}\to\mathcal{X}$, we use $\mathop{\mathrm{Lim~sup}}_{\mathcal{X}\ni\bm{x}\to\widebar{\bm{x}}}S(\bm{x})\coloneqq\mathop{\mathrm{Lim~sup}}_{\mathcal{X}\ni\bm{x}\to\widebar{\bm{x}}}\{S(\bm{x})\}$ in order to distinguish the standard notation for the limit superior $\limsup$” in (9).
2 Preliminaries
2.1 Subdifferential and Normal Cone
We review some necessary notions, namely subdifferentials and normal cones, in nonsmooth analysis. We refer to the seminal book [57] for a thorough review.
Definition 2.1 (Subdifferential [57, Def. 8.3]).
For a possibly nonconvex function $J:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$, the limiting (or general) subdifferential of $J$ at $\widebar{\bm{x}}\in\mathcal{X}$ is defined by
$(\widebar{\bm{x}}\in\mathrm{dom}(J))\quad\partial_{\rm L}J(\widebar{\bm{x}})\coloneqq\mathop{\mathrm{Lim~sup}}_{\mathcal{X}\ni\bm{x}\to\widebar{\bm{x}},\ J(\bm{x})\to J(\widebar{\bm{x}})}\partial_{\rm F}J(\bm{x}),$ (13)
where $\partial_{\rm F}J(\widebar{\bm{x}})\coloneqq\left\{\bm{v}\in\mathcal{X}\mid\liminf_{\mathcal{X}\setminus\{\widebar{\bm{x}}\}\ni\bm{x}\to\widebar{\bm{x}}}\frac{J(\bm{x})-J(\widebar{\bm{x}})-{\left\langle\bm{v},\bm{x}-\widebar{\bm{x}}\right\rangle}}{\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert}\geq 0\right\}$ is called the Fréchet (or regular) subdifferential of $J$ at $\widebar{\bm{x}}\in\mathrm{dom}(J)$, and $\partial_{\rm L}J(\widebar{\bm{x}})$ and $\partial_{\rm F}J(\widebar{\bm{x}})$ at $\widebar{\bm{x}}\notin\mathrm{dom}(J)$ are understood as $\emptyset$ (Note: for a convex function $J$, $\partial_{\rm L}J$ and $\partial_{\rm F}J$ are identical to the standard convex subdifferential of $J$ [57, Prop. 8.12]).
By using the subdifferentials, the general normal cone $N_{C}(\widebar{\bm{x}})$ and the regular normal cone $\widehat{N}_{C}(\widebar{\bm{x}})$ to a nonempty subset $C\subset\mathcal{X}$ at $\widebar{\bm{x}}\in C$ are defined (see [57, Def. 6.3, Prop. 6.5, and Ex. 8.14]) respectively by
$N_{C}(\widebar{\bm{x}})$ $\coloneqq\partial_{\rm L}\iota_{C}(\widebar{\bm{x}})=\mathop{\mathrm{Lim~sup}}_{C\ni\bm{x}\to\widebar{\bm{x}}}\partial_{\rm F}\iota_{C}(\widebar{\bm{x}})=\mathop{\mathrm{Lim~sup}}_{C\ni\bm{x}\to\widebar{\bm{x}}}\widehat{N}_{C}(\widebar{\bm{x}})$ (14)
$\widehat{N}_{C}(\widebar{\bm{x}})$ $\coloneqq\partial_{\rm F}\iota_{C}(\widebar{\bm{x}})=\left\{\bm{v}\in\mathcal{X}\mid\limsup_{C\setminus\{\widebar{\bm{x}}\}\ni\bm{x}\to\widebar{\bm{x}}}\frac{{\left\langle\bm{v},\bm{x}-\widebar{\bm{x}}\right\rangle}}{\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert}\leq 0\right\},$ (15)
where $N_{C}(\widebar{\bm{x}})$ and $\widehat{N}_{C}(\widebar{\bm{x}})$ at $\widebar{\bm{x}}\notin C$ are understood as $\emptyset$.
With notions of normal cone operators, we define regularities of a subset and functions as follows. Such regularities ensure the chain and sum rules for the limiting subdifferential as shown in Lemma 2.4 below (see also Fact A.4 in Appendix A).
Definition 2.2 (Regularity).
  1. (a)
    (Clarke regularity [57, Def. 6.4]) A nonempty set $C\subset\mathcal{X}$ is said to be Clarke regular at $\widebar{\bm{x}}\in C$ if (i) $N_{C}(\widebar{\bm{x}})=\widehat{N}_{C}(\widebar{\bm{x}})$ (Note: the inclusion $N_{C}(\widebar{\bm{x}})\supset\widehat{N}_{C}(\widebar{\bm{x}})$ always holds), and (ii) $C\cap\overline{\mathcal{N}}_{\widebar{\bm{x}}}$ is closed for some closed neighborhood [^3] In other words, $\overline{\mathcal{N}}_{\widebar{\bm{x}}}$ is a closed set that contains $\widebar{\bm{x}}$ as an interior point of $\overline{\mathcal{N}}_{\widebar{\bm{x}}}$. $\overline{\mathcal{N}}_{\widebar{\bm{x}}}\subset\mathcal{X}$ of $\widebar{\bm{x}}$. Moreover, $C$ is said to be Clarke regular if it is Clarke regular at every $\widebar{\bm{x}}\in C$.
  2. (b)
    (Subdifferential regularity [57, Def. 7.25]) A proper lower semicontinuous function $J:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ is said to be subdifferentially regular at $\widebar{\bm{x}}\in\mathrm{dom}(J)$ if $\mathrm{epi}(J)\subset\mathcal{X}\times\mathbb{R}$ is Clarke regular at $(\widebar{\bm{x}},J(\widebar{\bm{x}}))$. Moreover, $J$ is said to be subdifferentially regular if it is subdifferentially regular at every $\widebar{\bm{x}}\in\mathrm{dom}(J)$. We note that all functions in this paper are shown to be subdifferentially regular (see, e.g., Lemma 2.4 below and also Example A.2).
A typical Clarke regular set is a closed convex set [57, Thm. 6.9]. Another important example of Clarke regular sets is an embedded submanifold of the Euclidean space $\mathcal{X}$ [57, Exm. 6.8] as follows.
Example 2.3 (Embedded submanifold of $\mathcal{X}$ [57, Exm. 6.8]).
Let $(\emptyset\neq)C\subset\mathcal{X}$ be a $d(\leq\dim(\mathcal{X}))$-dimensional embedded submanifold of $\mathcal{X}$, i.e., for each $\widebar{\bm{x}}\in C$, there exist (i) an open neighborhood $\mathcal{N}_{\widebar{\bm{x}}}\subset\mathcal{X}$ of $\widebar{\bm{x}}$ in $\mathcal{X}$ and (ii) a continuously differentiable mapping $H_{\widebar{\bm{x}}}:\mathcal{N}_{\widebar{\bm{x}}}\to\mathbb{R}^{\dim(\mathcal{X})-d}$, such that $C\cap\mathcal{N}_{\widebar{\bm{x}}}=H_{\widebar{\bm{x}}}^{-1}(\bm{0})$ and $\mathrm{Rank}(\mathrm{D}H_{\widebar{\bm{x}}}(\widebar{\bm{x}}))=\dim(\mathcal{X})-d$ (Note: by Whitney’s embedding theorem (see, e.g., [62, Thm. 6.15]), every $d$-dimensional smooth manifold can be seen as an embedded submanifold in $\mathbb{R}^{2d+1}$). Then, by [57, Def. 6.4 and Exm. 6.8], the embedded submanifold $C$ is Clarke regular, and we have the following expression:
$(\widebar{\bm{x}}\in C)\quad N_{C}(\widebar{\bm{x}})=\widehat{N}_{C}(\widebar{\bm{x}})=\mathrm{Ran}((\mathrm{D}H_{\widebar{\bm{x}}}(\widebar{\bm{x}}))^{*})\subset\mathcal{X},\quad T_{C}(\widebar{\bm{x}})=\left(N_{C}(\widebar{\bm{x}})\right)^{\perp}\subset\mathcal{X},$ (16)
where the orthogonal complement subspace $T_{C}(\widebar{\bm{x}})$ of $N_{C}(\widebar{\bm{x}})$ is called the tangent cone to $C$ at $\widebar{\bm{x}}$ (Note: $N_{C}(\widebar{\bm{x}})$ and $T_{C}(\widebar{\bm{x}})$ are also called respectively the normal space and the tangent space to the manifold $C$ at $\widebar{\bm{x}}\in C$ [32, Thm. 3.15, Def. 3.16, and p.107]).
Lemma 2.4 (Subdifferential calculus for Problems 1.1 and 1.2).
Consider $F=h+g\circ\mathfrak{S}$, $C\subset\mathcal{X}$ and $\bm{\varphi}$ in Problems 1.1 and 1.2. Then, $F$, $F+\iota_{C}$ and $F\circ\bm{\varphi}$ are subdifferentially regular, and we have the following expressions:
$(\widebar{\bm{x}}\in C)\quad$ $\partial_{\rm L}(F+\iota_{C})(\widebar{\bm{x}})=\partial_{\rm L}F(\widebar{\bm{x}})+N_{C}(\widebar{\bm{x}});$ (17)
$(\widebar{\bm{y}}\in\mathcal{Y})\quad$ $\partial_{\rm L}(F\circ\bm{\varphi})(\widebar{\bm{y}})=(\mathrm{D}\bm{\varphi}(\widebar{\bm{y}}))^{*}[\partial_{\rm L}F(\bm{\varphi}(\widebar{\bm{y}}))].$ (18)
Proof.
See Appendix A. ∎
The following states a relation between the normal cone to the parameterized set, and the parametrization $\bm{\varphi}$.
Fact 2.5 (Normal cone to image sets [57, Thm. 6.43]).
Let $\mathcal{X},\mathcal{Y}$ be Euclidean spaces, and $\bm{\varphi}:\mathcal{Y}\to\mathcal{X}$ a continuously differentiable mapping. Then, for every $\widebar{\bm{x}}\in\bm{\varphi}(\mathcal{Y})$,
$\widehat{N}_{\bm{\varphi}(\mathcal{Y})}(\widebar{\bm{x}})\subset\bigcap_{\widebar{\bm{y}}\in\bm{\varphi}^{-1}(\widebar{\bm{x}})}\left\{\bm{v}\in\mathcal{X}\mid(\mathrm{D}\bm{\varphi}(\widebar{\bm{y}}))^{*}[\bm{v}]\in\widehat{N}_{\mathcal{Y}}(\widebar{\bm{y}})=\{\bm{0}\}\right\}=\bigcap_{\widebar{\bm{y}}\in\bm{\varphi}^{-1}(\widebar{\bm{x}})}\mathrm{Ker}\left((\mathrm{D}\bm{\varphi}(\widebar{\bm{y}}))^{*}\right),$ (19)
where $\bm{\varphi}^{-1}(\widebar{\bm{x}})\coloneqq\{\widebar{\bm{y}}\in\mathcal{Y}\mid\bm{\varphi}(\widebar{\bm{y}})=\widebar{\bm{x}}\}\subset\mathcal{Y}$ is the preimage of $\widebar{\bm{x}}\in\mathcal{X}$ under $\bm{\varphi}$.
Proof.
The first inclusion is verified by [57, Thm. 6.43]. The equality follows by the relation $(\mathrm{D}\bm{\varphi}(\widebar{\bm{y}}))^{*}[\bm{v}]\in\{\bm{0}\}\Leftrightarrow\bm{v}\in\mathrm{Ker}\left((\mathrm{D}\bm{\varphi}(\widebar{\bm{y}}))^{*}\right)\ (\widebar{\bm{y}}\in\bm{\varphi}^{-1}(\widebar{\bm{x}}))$. ∎
2.2 Proximity Operator and Moreau Envelope
The proximity operator and the Moreau envelope have been serving as key ingredients for minimization of nonsmooth convex functions (see, e.g., [6, 7, 8]). Even for weakly convex functions, they are useful in design of nonsmooth optimization algorithms. We note that closed-form expressions of the proximity operators for commonly-used weakly convex functions have been found, e.g., $\ell_{1}$-norm [6], Minimax Concave Penalty (MCP) [16], Smoothly Clipped Absolute Deviation (SCAD) [63], log-sum penalty [64], piece-wise exponential function [65], and arctangent penalty [66].
Definition 2.6 (Proximity operator and Moreau envelope [57, Def. 1.22]).
Let $g:\mathcal{Z}\to\mathbb{R}\cup\{+\infty\}$ be proper lower semicontinuous, and $\eta$-weakly convex, i.e., $g+\eta/2\left\lVert\cdot\right\rVert^{2}$ is convex with $\eta>0$. For $\mu\in(0,\eta^{-1})$, the proximity operator and the Moreau envelope of $g$ of index $\mu$ are defined respectively as
$(\widebar{\bm{z}}\in\mathcal{Z})\$$\mathrm{prox}_{\mu g}(\widebar{\bm{z}})\coloneqq\mathop{\mathrm{argmin}}\limits_{\bm{z}\in\mathcal{Z}}\left(g(\bm{z})+\frac{1}{2\mu}\left\lVert\bm{z}-\widebar{\bm{z}}\right\rVert^{2}\right);$ (20)
$(\widebar{\bm{z}}\in\mathcal{Z})\$${}^{\mu}g(\widebar{\bm{z}})\coloneqq\inf_{\bm{z}\in\mathcal{Z}}\left(g(\bm{z})+\frac{1}{2\mu}\left\lVert\bm{z}-\widebar{\bm{z}}\right\rVert^{2}\right)=g(\mathrm{prox}_{\mu g}(\widebar{\bm{z}}))+\frac{1}{2\mu}\left\lVert\mathrm{prox}_{\mu g}(\widebar{\bm{z}})-\widebar{\bm{z}}\right\rVert^{2},$ (21)
where $\mathrm{prox}_{\mu g}(\widebar{\bm{z}})$ is single-valued because the unique existence of the minimizer is guaranteed by the strong convexity of $g(\cdot)+\frac{1}{2\mu}\left\lVert\cdot-\widebar{\bm{z}}\right\rVert^{2}$.
Fact 2.7 (Properties of Moreau envelope of weakly convex function).
Let $g:\mathcal{Z}\to\mathbb{R}\cup\{+\infty\}$ be proper lower semicontinuous, and $\eta$-weakly convex with some $\eta>0$. Then, the following hold:
  1. (a)
    For any $\widebar{\bm{z}}\in\mathcal{Z}$ and $(\bm{z}_{n})_{n=1}^{\infty}\subset\mathcal{Z}$ satisfying $\lim_{n\to\infty}\bm{z}_{n}=\widebar{\bm{z}}$, we have $\lim_{n\to\infty}\mathrm{prox}_{\mu_{n}g}(\bm{z}_{n})=P_{\overline{\mathrm{dom}(\partial_{\rm L}g)}}(\widebar{\bm{z}})$ if $(\mu_{n})_{n=1}^{\infty}\searrow 0$ with $\mu_{n}\in(0,\eta^{-1})$ [67, Lemma 2.3], where $\mathrm{dom}(\partial_{\rm L}g)\coloneqq\{\bm{z}\in\mathcal{Z}\mid\partial_{\rm L}g(\bm{z})\neq\emptyset\}\subset\mathcal{Z}$, $\overline{(\cdot)}$ denotes the closure of a given set, and $P_{\overline{\mathrm{dom}(\partial_{\rm L}g)}}=\mathrm{prox}_{\iota_{\overline{\mathrm{dom}(\partial_{\rm L}g)}}}$ denotes the projection mapping onto the closed convex set $\overline{\mathrm{dom}(\partial_{\rm L}g)}\subset\mathcal{Z}$.
  2. (b)
    For any $\mu\in(0,\eta^{-1})$, (i) $^{\mu}g:\mathcal{Z}\to\mathbb{R}$ is continuously differentiable with $\nabla{}^{\mu}g=\frac{\mathrm{Id}-\mathrm{prox}_{\mu g}}{\mu}$, and (ii) $\nabla{}^{\mu}g$ is Lipschitz continuous with a Lipschitz constant $\max\left\{\mu^{-1},\frac{\eta}{1-\eta\mu}\right\}$ [68, Cor. 3.4].
  3. (c)
    $(\widebar{\bm{z}}\in\mathcal{Z},\mu\in(0,\eta^{-1}))\quad\nabla{}^{\mu}g(\widebar{\bm{z}})\in\partial_{\rm L}g(\mathrm{prox}_{\mu g}(\widebar{\bm{z}}))$ [23, Lemma 3.2].
  4. (d)
    Assume that $g$ is Lipschitz continuous with a Lipschitz constant $L_{g}>0$. By [23, Lemma 3.3] and [69, Prop. 3.4] [^4] In [69, Prop. 3.4], the inequality (23) is proved under the convexity of $g$. However, the proof in [69, Prop. 3.4] can be applied to the weakly convex case as well. , the following hold:
    $(\widebar{\bm{z}}\in\mathcal{Z},\mu\in(0,\eta^{-1}))\quad$ $\left\lVert\nabla{}^{\mu}g(\widebar{\bm{z}})\right\rVert\leq L_{g};$ (22)
    $(\widebar{\bm{z}}\in\mathcal{Z},\mu\in(0,\eta^{-1}))\quad$ ${}^{\mu}g(\widebar{\bm{z}})\leq g(\widebar{\bm{z}})\leq{}^{\mu}g(\widebar{\bm{z}})+\frac{1}{2}\mu L_{g}^{2}.$ (23)
    Moreover, for $0<\mu_{2}<\mu_{1}<\eta^{-1}$, we have [57, Thm. 1.25], [23, Lemma 4.1]
    $(\widebar{\bm{z}}\in\mathcal{Z})\quad{}^{\mu_{1}}g(\widebar{\bm{z}})\leq{}^{\mu_{2}}g(\widebar{\bm{z}})\leq{}^{\mu_{1}}g(\widebar{\bm{z}})+\frac{1}{2}\frac{\mu_{1}-\mu_{2}}{\mu_{2}}\mu_{1}L_{g}^{2}.$ (24)
3 Relation between First-order Optimality Conditions of Problems 1.1 and 1.2
This section presents a relation between the original Problem 1.1 and the parameterized Problem 1.2. We begin with the notions of optimality.
Definition 3.1 (Optimality).
Let $J:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ be proper.
  1. (a)
    (Local minimizer) $\bm{x}^{\star\star}\in\mathrm{dom}(J)$ is said to be a local minimizer of $J$ if there exists an open neighborhood $\mathcal{N}_{\bm{x}^{\star\star}}\subset\mathcal{X}$ of $\bm{x}^{\star\star}$ such that $J(\bm{x}^{\star\star})\leq J(\bm{x})\ (\bm{x}\in\mathcal{N}_{\bm{x}^{\star\star}})$.
  2. (b)
    (Stationary point) $\bm{x}^{\star}\in\mathcal{X}$ is said to be a stationary point [^5] A stationary point $\bm{x}^{\star}\in\mathcal{X}$ is also called a limiting stationary point of $J$ in order to distinguish from a Fréchet stationary point where $\partial_{\rm F}J(\bm{x}^{\star})\ni\bm{0}$ holds. However, these stationarities are identical in a case where $J$ is subdifferentially regular in Definition 2.2 (b) (see [57, Cor. 8.11]). Since all functions in this paper are subdifferentially regular, we do not distinguish these stationarities in this paper. of $J$ if $\partial_{\rm L}J(\bm{x}^{\star})\ni\bm{0}$.
By Fermat’s rule [57, Thm. 10.1], for a proper function $J:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$, $\bm{x}^{\star\star}\in\mathcal{X}$ is a stationary point of $J$ if $\bm{x}^{\star\star}$ is a local minimizer of $J$. Thus, the condition $\partial_{\rm L}J(\bm{x}^{\star})\ni\bm{0}$ at $\bm{x}^{\star}\in\mathcal{X}$ seems to be the most versatile expressions of the first-order optimality condition for minimization of $J$ at $\bm{x}^{\star}$ (see, e.g., [57, p.207]). For this reason, we focus on finding such stationary points of Problem 1.1 and Problem 1.2 throughout this paper. By Lemma 2.4, the first-order optimality conditions can be expressed respectively with $F=h+g\circ\mathfrak{S}$ as:
[…]
for $\bm{x}^{\star}\in C$ to be a stationary point of Problem 1.1, and
[…]
${\rm or\ equivalently},\ d\left(\bm{0},\partial_{\rm L}(F\circ\bm{\varphi})(\bm{y}^{\star})\right)=0$ (27)
for $\bm{y}^{\star}\in\mathcal{Y}$ to be a stationary point of Problem 1.2.
Theorem 3.2 below presents a relation between (25) and (26). Based on Theorem 3.2, the conditions (25) and (26) are equivalent with $\bm{x}^{\star}=\bm{\varphi}(\bm{y}^{\star})$ under a mild condition (see Corollary 3.4). In a special case where the cost function $F$ is continuously differentiable, a similar relation between the first-order optimality conditions is found in [55, Thm. 2.4]. We note that, even if $C$ is not an embedded submanifold of $\mathcal{X}$ but is a closed Clarke regular subset of $\mathcal{X}$, the same statements in Theorem 3.2 hold.
Theorem 3.2 (Point-wise relation between the first-order optimality conditions (25) and (26)).
Consider Problems 1.1 and 1.2. Let $\bm{x}^{\star}\coloneqq\bm{\varphi}(\bm{y}^{\star})\in C$ with some $\bm{y}^{\star}\in\mathcal{Y}$. Then, the following hold:
  1. (a)
    If $\bm{x}^{\star}$ satisfies (25), then $\bm{y}^{\star}$ satisfies (26).
  2. (b)
    Assume $N_{C}(\bm{x}^{\star})\supset\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*})$. If $\bm{y}^{\star}$ satisfies (26), then $\bm{x}^{\star}$ satisfies (25) holds.
To prove Theorem 3.2, we use the following lemma.
Lemma 3.3.
Consider Problem 1.1. Let $\bm{\varphi}:\mathcal{Y}\to C$ be continuously differentiable, and let $\widebar{\bm{x}}\coloneqq\bm{\varphi}(\widebar{\bm{y}})\in C$ with some $\widebar{\bm{y}}\in\mathcal{Y}$. Then, $N_{C}(\widebar{\bm{x}})\subset\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\widebar{\bm{y}}))^{*})$.
Proof.
Recall $\widehat{N}_{\bm{\varphi}(\mathcal{Y})}(\widebar{\bm{x}})\subset\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\widebar{\bm{y}}))^{*})$ from Fact 2.5. Moreover, we have $\widehat{N}_{\bm{\varphi}(\mathcal{Y})}(\widebar{\bm{x}})\supset\widehat{N}_{C}(\widebar{\bm{x}})=N_{C}(\widebar{\bm{x}})$, where $\supset$” follows [^6] We can check the inclusion by showing $\limsup_{\bm{\varphi}(\mathcal{Y})\setminus\{\widebar{\bm{x}}\}\ni\bm{x}\to\widebar{\bm{x}}}\frac{{\left\langle\bm{v},\bm{x}-\widebar{\bm{x}}\right\rangle}}{\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert}\leq\limsup_{C\setminus\{\widebar{\bm{x}}\}\ni\bm{x}\to\widebar{\bm{x}}}\frac{{\left\langle\bm{v},\bm{x}-\widebar{\bm{x}}\right\rangle}}{\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert}$ from (15). The inequality can be verified with (9) from $\inf\limits_{\epsilon>0}\left(\sup\limits_{\begin{subarray}{c}0<\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert<\epsilon\\ \bm{x}\in\bm{\varphi}(\mathcal{Y})\end{subarray}}\frac{{\left\langle\bm{v},\bm{x}-\widebar{\bm{x}}\right\rangle}}{\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert}\right)\overset{\bm{\varphi}(\mathcal{Y})\subset C}{\leq}\inf\limits_{\epsilon>0}\left(\sup\limits_{\begin{subarray}{c}0<\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert<\epsilon\\ \bm{x}\in C\end{subarray}}\frac{{\left\langle\bm{v},\bm{x}-\widebar{\bm{x}}\right\rangle}}{\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert}\right)$. by $C\supset\bm{\varphi}(\mathcal{Y})$ (Note: $\supset$” can be replaced by “==” if $\bm{\varphi}(\mathcal{Y})=C$), and the equality follows by the Clarke regularity of $C$ (see (16) in Example 2.3). By combining these inclusions, we complete the proof. ∎
Proof of Theorem 3.2.
For simplicity, let $F\coloneqq h+g\circ\mathfrak{S}$.
(a) Let $\bm{x}^{\star}$ satisfy (25), i.e., there exists $\bm{v}\in\partial_{\rm L}F(\bm{x}^{\star})$ such that $\bm{v}\in N_{C}(\bm{x}^{\star})$. Then, Lemma 3.3 implies $\bm{v}\in N_{C}(\bm{x}^{\star})\subset\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*})$. By combining this inclusion with the linearity of $\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*})$, we have $\bm{v}\in\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*})$. Thus, $(\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*}\left[\partial_{\rm L}F\left(\bm{\varphi}(\bm{y}^{\star})\right)\right]\ni\bm{0}$, i.e., $\bm{y}^{\star}$ satisfies (26).
(b) Let $\bm{y}^{\star}\in\mathcal{Y}$ satisfy (26), i.e., there exists $\bm{v}\in\partial_{\rm L}F(\bm{\varphi}(\bm{y}^{\star}))$ such that $(\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*}[\bm{v}]=\bm{0}$. The linearity of $\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*})$ implies $\bm{v}\in\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*})$, and then the assumption $N_{C}(\bm{x}^{\star})\supset\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*})$ yields $\bm{v}\in N_{C}(\bm{\varphi}(\bm{y}^{\star}))$. Thus, we have $\bm{0}\in\partial_{\rm L}F(\bm{\varphi}(\bm{y}^{\star}))+N_{C}(\bm{\varphi}(\bm{y}^{\star}))=\partial_{\rm L}F(\bm{x}^{\star})+N_{C}(\bm{x}^{\star})$, i.e., $\bm{x}^{\star}$ satisfies (25). ∎
Corollary 3.4 (Special case for (25) and (26) to be equivalent).
[…]
Proof.
[…]
Remark 3.5 (Case for $\bm{\varphi}(\mathcal{Y})\subsetneq C$).
Even in a case where $\bm{\varphi}:\mathcal{Y}\to C$ is not surjective onto $C$, the same conclusions as in Theorem 3.2 (a) and (b), and in Corollary 3.4 can be achieved because the surjectivity of $\bm{\varphi}$ is not used in their proofs at all.
Theorem 3.2 and Corollary 3.4 encourage us to focus on finding a stationary point of Problem 1.2 for our goal of finding a stationary point of Problem 1.1. Since Problem 1.2 is an unconstrained optimization problem thanks to the available parametrization $\bm{\varphi}$, Problem 1.2 seems to be more tractable than Problem 1.1 in terms of designing iterative algorithms. Indeed, iterative algorithms with parametrization has been proposed in a case where the cost function $F=h+g\circ\mathfrak{S}$ is smooth [48, 49, 50, 51, 52, 70, 53, 54, 71, 72]. In the next section, we propose an iterative algorithm for Problem 1.2 of nonsmooth cost function $F$ and parametrization $\bm{\varphi}$.
4 Variable Smoothing Algorithm for Problem 1.2
4.1 Properties of smoothed surrogate function for Problem 1.2
In this section, we investigate properties of our smoothed surrogate function $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}\ (\mu\in(0,\eta^{-1}))$ of the cost function $(h+g\circ\mathfrak{S})\circ\bm{\varphi}$ in Problem 1.2, where $^{\mu}g$ is the Moreau envelope of $g$ in (21). The goal of this section is to present an asymptotic property, called gradient (sub-)consistency, of the gradient of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}\ (\mu\searrow 0)$ to approximate subgradients of $(h+g\circ\mathfrak{S})\circ\bm{\varphi}$ (see Theorem 4.4 and Remark 4.5). By exploiting the gradient (sub-)consistency of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$, we will establish an asymptotic convergence analysis for the proposed variable smoothing algorithm in Section 4.2. We summarize basic properties of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$, with $\mu\in(0,\eta^{-1})$ fixed, which follow from direct applications of Fact 2.7 (d).
Lemma 4.1.
For a smoothed surrogate function $F^{\langle\mu\rangle}\circ\bm{\varphi}\coloneqq(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}\ (\mu\in(0,\eta^{-1}))$ of $F\circ\bm{\varphi}=(h+g\circ\mathfrak{S})\circ\bm{\varphi}$ in Problem 1.2, we have
$(\widebar{\bm{y}}\in\mathcal{Y},\mu\in(0,\eta^{-1}))\quad$ $F^{\langle\mu\rangle}\circ\bm{\varphi}(\widebar{\bm{y}})\leq F\circ\bm{\varphi}(\widebar{\bm{y}})\leq F^{\langle\mu\rangle}\circ\bm{\varphi}(\widebar{\bm{y}})+\frac{1}{2}\mu L_{g}^{2},$ (28)
$(\mu\in(0,\eta^{-1}))\quad$ $-\infty<\inf_{\bm{y}\in\mathcal{Y}}F^{\langle\mu\rangle}\circ\bm{\varphi}(\bm{y})$ (29)
with a Lipschitz constant $L_{g}>0$ of $g$. Moreover, with $0<\mu_{2}<\mu_{1}<\eta^{-1}$, we have
$(\widebar{\bm{y}}\in\mathcal{Y})\quad F^{\langle\mu_{1}\rangle}\circ\bm{\varphi}(\widebar{\bm{y}})\leq F^{\langle\mu_{2}\rangle}\circ\bm{\varphi}(\widebar{\bm{y}})\leq F^{\langle\mu_{1}\rangle}\circ\bm{\varphi}(\widebar{\bm{y}})+L_{g}^{2}(\mu_{1}-\mu_{2}).$ (30)
Proof.
The inequality (28) follows from (23). The inequality (29) is verified by (28) and $\inf_{\bm{y}\in\mathcal{Y}}F\circ\bm{\varphi}(\bm{y})>$-$\infty$ from the condition (iv) in Problem 1.1. It suffices for deriving (30) to show
$(0<\mu_{2}<\mu_{1}<\eta^{-1},\ \widebar{\bm{z}}\in\mathcal{Z})\quad{}^{\mu_{1}}g(\widebar{\bm{z}})\leq{}^{\mu_{2}}g(\widebar{\bm{z}})\leq{}^{\mu_{1}}g(\widebar{\bm{z}})+L_{g}^{2}(\mu_{1}-\mu_{2}).$ (31)
For the case $\frac{\mu_{1}}{\mu_{2}}\leq 2$, the inequality (31) follows from (24). Now suppose $\frac{\mu_{1}}{\mu_{2}}>2$, and define $\widetilde{\mu}_{t}\coloneqq 2^{-t}\mu_{1}\ (t=0,1,2,\ldots,l-1)$ and $\widetilde{\mu}_{l}\coloneqq\mu_{2}(\geq 2^{-l}\mu_{1})$ with $l\coloneqq\lceil\log_{2}(\frac{\mu_{1}}{\mu_{2}})\rceil\geq 2$, where $\lceil\cdot\rceil$ is the ceiling function. Then, each pair $(\widetilde{\mu}_{t},\widetilde{\mu}_{t+1})$ satisfies $\frac{\widetilde{\mu}_{t}}{\widetilde{\mu}_{t+1}}\leq 2$, and thus the inequality (31) with $(\mu_{1},\mu_{2})\coloneqq(\widetilde{\mu}_{t},\widetilde{\mu}_{t+1})$ (already proved for the case $\frac{\mu_{1}}{\mu_{2}}\leq 2$) yields $0\leq{}^{\widetilde{\mu}_{t+1}}g(\widebar{\bm{z}})-{}^{\widetilde{\mu}_{t}}g(\widebar{\bm{z}})\leq L_{g}^{2}(\widetilde{\mu}_{t}-\widetilde{\mu}_{t+1})\ (\widebar{\bm{z}}\in\mathcal{Z})$. Hence, we get $0\leq{}^{\mu_{2}}g(\widebar{\bm{z}})-{}^{\mu_{1}}g(\widebar{\bm{z}})=\sum_{t=0}^{l-1}({}^{\widetilde{\mu}_{t+1}}g(\widebar{\bm{z}})-{}^{\widetilde{\mu}_{t}}g(\widebar{\bm{z}}))\leq L_{g}^{2}\sum_{t=0}^{l-1}(\widetilde{\mu}_{t}-\widetilde{\mu}_{t+1})=L_{g}^{2}(\mu_{1}-\mu_{2})$, implying the inequality (31) also follows for the case $\frac{\mu_{1}}{\mu_{2}}>2$. ∎
The Lipschitz continuity of the gradient of a cost function is one of important ingredients for optimization to ensure the descent lemma (see, e.g., [73, Lemma 5.7]). The following presents a sufficient condition for the Lipschitz constant of the gradient of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ (see also Remark 4.3 for its further sufficient conditions).
Proposition 4.2 (Lipschitzian of $\nabla((h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi})$).
Consider a smoothed surrogate function $F^{\langle\mu\rangle}\circ\bm{\varphi}\coloneqq(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}\ (\mu\in(0,\eta^{-1}))$ of $F\circ\bm{\varphi}=(h+g\circ\mathfrak{S})\circ\bm{\varphi}$ in Problem 1.2, and the following conditions:
  1. (A1)
    The operator norm of $(\mathrm{D}\mathfrak{S}(\cdot))^{*}$ is bounded above by $\kappa_{\mathfrak{S}}>0$ over the convex hull $\mathrm{Conv}(C)\coloneqq\{t\bm{x}_{1}+(1-t)\bm{x}_{2}\in\mathcal{X}\mid\bm{x}_{1},\bm{x}_{2}\in C,t\in[0,1]\}\subset\mathcal{X}$ of $C$, i.e.,
    $(\exists\kappa_{\mathfrak{S}}>0,\forall\bm{x}\in\mathrm{Conv}(C))\quad\left\lVert(\mathrm{D}\mathfrak{S}(\bm{x}))^{*}\right\rVert_{\mathrm{op}}\leq\kappa_{\mathfrak{S}};$ (32)
  2. (A2)
    $(\mathrm{D}\mathfrak{S}(\cdot))^{*}$ is Lipschitz continuous with a Lipschitz constant $L_{\mathrm{D}\mathfrak{S}}>0$ over $C$, i.e.,
    $(\exists L_{\mathrm{D}\mathfrak{S}}>0,\forall\bm{x}_{1},\bm{x}_{2}\in C)\ \left\lVert(\mathrm{D}\mathfrak{S}(\bm{x}_{1}))^{*}-(\mathrm{D}\mathfrak{S}(\bm{x}_{2}))^{*}\right\rVert_{\mathrm{op}}\leq L_{\mathrm{D}\mathfrak{S}}\left\lVert\bm{x}_{1}-\bm{x}_{2}\right\rVert;$ (33)
  3. (A3)
    The norm of $\nabla h$ is bounded above by $\kappa_{h}>0$ over $C$, i.e.,
    $(\exists\kappa_{h}>0,\forall\bm{x}\in C)\ \left\lVert\nabla h(\bm{x})\right\rVert\leq\kappa_{h};$ (34)
  4. (A4)
    The operator norm of $(\mathrm{D}\bm{\varphi}(\cdot))^{*}$ is bounded above by $\kappa_{\bm{\varphi}}>0$ over $\mathcal{Y}$;
  5. (A5)
    $(\mathrm{D}\bm{\varphi}(\cdot))^{*}$ is Lipschitz continuous with a Lipschitz constant $L_{\mathrm{D}\bm{\varphi}}>0$ over $\mathcal{Y}$.
Then, for $\mu\in(0,2^{-1}\eta^{-1}]$, $F^{\langle\mu\rangle}\circ\bm{\varphi}$ is continuously differentiable, and the following hold:
  1. (a)
    If the conditions (A1) and (A2) hold, then $\nabla F^{\langle\mu\rangle}$ is Lipschitz continuous with a Lipschitz constant $L_{\nabla F^{\langle\mu\rangle}}\coloneqq L_{\nabla h}+L_{g}L_{\mathrm{D}\mathfrak{S}}+\kappa_{\mathfrak{S}}^{2}\mu^{-1}>0$ over $C$.
  2. (b)
    If the conditions (A1)-(A5) hold, then $\nabla(F^{\langle\mu\rangle}\circ\bm{\varphi})$ is Lipschitz continuous with a Lipschitz constant $L_{\nabla(F^{\langle\mu\rangle}\circ\bm{\varphi})}\coloneqq\varpi_{1}+\varpi_{2}\mu^{-1}>0$ over $\mathcal{Y}$, where constants $\varpi_{1}\coloneqq\kappa_{\bm{\varphi}}^{2}(L_{\nabla h}+L_{g}L_{\mathrm{D}\mathfrak{S}})+L_{\mathrm{D}\bm{\varphi}}(\kappa_{h}+\kappa_{\mathfrak{S}}L_{g})$ and $\varpi_{2}\coloneqq\kappa_{\bm{\varphi}}^{2}\kappa_{\mathfrak{S}}^{2}$ do not depend on $\mu$.
Proof.
See Appendix B. ∎
Remark 4.3 (Sufficient conditions for assumptions in Proposition 4.2).
  1. (a)
    (Sufficient conditions regarding $\mathfrak{S}$) In a case where $\mathfrak{S}$ is a linear operator, the conditions (A1) and (A2) in Proposition 4.2 are automatically satisfied. In another case where $C$ is compact and $\mathfrak{S}$ is twice continuously differentiable, the conditions (A1)(A2) and (A3) are automatically satisfied.
  2. (b)
    (Sufficient conditions regarding $\bm{\varphi}$) The conditions (A4) and (A5) are automatically satisfied in a case where $C$ is a linear subspace of $\mathcal{X}$ and $\bm{\varphi}$ is a linear operator. Even for the other cases, we will present an example of $(C,\bm{\varphi})$, in Corollary 5.1 (see also Lemma D.2), achieving the conditions (A4) and (A5).
From (27), an approximation of a stationary point of Problem 1.2 can be reformulated as finding $\bm{y}^{\star}\in\mathcal{Y}$ where the value $d\left(\bm{0},\partial_{\rm L}(F\circ\bm{\varphi})(\bm{y}^{\star})\right)$ with $F=h+g\circ\mathfrak{S}$ is sufficiently small. Theorem 4.4 below presents an upper bound of this distance with the gradient of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ in (37) by showing the gradient sub-consistency of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ for $F$ (see also Remark 4.5).
Theorem 4.4 (Gradient sub-consistency of smoothed surrogate functions).
  1. (a)
    Let $\mathcal{Y}$ and $\mathcal{Z}$ be Euclidean spaces. Let $J:\mathcal{Z}\to\mathbb{R}$ be $\eta$-weakly convex, and $\mathcal{F}:\mathcal{Y}\to\mathcal{Z}$ be continuously differentiable. Then, $^{\mu}J\circ\mathcal{F}$ enjoys the gradient sub-consistency property (see Remark 4.5) for $J\circ\mathcal{F}$, i.e.,
    $(\widebar{\bm{y}}\in\mathcal{Y})\quad\mathop{\mathrm{Lim~sup}}_{\mathbb{R}_{++}\ni\mu\searrow 0,\ \mathcal{Y}\ni\bm{y}\to\widebar{\bm{y}}}\nabla({}^{\mu}J\circ\mathcal{F})(\bm{y})\subset\partial_{\rm L}(J\circ\mathcal{F})(\widebar{\bm{y}}).$ (35)
  2. (b)
    Consider a smoothed surrogate function $F^{\langle\mu\rangle}\circ\bm{\varphi}\coloneqq(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}\ (\mu\in(0,\eta^{-1}))$ of $F\circ\bm{\varphi}=(h+g\circ\mathfrak{S})\circ\bm{\varphi}$ in Problem 1.2. Then, we have
    $\mathop{\mathrm{Lim~sup}}_{\mathbb{R}_{++}\ni\mu\searrow 0,\ \mathcal{Y}\ni\bm{y}\to\widebar{\bm{y}}}\nabla(F^{\langle\mu\rangle}\circ\bm{\varphi})(\bm{y})\subset\partial_{\rm L}(F\circ\bm{\varphi})(\widebar{\bm{y}}).$ (36)
    Moreover, if a given sequence $(\bm{y}_{n})_{n=1}^{\infty}\subset\mathcal{Y}$ converges to some $\widebar{\bm{y}}\in\mathcal{Y}$, and $(\mu_{n})_{n=1}^{\infty}\searrow 0$ with $\mu_{n}\in(0,\eta^{-1})$, then
    $d(\bm{0},\partial_{\rm L}(F\circ\bm{\varphi})(\widebar{\bm{y}}))\leq\liminf_{n\to\infty}\left\lVert\nabla(F^{\langle\mu_{n}\rangle}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert.$ (37)
Proof.
See Appendix C. ∎
Remark 4.5 (Gradient (sub-)consistency).
  1. (a)
    Although the converse inclusion of (35) is not required for this paper, the equality
    $(\widebar{\bm{y}}\in\mathcal{Y})\quad\mathop{\mathrm{Lim~sup}}_{\mathbb{R}_{++}\ni\mu\searrow 0,\ \mathcal{Y}\ni\bm{y}\to\widebar{\bm{y}}}\nabla({}^{\mu}J\circ\mathcal{F})(\bm{y})=\partial_{\rm L}(J\circ\mathcal{F})(\widebar{\bm{y}}),$ (38)
    called the gradient consistency property [35] of $^{\mu}J\circ\mathcal{F}$ for $J\circ\mathcal{F}$, can also be shown (i) by checking that $^{\mu}J\circ\mathcal{F}$ epi-converges to $J\circ\mathcal{F}$ (see [57, Def. 7.1] for the definition of epi-convergence); and (ii) by applying [61, Lemma 3.4] to $^{\mu}J\circ\mathcal{F}$. Indeed, (i) can be verified by [57, Thm. 7.4(d)] and $^{\mu}J\circ\mathcal{F}\nearrow J\circ\mathcal{F}$ as $\mu\searrow 0$.
  2. (b)
    In a special case where $J$ is convex, we can check that $^{\mu}J\circ\mathcal{F}$ in Theorem 4.4 (a) has the gradient consistency property for $J\circ\mathcal{F}$ by using [61, Cor. 5.8]. However, the discussion in the proof of [61, Cor. 5.8] can not apply to our case where $J$ is just weakly convex because the proof of [61, Cor. 5.8] relies on the structure of the convexity of $J$ (see Attouch’s theorem [57, Thm. 12.35]). In order to show the gradient consistency of $^{\mu}J\circ\mathcal{F}$ even with $J$ weakly convex, we take a different approach in the proof of Theorem 4.4.
4.2 Variable Smoothing Algorithm with Convergence Analysis
The inequality (37) in Theorem 4.4 (b) implies that finding $\bm{y}^{\star}\in\mathcal{Y}$ satisfying (27) can be reduced to
$\begin{split}&{\rm find\ a\ convergent\ sequence}\ (\bm{y}_{n})_{n=1}^{\infty}\subset\mathcal{Y}\ {\rm such\ that}\\ &\quad\liminf_{n\to\infty}\left\lVert\nabla((h+{}^{\mu_{n}}g\circ\mathfrak{S})\circ\bm{\varphi})(\bm{y}_{n})\right\rVert=0\ {\rm with\ }(\mu_{n})_{n=1}^{\infty}(\subset(0,\eta^{-1}))\searrow 0.\end{split}$ (39)
In this section, for Problem 1.2 via the problem in (39), we present a variable smoothing algorithm in Algorithm 1 as an extension of the variable smoothing algorithm [23] that has been proposed for Problem 1.1 with a linear operator $\mathfrak{S}$ and $C\coloneqq\mathcal{X}$. The proposed algorithm is designed with a time-varying gradient descent update for minimization of
$F_{n}\circ\bm{\varphi}\coloneqq(h+{}^{\mu_{n}}g\circ\mathfrak{S})\circ\bm{\varphi}\ \mathrm{with}\ F_{n}:\mathcal{X}\to\mathbb{R}:\widebar{\bm{x}}\mapsto(h+{}^{\mu_{n}}g\circ\mathfrak{S})(\widebar{\bm{x}})$ (40)
at $n$th update from $\bm{y}_{n}\in\mathcal{Y}$ to $\bm{y}_{n+1}\coloneqq\bm{y}_{n}-\gamma_{n}\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\in\mathcal{Y}$ with $\mu_{n}\searrow 0$ and a stepsize $\gamma_{n}>0$. In a case where $\mathrm{prox}_{\mu_{n}g}$ has a closed-form expression, the values of $F_{n}\circ\bm{\varphi}(\widebar{\bm{y}})$ and $\nabla(F_{n}\circ\bm{\varphi})(\widebar{\bm{y}})$ are available for every $\widebar{\bm{y}}\in\mathcal{Y}$ (see (21) and Fact 2.7 (b)).
The index $\mu_{n}\in(0,(2\eta)^{-1}]$ in Algorithm 1 is designed to establish a convergence analysis of Algorithm 1 as
$(\mu_{n})_{n=1}^{\infty}\subset(0,(2\eta)^{-1}]\ \mathrm{satisfies}\ \begin{cases}&{\rm(i)}\ \lim_{n\to\infty}\mu_{n}=0,{\rm(ii)}\ \sum_{n=1}^{\infty}\mu_{n}=+\infty,\\ &{\rm(iii)}\ \mu_{n+1}\leq\mu_{n}\ (n\in\mathbb{N}),\end{cases}$ (41)
where $\eta>0$ is the level of the weak convexity of $g$ (see Problem 1.1). We have many choices of $(\mu_{n})_{n=1}^{\infty}$ enjoying (41) as shown in Example 4.6 (see Remark 4.10 for a reasonable choice of $(\mu_{n})_{n=1}^{\infty}$).
Algorithm 1 Variable smoothing for Problem 1.2 (VSmooth)
[…]
3:while stopping criteria are not satisfied do
4:  Set $F_{n}\coloneqq(h+{}^{\mu_{n}}g\circ\mathfrak{S})$
5:  Find $\gamma_{n}>0$ satisfying (48) with $c$ (see Example 4.8)
6:  $\bm{y}_{n+1}\leftarrow\bm{y}_{n}-\gamma_{n}\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})$
7:  $n\leftarrow n+1$
8:end while
[…]
Example 4.6 ($(\mu_{n})_{n=1}^{\infty}$ satisfying (41)).
  1. (a)
    For any $\alpha\geq 1$, $(\mu_{n})_{n=1}^{\infty}\coloneqq((2\eta)^{-1}n^{-1/\alpha})_{n=1}^{\infty}$ satisfies (41).
  2. (b)
    $(\mu_{n})_{n=1}^{\infty}\coloneqq\left(\frac{1}{2\eta(n+1)\log(n+1)}\right)_{n=1}^{\infty}$ satisfies (41).
Proof.
(a) Let $\alpha\geq 1$ be chosen arbitrarily. Clearly, $(\mu_{n})_{n=1}^{\infty}\subset(0,(2\eta)^{-1}]$ satisfies the conditions (i) and (iii) in (41). The condition (ii) can be checked by
$\sum_{n=1}^{K}n^{-1/\alpha}=\sum_{n=1}^{K}\int_{n}^{n+1}n^{-1/\alpha}dx\geq\int_{1}^{K+1}x^{-1/\alpha}dx=\begin{cases}\frac{(K+1)^{1-\alpha^{-1}}-1}{1-\alpha^{-1}}&(\alpha>1)\\ \log(K+1)&(\alpha=1)\end{cases}\overset{K\to\infty}{\to}+\infty.$ (42)
(b) Clearly, $(\mu_{n})_{n=1}^{\infty}\subset(0,(2\eta)^{-1}]$ satisfies the conditions (i) and (iii). The condition (ii) can be checked by
$\sum_{n=1}^{K}\frac{1}{(n+1)\log(n+1)}=\sum_{n=1}^{K}\int_{n}^{n+1}\frac{dx}{(n+1)\log(n+1)}\geq\int_{1}^{K+1}\frac{dx}{(x+1)\log(x+1)}$ (43)
$\overset{u=\log(x+1)}{=}\int_{\log(2)}^{\log(K+2)}\frac{du}{u}=\log(\log(K+2))-\log(\log(2))\overset{K\to\infty}{\to}+\infty.$ (44)
Every stepsize $\gamma_{n}>0$ in Algorithm 1 is chosen to enjoy the so-called Armijo condition (see, e.g., [74, Thm. 1.1]) with a predetermined constant $c\in(0,1)$:
$F_{n}\circ\bm{\varphi}(\bm{y}_{n}-\gamma_{n}\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n}))\leq F_{n}\circ\bm{\varphi}(\bm{y}_{n})-c\gamma_{n}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert^{2}.$ (45)
In a case where $\gamma_{n}$ satisfies the Armijo condition (45), the descent condition $F_{n}\circ\bm{\varphi}(\bm{y}_{n}-\gamma_{n}\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n}))<F_{n}\circ\bm{\varphi}(\bm{y}_{n})$ is achieved if $\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\neq\bm{0}$, and thus we update $\bm{y}_{n+1}\coloneqq\bm{y}_{n}-\gamma_{n}\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})$.
To use $\gamma_{n}$ satisfying (45) in Algorithm 1, we assume Assumption 4.7 below. Sufficient conditions for Assumption 4.7 (a) are found in Proposition 4.2 and Remark 4.3. Moreover, Example 4.8 presents typical choices of $\gamma_{n}$ achieving Assumption 4.7 (b) under Assumption 4.7 (a).
Assumption 4.7.
Consider Problem 1.2 and Alg. 1. For $F_{n}=h+{}^{\mu_{n}}g\circ\mathfrak{S}$, we assume:
  1. (a)
    (Gradient Lipschitz continuity condition)
    $\hskip-18.49988pt(\exists\varpi_{1}\in\mathbb{R}_{+},\exists\varpi_{2}\in\mathbb{R}_{++},\forall n\in\mathbb{N})$ (46)
    $\hskip-18.49988pt\nabla(F_{n}\circ\bm{\varphi})~\mathrm{is~Lipschitz~continuous}\mathrm{~with~a~Lipschitz~constant~}L_{\nabla(F_{n}\circ\bm{\varphi})}\coloneqq\varpi_{1}+\frac{\varpi_{2}}{\mu_{n}}.$ (47)
    (Note: under the condition (47), we have $L_{\nabla(F_{n+1}\circ\bm{\varphi})}\geq L_{\nabla(F_{n}\circ\bm{\varphi})}\geq\cdots\geq L_{\nabla(F_{1}\circ\bm{\varphi})}$ by $\mu_{n}\searrow 0$ [see the condition (iii) in (41)] for $(\mu_{n})_{n=1}^{\infty}$)
  2. (b)
    (Lower bound condition for stepsizes)
[…]
Example 4.8 (Stepsize achieving Assumption 4.7 (b)).
Under Assumption 4.7 (a), we have mainly two choices of $\gamma_{n}$ achieving (48) (see, e.g., [74, Thm. 1.1]):
  1. (a)
    With the Lipschitz constant $L_{\nabla(F_{n}\circ\bm{\varphi})}\in\mathbb{R}_{++}$,
    $(n\in\mathbb{N})\quad\gamma_{n}\coloneqq\beta L_{\nabla(F_{n}\circ\bm{\varphi})}^{-1}\coloneqq 2(1-c)L_{\nabla(F_{n}\circ\bm{\varphi})}^{-1}\in\mathbb{R}_{++}$ (49)
    satisfies the Armijo condition (45) with $c\in(0,1)$, and thus (48).
  2. (b)
    The so-called backtracking algorithm in Algorithm 2 returns $\gamma_{n}$ enjoying the Armijo condition (45) with $c\in(0,1)$, where $\gamma_{\rm initial}\in\mathbb{R}_{++}$ is an initial guess for $\gamma_{n}$, and $\rho\in(0,1)$. In this case, $\gamma_{n}$ is known to have the following lower bound:
    $(n\in\mathbb{N})\quad\gamma_{n}\geq\min\left\{\gamma_{\rm initial}L_{\nabla(F_{n}\circ\bm{\varphi})},2\rho(1-c)\right\}L_{\nabla(F_{n}\circ\bm{\varphi})}^{-1}\geq\beta L_{\nabla(F_{n}\circ\bm{\varphi})}^{-1}$ (50)
    with $\beta\coloneqq\min\left\{\gamma_{\rm initial}L_{\nabla(F_{1}\circ\bm{\varphi})},2\rho(1-c)\right\}\in\mathbb{R}_{++}$, where the first inequality follows by [74, Thm. 1.1], and the second inequality follows [^9] The derivation of the second inequality is inspired by [24, the equality just after (4.12)]. by $L_{\nabla(F_{n}\circ\bm{\varphi})}\geq L_{\nabla(F_{1}\circ\bm{\varphi})}$ (see the sentence just after (47)). From (50), Algorithm 2 is guaranteed to terminate in at most $\max\left\{1,\left\lceil\log_{\rho}\left(\frac{2(1-c)}{\gamma_{\rm inital}L_{\nabla(F_{n}\circ\bm{\varphi})}}\right)\right\rceil\right\}$ updates, and $\gamma_{n}$ given by Algorithm 2 satisfies (48), where $\lceil\cdot\rceil$ stands for the ceiling function. We note that any knowledge on $L_{\nabla(F_{n}\circ\bm{\varphi})}$ is not required in Algorithm 2.
Algorithm 2 Backtracking algorithm to find $\gamma_{n}$ satisfying (45)
$c\in(0,1),\rho\in(0,1),\ \gamma_{\rm initial}\in\mathbb{R}_{++}$
$\gamma_{n}\leftarrow\gamma_{\rm initial}$
while $F_{n}\circ\bm{\varphi}(\bm{y}_{n}-\gamma_{n}\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n}))>F_{n}\circ\bm{\varphi}(\bm{y}_{n})-c\gamma_{n}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert^{2}$ do
  $\gamma_{n}\leftarrow\rho\gamma_{n}$
end while
$\gamma_{n}$
We present below a convergence analysis of Algorithm 1 under Assumption 4.7.
Theorem 4.9 (Convergence analysis of Algorithm 1).
Consider Problem 1.2. Choose arbitrarily $\bm{y}_{1}\in\mathcal{Y}$, $c\in(0,1)$, and $(\mu_{n})_{n=1}^{\infty}\subset(0,(2\eta)^{-1}]$ satisfying (41). Suppose that $(\bm{y}_{n})_{n=1}^{\infty}\subset\mathcal{Y}$ is generated by Algorithm 1 under Assumption 4.7 with $F_{n}\coloneqq F^{\langle\mu_{n}\rangle}\coloneqq h+{}^{\mu_{n}}g\circ\mathfrak{S}\ (n\in\mathbb{N})$. Then, $F_{n}\circ\bm{\varphi}\ (n\in\mathbb{N})$ and $(\bm{y}_{n})_{n=1}^{\infty}$ enjoy the following:
  1. (a)
    For any pair $(\underline{k},\overline{k})\in\mathbb{N}^{2}$ such that $\underline{k}\leq\overline{k}$, we have
    $\min_{\underline{k}\leq n\leq\overline{k}}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert\leq\sqrt{\frac{\chi}{\sum_{n=\underline{k}}^{\overline{k}}\mu_{n}}}.$ (51)
    with $\chi\coloneqq\frac{(F_{1}\circ\bm{\varphi}(\bm{y}_{1})-\inf_{\bm{y}\in\mathcal{Y}}F_{1}\circ\bm{\varphi}(\bm{y})+L_{g}^{2}\mu_{1})(\varpi_{1}+\eta\varpi_{2})}{c\beta\eta}\in\mathbb{R}_{++}$, where (I) $L_{g},\eta\in\mathbb{R}_{++}$ are respectively the Lipschitz constant and the level of the weak convexity of $g$ (see the condition (iii) in Problem 1.1), and (II) $\varpi_{1}\in\mathbb{R}_{+}$ and $\varpi_{2},\beta\in\mathbb{R}_{++}$ are given in Assumption 4.7.
  2. (b)
    $\liminf_{n\to\infty}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert=0.$ (52)
  3. (c)
    Choose a subsequence $(\bm{y}_{m(l)})_{l=1}^{\infty}\subset\mathcal{Y}$ of $(\bm{y}_{n})_{n=1}^{\infty}$ such that $\lim_{l\to\infty}\left\lVert\nabla(F_{m(l)}\circ\bm{\varphi})(\bm{y}_{m(l)})\right\rVert=0$, where $m:\mathbb{N}\to\mathbb{N}$ is monotonically increasing [^10] From Theorem 4.9 (b), we can construct such a subsequence $(\bm{y}_{m(l)})_{l=1}^{\infty}$ by using, e.g., $m(l)\coloneqq\begin{cases}1&(l=1)\\ \min\{n\in\mathbb{N}\mid n>m(l-1),\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert\leq 2^{-l}\}&(l>1).\end{cases}$ (53) . Then, every cluster point $\bm{y}^{\star}\in\mathcal{Y}$ of $(\bm{y}_{m(l)})_{l=1}^{\infty}$ is a stationary point of Problem 1.2 in the sense of (26).
  4. (d)
    Consider $\mu_{n}\coloneqq\tau n^{-1/\alpha}\ (n\in\mathbb{N})$ with $\tau\in(0,(2\eta)^{-1}]$ and $\alpha>1$. Then, $(\mu_{n})_{n=1}^{\infty}$ satisfies (41) by Example 4.6 (a), and, for a given tolerance $\epsilon\in(0,1)$, the conditions $\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert<\epsilon$ and $\mu_{n}<\epsilon$ are jointly achieved within at most
    $\left\lceil\left(\chi\tau^{-1}(1-\alpha^{-1})+\tau^{\alpha-1}+2\right)^{\frac{\alpha}{\alpha-1}}\epsilon^{\min\left\{\frac{-2\alpha}{\alpha-1},-\alpha\right\}}\right\rceil\ \mathrm{iterations},$ (54)
    where $\chi\in\mathbb{R}_{++}$ is given in Theorem 4.9 (a), and $\lceil\cdot\rceil$ stands for the ceiling function (see also Remark 4.10 below for an interpretation of this result).
Proof.
The proof for (a) is inspired partially by that for [23, Thm. 4.1].
(a) Since $\gamma_{n}>0$ satisfies the Armijo condition (45) by Assumption 4.7 (b), we have
$(n\in\mathbb{N})\quad F_{n}\circ\bm{\varphi}(\bm{y}_{n+1})$ $=F_{n}\circ\bm{\varphi}(\bm{y}_{n}-\gamma_{n}\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n}))$ (55)
$\leq F_{n}\circ\bm{\varphi}(\bm{y}_{n})-c\gamma_{n}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert^{2}.$ (56)
By (30) in Lemma 4.1 with $0<\mu_{n+1}\leq\mu_{n}<\eta^{-1}$, we have
[…]
By combining (56) and (57), we obtain
[…]
and thus
$(n\in\mathbb{N})\quad F_{n+1}\circ\bm{\varphi}(\bm{y}_{n+1})\leq F_{n}\circ\bm{\varphi}(\bm{y}_{n})+L_{g}^{2}(\mu_{n}-\mu_{n+1}).$ (60)
By summing (59) up from $n=\underline{k}$ to $\overline{k}$, the following inequality holds:
$c\sum_{n=\underline{k}}^{\overline{k}}\gamma_{n}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert^{2}$ $\leq F_{\underline{k}}\circ\bm{\varphi}(\bm{y}_{\underline{k}})-F_{\overline{k}+1}\circ\bm{\varphi}(\bm{y}_{\overline{k}+1})+L_{g}^{2}(\mu_{\underline{k}}-\mu_{\overline{k}+1})$ (61)
$\leq F_{\underline{k}}\circ\bm{\varphi}(\bm{y}_{\underline{k}})-\inf_{\bm{y}\in\mathcal{Y}}F_{1}\circ\bm{\varphi}(\bm{y})+L_{g}^{2}(\mu_{\underline{k}}-\mu_{\overline{k}+1})$ (62)
$\leq F_{1}\circ\bm{\varphi}(\bm{y}_{1})-\inf_{\bm{y}\in\mathcal{Y}}F_{1}\circ\bm{\varphi}(\bm{y})+L_{g}^{2}(\mu_{1}-\mu_{\overline{k}+1})$ (63)
[…]
By Assumption 4.7, recall that (i) $\nabla(F_{n}\circ\bm{\varphi})$ is Lipschitz continuous with a Lipschitz constant $L_{\nabla(F_{n}\circ\bm{\varphi})}=\varpi_{1}+\varpi_{2}\mu_{n}^{-1}$ with some constants $\varpi_{1}\geq 0,\varpi_{2}>0$; (ii) every $\gamma_{n}$ is bounded below by $\beta L_{\nabla(F_{n}\circ\bm{\varphi})}^{-1}$ with some $\beta\in\mathbb{R}_{++}$. We have
$(n\in\mathbb{N})\quad\gamma_{n}\geq\beta L_{\nabla(F_{n}\circ\bm{\varphi})}^{-1}=\frac{\beta\mu_{n}}{\varpi_{1}\mu_{n}+\varpi_{2}}=\frac{\beta\eta\mu_{n}}{\varpi_{1}\eta\mu_{n}+\eta\varpi_{2}}\geq\frac{\beta\eta}{\varpi_{1}+\eta\varpi_{2}}\mu_{n},$ (65)
where the last inequality follows by $\eta\mu_{n}\in(0,2^{-1}]$. Then, the LHS in (64) can be further bounded below as
$c\sum_{n=\underline{k}}^{\overline{k}}\gamma_{n}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert^{2}$ $\geq\frac{c\beta\eta}{\varpi_{1}+\eta\varpi_{2}}\sum_{n=\underline{k}}^{\overline{k}}\mu_{n}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert^{2}$ (66)
$\geq\frac{c\beta\eta}{\varpi_{1}+\eta\varpi_{2}}\min_{\underline{k}\leq n\leq\overline{k}}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert^{2}\sum_{n=\underline{k}}^{\overline{k}}\mu_{n}.$ (67)
By (64) and (67), we get the desired inequality in (51).
(b) Let $a_{n}\coloneqq\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert$, and $b_{n}\coloneqq\inf\{a_{m}\mid m\geq n\}$. For every $\underline{k}\in\mathbb{N}$, the inequality (51) with $\overline{k}\to\infty$ implies $b_{\underline{k}}=0$ thanks to the nonsummablility of $(\mu_{n})_{n=1}^{\infty}$ (see the condition (ii) in (41)). This implies $\liminf_{n\to\infty}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert=\liminf_{n\to\infty}a_{n}=\lim_{n\to\infty}b_{n}=0$.
(c) Let $\bm{y}^{\star}\in\mathcal{Y}$ be a cluster point of $(\bm{y}_{m(l)})_{l=1}^{\infty}$. By (37) and Theorem 4.9 (b), we have $d(\bm{0},\partial_{\rm L}(F\circ\bm{\varphi})(\bm{y}^{\star}))=0$ with $F=h+g\circ\mathfrak{S}$, implying $\bm{y}^{\star}$ satisfies (26).
(d) By $\mu_{n}=\tau n^{-1/\alpha}$, we have $\mu_{n}<\epsilon$ for all $n\geq\underline{k}\coloneqq\lceil\tau^{\alpha}\epsilon^{-\alpha}\rceil+1$. We derive $\overline{k}\in\mathbb{N}$ below such that $\overline{k}\geq\underline{k}$ and $\min_{\underline{k}\leq n\leq\overline{k}}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert<\epsilon$. By (51), we have
$\min_{\underline{k}\leq n\leq\overline{k}}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert\leq\sqrt{\frac{\chi}{\tau\sum_{n=\underline{k}}^{\overline{k}}n^{-1/\alpha}}}\leq\sqrt{\frac{\chi(1-\alpha^{-1})}{\tau\left((\overline{k}+1)^{1-\alpha^{-1}}-\underline{k}^{1-\alpha^{-1}}\right)}},$ (68)
where we used $\sum_{n=\underline{k}}^{\overline{k}}n^{-1/\alpha}\geq\frac{(\overline{k}+1)^{1-\alpha^{-1}}-\underline{k}^{1-\alpha^{-1}}}{1-\alpha^{-1}}$ by a discussion similar to (42). Hence, $\min_{\underline{k}\leq n\leq\overline{k}}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert<\epsilon$ holds if $\overline{k}$ satisfies $\sqrt{\frac{\chi(1-\alpha^{-1})}{\tau\left((\overline{k}+1)^{1-\alpha^{-1}}-\underline{k}^{1-\alpha^{-1}}\right)}}<\epsilon$, or equivalently
$\overline{k}>\left(\chi\tau^{-1}(1-\alpha^{-1})\epsilon^{-2}+\underline{k}^{1-\alpha^{-1}}\right)^{\frac{\alpha}{\alpha-1}}-1.$ (69)
[…]
$\left(\chi\tau^{-1}(1-\alpha^{-1})\epsilon^{-2}+\underline{k}^{1-\alpha^{-1}}\right)^{\frac{\alpha}{\alpha-1}}\leq\left(\left(\chi\tau^{-1}(1-\alpha^{-1})+\tau^{\alpha-1}+2\right)\epsilon^{\min\{-2,1-\alpha\}}\right)^{\frac{\alpha}{\alpha-1}}$ (70)
$\leq\left\lceil\left(\chi\tau^{-1}(1-\alpha^{-1})+\tau^{\alpha-1}+2\right)^{\frac{\alpha}{\alpha-1}}\epsilon^{\min\left\{\frac{-2\alpha}{\alpha-1},-\alpha\right\}}\right\rceil\eqqcolon n_{0}.$ (71)
Hence, (69) is achieved by $\overline{k}\coloneqq n_{0}$, and thus we get $\min_{\underline{k}\leq n\leq n_{0}}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert<\epsilon$, i.e., there exists $\widehat{n}\in[\underline{k},n_{0}]\cap\mathbb{N}$ such that $\left\lVert\nabla(F_{\widehat{n}}\circ\bm{\varphi})(\bm{y}_{\widehat{n}})\right\rVert<\epsilon$. Since $\mu_{n}<\epsilon$ holds for all $n\geq\underline{k}$, the conditions $\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert<\epsilon$ and $\mu_{n}<\epsilon$ are jointly achieved within at most $n_{0}$ iterations. ∎
Remark 4.10 (Interpretation of Theorem 4.9 (d)).
For a given $\epsilon>0$, a point $\widehat{\bm{y}}\in\mathcal{Y}$ satisfying $\mu<\epsilon$ and $\left\lVert\nabla(F^{\langle\mu\rangle}\circ\bm{\varphi})(\widehat{\bm{y}})\right\rVert<\epsilon$ with $F^{\langle\mu\rangle}=h+{}^{\mu}g\circ\mathfrak{S}$ is called an $\epsilon$-approximate stationary point [75, 41, 39] of Problem 1.2, where $F^{\langle\mu\rangle}\circ\bm{\varphi}$ has the gradient consistency property for $F\circ\bm{\varphi}$ (see Theorem 4.4 (b) and Remark 4.5). By Theorem 4.9 (d), for a given $\epsilon\in(0,1)$, a generated sequence $(\bm{y}_{n})_{n=1}^{\infty}$ by Algorithm 1 with $\mu_{n}\coloneqq\tau n^{-1/\alpha}\ (\tau\in(0,2\eta)^{-1}],\ \alpha>1)$ achieves an $\epsilon$-approximate stationary point of Problem 1.2 within $\mathcal{O}(\epsilon^{\min\{-2\alpha/(\alpha-1),-\alpha\}})$ iterations, where $\mathcal{O}$ stands for the Landau’s Big O notation. Among $\mathcal{O}(\epsilon^{\min\{-2\alpha/(\alpha-1),-\alpha\}})$ for $\alpha>1$, the parameter $\alpha\coloneqq 3$ achieves the best iteration complexity $\mathcal{O}(\epsilon^{-3})$. The same complexity $\mathcal{O}(\epsilon^{-3})$ is shown for a smoothing method in [41] for the minimization of a nonconvex nonsmooth function $h(\bm{x})+\sum_{i=1}^{N}\psi(\lvert[\bm{x}]_{i}\rvert^{p})\ (\bm{x}\in\mathbb{R}^{N})$, where $h$ is continuously differentiable, and its gradient is Lipschitz continuous; $0<p\leq 1$; and $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is a penalty function such that $\psi$ is continuously differentiable, nondecreasing, and $\psi^{\prime}$ is locally Lipschitz continuous. We note that (i) the parameter $\alpha\coloneqq 3$ in $\mu_{n}\coloneqq\tau n^{-1/\alpha}$ has been utilized in variable smoothing-type algorithms [23, 25, 24] as a reasonable smoothing parameter; and (ii) by Theorem 4.9 (c), an asymptotic convergence property of Algorithm 1 is guaranteed with smoothing parameters $(\mu_{n})_{n=1}^{\infty}$ satisfying (41) other than $\mu_{n}\coloneqq\tau n^{-1/\alpha}$.
5 Numerical experiments
We demonstrate the efficacy of Algorithm 1 with $\gamma_{n}$ given by Algorithm 2121212In our preliminary experiments, we examined that Algorithm 1 with $\gamma_{n}$ given by Algorithm 2 achieves faster convergence than Algorithm 1 with $\gamma_{n}$ in (49). Therefore, in the numerical experiments, we demonstrate the efficacy of Algorithm 1 with $\gamma_{n}$ given by Algorithm 2. by numerical experiments in scenarios of two applications: (i) the sparse spectral clustering (SSC) and (ii) the sparse principal component analysis (SPCA). In the SSC application (Section 5.1), we present a new formulation of SSC, via Problem 1.1 with a weakly convex and smooth composition, and demonstrate how this new formulation improves estimation performance of the existing SSC algorithm. In the SPCA application (Section 5.2), we demonstrate the efficient numerical performance of the proposed Algorithm 1 by comparing its empirical results, e.g., convergence speed, with those of fairly standard Riemmannian nonsmooth optimization algorithms [24, 21, 20] (Note: the proposed Algorithm 1 and the standard algorithms [24, 21, 20] are applicable to the SPCA problem). All experiments were performed by MATLAB on MacBookPro (Apple M3, 16GB).
Both applications are reduced to nonsmooth optimization problems over an embedded submanifold, say the Stiefel manifold ${\rm St}(p,N)\coloneqq\{\bm{U}\in\mathbb{R}^{N\times p}\mid\bm{U}^{\mathsf{T}}\bm{U}=\bm{I}_{p}\}\ (p\leq N)$, of $\mathbb{R}^{N\times p}$ [30, Section 3.3.2]. In order to apply Algorithm 1 to such problems, we employ the generalized inverse Cayley transform, denoted by $\Phi_{\bm{S}}^{-1}$ in [54], as a parametrization $\bm{\varphi}$ of ${\rm St}(p,N)$:
$\bm{\varphi}_{\bm{S}}^{\rm Cay}:(\mathcal{Y}\coloneqq)Q_{N,p}\to{\rm St}(p,N)\setminus E_{N,p}(\bm{S}):\bm{V}\mapsto\bm{S}(\bm{I}-\bm{V})(\bm{I}+\bm{V})^{-1}\bm{I}_{N\times p}$ (72)
with $\bm{S}\in{\rm O}(N)\coloneqq{\rm St}(N,N)$, where
$Q_{N,p}\coloneqq\left\{\left.\begin{bmatrix}\bm{A}&-\bm{B}^{\mathsf{T}}\\ \bm{B}&\bm{0}\end{bmatrix}\in\mathbb{R}^{N\times N}\;\right|\;\begin{subarray}{c}-\bm{A}^{\mathsf{T}}=\bm{A}\in\mathbb{R}^{p\times p},\\ \bm{B}\in\mathbb{R}^{(N-p)\times p}\end{subarray}\right\}\equiv\mathbb{R}^{Np-p(p+1)/2}$ (73)
is the set of skew-symmetric matrices, $E_{N,p}(\bm{S})\coloneqq\{\bm{U}\in{\rm St}(p,N)\mid\det(\bm{I}_{p}+\bm{I}_{N\times p}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{U})=0\}$ is called the singular-point set, and $\bm{I}_{N\times p}\in\mathbb{R}^{N\times p}$ is the first $p$ column vectors of the NN-by-NN identity matrix $\bm{I}$. Although the image $\bm{\varphi}_{\bm{S}}^{\rm Cay}(Q_{N,p})={\rm St}(p,N)\setminus E_{N,p}(\bm{S})$ is a proper subset of ${\rm St}(p,N)$, ${\rm St}(p,N)\setminus E_{N,p}(\bm{S})$ is open [^13] There exists an open subset $O\subset\mathbb{R}^{N\times p}$ in the Euclidean space $\mathbb{R}^{N\times p}$ such that ${\rm St}(p,N)\setminus E_{N,p}(\bm{S})=O\cap{\rm St}(p,N)$. and dense [^14] For any $\bm{U}\in{\rm St}(p,N)$, there exists $(\bm{U}_{n})_{n=1}^{\infty}\subset{\rm St}(p,N)\setminus E_{N,p}(\bm{S})$ such that $\lim_{n\to\infty}\bm{U}_{n}=\bm{U}$. in ${\rm St}(p,N)$ [54, Thm. 2.3 (b)], implying thus almost all points in ${\rm St}(p,N)$ can be parametrized in terms of the Euclidean space $Q_{N,p}$ with $\bm{\varphi}_{\bm{S}}^{\rm Cay}$.
Corollary 5.1 below implies that $\bm{\varphi}_{\bm{S}}^{\rm Cay}$ possesses well-suited properties for its application to Algorithm 1 as a parametrization $\bm{\varphi}$. More precisely, with $C\coloneqq{\rm St}(p,N)$ and $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$, (i) Corollary 5.1 (a) implies the equivalence of the first-order optimality conditions for Problem 1.1 and Problem 1.2; (ii) Corollary 5.1 (b) and Theorem 4.9 imply that Algorithm 1 is guaranteed to approximate a stationary point of Problem 1.2.
Corollary 5.1 (Properties of $\bm{\varphi}_{\bm{S}}^{\rm Cay}$ in (72) as parametrization of ${\rm St}(p,N)$).
Consider Problem 1.1 with $C\coloneqq{\rm St}(p,N)$. Let $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}:(\mathcal{Y}\coloneqq)Q_{N,p}\to{\rm St}(p,N)\setminus E_{N,p}(\bm{S})$ in (72) with arbitrarily chosen $\bm{S}\in{\rm O}(N)$. Then, the following hold:
  1. (a)
    For $\bm{V}^{\star}\in Q_{N,p}$ and $\bm{U}^{\star}\coloneqq\bm{\varphi}(\bm{V}^{\star})=\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}^{\star})\in{\rm St}(p,N)\setminus E_{N,p}(\bm{S})$,
    $\partial_{\rm L}(h+g\circ\mathfrak{S}+\iota_{C})(\bm{U}^{\star})\ni\bm{0}\Leftrightarrow\partial_{\rm L}((h+g\circ\mathfrak{S})\circ\bm{\varphi})(\bm{V}^{\star})\ni\bm{0}.$ (74)
  2. (b)
    If $\mathfrak{S}$ is twice continuously differentiable, then Assumption 4.7 (a) with $C\coloneqq{\rm St}(p,N)$ and $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$ is satisfied.
Proof.
See Appendix D. ∎
Remark 5.2 (Singular-point issue for minimizing $F\circ\bm{\varphi}_{\bm{S}}^{\rm Cay}$).
The procedure for minimization of $F\circ\bm{\varphi}_{\bm{S}}^{\rm Cay}$ with $\bm{S}\in{\rm O}(N)$ fixed may suffer from slow convergence of updating iterates [54] in a case where a minimizer $\bm{U}^{\star}\in{\rm St}(p,N)$ of $f:\mathbb{R}^{N\times p}\to\mathbb{R}$ over ${\rm St}(p,N)$ is close to the singular-point set $E_{N,p}(\bm{S})$. To suppress such a performance degradation called a singular-point issue, we recently proposed an adaptive parametrization strategy [76, 72] that updates $\bm{S}$ adaptively so that $E_{N,p}(\bm{S})$ is located far away from $\bm{U}^{\star}$, and its efficacy has been demonstrated numerically therein. This indicates that a similar adaptive parametrization strategy may improve the convergence speed of Algorithm 1, however this improvement is beyond the scope of this paper, and thus will be addressed in future work.
5.1 Sparse Spectral Clustering (SSC)
The spectral clustering (SC) [77] is one of powerful modern clustering algorithms, where the goal of clustering is to split given data $(\bm{\xi}_{i})_{i=1}^{N}\subset\mathbb{R}^{d}$ into $K\leq N$ groups without requiring labels of data. The SC has been used not only for clustering on graphs [78] in graph signal processing, but also for single-cell RNA sequence [19], remote sensing image analysis [79], and detecting clusters in networks [80].
Algorithm 3 (Sparse) Spectral clustering
1:Data $(\bm{\xi}_{i})_{i=1}^{N}\subset\mathbb{R}^{d}$.
2:Construct an affinity matrix $\bm{W}\in\mathbb{R}^{N\times N}$ whose entries $\bm{W}]_{i,j}\geq 0$ stands for the similarity between $\bm{\xi}_{i}$ and $\bm{\xi}_{j}$.
3:Compute the normalized Laplacian $\bm{L}\coloneqq\bm{I}-\bm{D}^{-1/2}\bm{W}\bm{D}^{-1/2}\in\mathbb{R}^{N\times N}$ by regarding $\bm{W}$ as the adjacency matrix of a certain graph $\mathcal{G}$, where $\bm{D}\in\mathbb{R}^{N\times N}$ is the degree (diagonal) matrix of $\bm{W}$, i.e., $\bm{D}]_{i,i}=\sum_{j=1}^{N}[\bm{W}]_{i,j}$.
4:Compute $\bm{U}^{\star}\in{\rm St}(K,N)$ by solving (75) ((77) in case of Sparse Spectral Clustering).
5:Form $\widehat{\bm{U}}^{\star}\in\mathbb{R}^{N\times K}$ by normalizing each row $\widehat{\bm{u}}^{\star}_{i}\in\mathbb{R}^{K}$ of $\bm{U}^{\star}$ to length 11.
6: Treat each row $\widehat{\bm{u}}^{\star}_{i}\in\mathbb{R}^{K}$ of $\widehat{\bm{U}}^{\star}$ as a feature vector of $\bm{\xi}_{i}$, and cluster $(\widehat{\bm{u}}_{i})_{i=1}^{N}$ into $K$ clusters by the k-means algorithm.
To split given data $(\bm{\xi}_{i})_{i=1}^{N}$ into $K$ groups, the SC is executed as in Algorithm 3 based on the graph theory (see, e.g., [81]). The SC consists of two parts: (i) constructing an affinity matrix $\bm{W}\in\mathbb{R}^{N\times N}$ and an affinity graph $\mathcal{G}$, e.g., k-nearest neighborhood graph [82], of $(\bm{\xi}_{i})_{i=1}^{N}$; (ii) decomposing $\mathcal{G}$ into $K$ connected subgraphs based on a certain spectral behavior of the normalized Laplacian $\bm{L}\coloneqq\bm{I}-\bm{D}^{-1/2}\bm{W}\bm{D}^{-1/2}\in\mathbb{R}^{N\times N}$ of $\mathcal{G}$, where $\bm{D}\in\mathbb{R}^{N\times N}$ is the degree matrix of $\mathcal{G}$. More precisely for (ii), the SC exploits the fact that if $\bm{L}$ has an eigenvalue 0 with the multiplicity $K$, then $\mathcal{G}$ can be decomposed into $K$ connected subgraphs by clustering the eigenvectors associated with the eigenvalue 0 (see, e.g., [81]). By leveraging this fact, the SC computes the $K$ smallest eigenvectors of $L$, in the third step of Algorithm 3, which can be characterized as a minimizer of the following optimization problem:
$\mathrm{find}\ \bm{U}^{\star}\in\mathop{\mathrm{argmin}}\limits_{\bm{U}\in{\rm St}(K,N)}{\rm Tr}(\bm{U}^{\mathsf{T}}\bm{L}\bm{U}).$ (75)
In the fourth step of Algorithm 3, we normalize each row vector, say $\widehat{\bm{u}}^{\star}_{i}\in\mathbb{R}^{K}$, of $\bm{U}^{\star}$ as $\left\lVert\widehat{\bm{u}}^{\star}_{i}\right\rVert_{2}=1$. Finally, we cluster $(\widehat{\bm{u}}^{\star}_{i})_{i=1}^{N}$ into $K$ groups by using the k-means algorithm.
In order to improve the numerical performance of the SC, the sparse SC (SSC) [5, 19] exploits a prior knowledge that the affinity matrix $\bm{W}$ and Laplacian $\bm{L}$ can be block diagonal, in the ideal case where there is no edge between nodes in different groups, by using a certain permutation $p:\{1,2,\ldots,N\}\to\{1,2,\ldots,N\}$ of indices of $(\bm{\xi}_{i})_{i=1}^{N}$ as
$(\underbrace{\bm{\xi}_{p(1)},\bm{\xi}_{p(2)},\ldots,\bm{\xi}_{p(\mathcal{I}_{1})}}_{1\mathrm{st\ group}},\underbrace{\bm{\xi}_{p(\mathcal{I}_{1}+1)},\ldots,\bm{\xi}_{p(\mathcal{I}_{1}+\mathcal{I}_{2})}}_{2\mathrm{nd\ group}},\ldots,\underbrace{\bm{\xi}_{p(\sum_{k=1}^{K-1}\mathcal{I}_{k}+1)},\ldots,\bm{\xi}_{p(N)}}_{K\mathrm{th\ group}}),$ (76)
where $\mathcal{I}_{k}\in\mathbb{N}$ denotes the size of $k$th group. The block diagonality of $\bm{L}$ indicates the block diagonality of $\bm{U}^{\star}\bm{U}^{\star\mathsf{T}}\in\mathbb{R}^{N\times N}$ with $\bm{U}^{\star}\in{\rm St}(K,N)$ in (75), thereby inducing the sparsity of $\bm{U}^{\star}\bm{U}^{\star\mathsf{T}}\in\mathbb{R}^{N\times N}$. Although the above permutation is not available in general, $\bm{U}^{\star}\bm{U}^{\star\mathsf{T}}$ is expected to be sparse even if any permutation is employed. Therefore, in the SSC, the sparsity of $\bm{U}^{\star}\bm{U}^{\star\mathsf{T}}$ is promoted via solving the following problem in place of (75):
$\mathrm{find}\ \bm{U}^{\star}\in\mathop{\mathrm{argmin}}\limits_{\bm{U}\in{\rm St}(K,N)}{\rm Tr}(\bm{U}^{\mathsf{T}}\bm{L}\bm{U})+\lambda\psi\circ\mathfrak{S}(\bm{U}),$ (77)
where $\lambda>0$, $\mathfrak{S}:\mathbb{R}^{N\times K}\to\mathbb{R}^{N\times N}:\bm{U}\mapsto\bm{U}\bm{U}^{\mathsf{T}}$, and $\psi:\mathbb{R}^{N\times N}\to\mathbb{R}$ is a sparsity promoting function, e.g., $\ell_{1}$-norm, MCP, and SCAD (see, e.g., [6, 9]).
By Corollary 5.1, the problem (77) can be reformulated as Problem 1.2 with $C\coloneqq{\rm St}(K,N)$, $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$ in (72), $h(\bm{U})\coloneqq{\rm Tr}(\bm{U}^{\mathsf{T}}\bm{L}\bm{U})$ and $g\coloneqq\lambda\psi$:
$\mathrm{find}\ \bm{V}^{\star}\in\mathop{\mathrm{argmin}}\limits_{\bm{V}\in Q_{N,K}}(\underbrace{h+g\circ\mathfrak{S}}_{\eqqcolon F})\circ\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}).$ (78)
By employing a Lipschitz continuous and weakly convex function as $\psi$, we can find a stationary point of (77) by Algorithm 1 (see Theorem 4.9 with Corollary 5.1). We note that (i) standard Riemannian nonsmooth optimization algorithms, e.g., [20, 24, 25], can not be applied to the problem (77) because $\mathfrak{S}$ is nonlinear; (ii) [19] reported a prox-linear method for the problem (77) in a case where $\psi$ is the convex $\ell_{1}$-norm.
In this numerical experiment, we evaluated the numerical performance of SSC via the formulation in (78) by applying Algorithm 1 with $\gamma_{n}$ given by Algorithm 2 under the setting (for a choice of $(\mu_{n})_{n=1}^{\infty}$, see Remark 4.10)
$(\mu_{n})_{n=1}^{\infty}=((2\eta)^{-1}n^{-1/3})_{n=1}^{\infty},c=2^{-13},\rho=0.5,\gamma_{\rm initial}=\min\left\{1,\left\lVert\nabla(F_{1}\circ\bm{\varphi}_{\bm{S}}^{\rm Cay})(\bm{V}_{1})\right\rVert_{F}^{-1}\right\},$ (79)
where $\left\lVert\cdot\right\rVert_{F}$ is the Frobenius norm, $F_{1}$ is defined as in Algorithm 1, and $(\bm{S},\bm{V}_{1})\in{\rm O}(N)\times Q_{N,K}$ was chosen [^15] By following [54, Thm. 2.7], a pair $(\bm{S},\bm{V}_{1})\in{\rm O}(N)\times Q_{N,K}$ satisfying $\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}_{1})=\bm{U}_{1}$ can be chosen as $\bm{S}\coloneqq\mathrm{diag}(\bm{Q}_{1}\bm{Q}_{2}^{\mathsf{T}},\bm{I}_{N-K})$ and $\bm{V}_{1}\coloneqq\Phi_{\bm{S}}(\bm{U}_{1})$, where $\bm{Q}_{1},\bm{Q}_{2}\in{\rm O}(K)$ are obtained by a singular-value decomposition of the upper block matrix $\bm{U}_{0{\rm up}}=\bm{Q}_{1}\bm{\Sigma}\bm{Q}_{2}\in\mathbb{R}^{K\times K}$ of $\bm{U}_{1}$ with a nonnegative-valued diagonal matrix $\bm{\Sigma}\in\mathbb{R}^{K\times K}$, and $\Phi_{\bm{S}}:{\rm St}(K,N)\setminus E_{N,K}(\bm{S})\to Q_{N,K}$ is the generalized Cayley transform [54, (10)-(12)] (Note: $\Phi_{\bm{S}}\circ\bm{\varphi}_{\bm{S}}^{\rm Cay}=\mathrm{Id}$ and $\bm{\varphi}_{\bm{S}}^{\rm Cay}\circ\Phi_{\bm{S}}=\mathrm{Id}$ hold respectively on their domains). to satisfy $\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}_{1})=\bm{U}_{1}$ for an initial guess $\bm{U}_{1}\in{\rm St}(K,N)$. In order to examine how much weakly convex regularizers can enhance the performance, we compared (i) the proposed SSC (SSC+MCP) by employing MCP [16]
$(\bm{X}\in\mathbb{R}^{N\times N})\quad$ $\psi^{\rm MCP}_{\theta}(\bm{X})\coloneqq\sum_{1\leq i\leq N}\sum_{1\leq j\leq N}r^{\rm MCP}_{\theta}([\bm{X}]_{i,j})$ (80)
$\mathrm{with}\ r^{\rm MCP}_{\theta}([\bm{X}]_{i,j})\coloneqq\begin{cases}\lvert[\bm{X}]_{i,j}\rvert-\frac{[\bm{X}]_{i,j}^{2}}{2\theta}&(\mathrm{if}\ \lvert[\bm{X}]_{i,j}\rvert\leq\theta)\\ \frac{\theta}{2}&(\mathrm{otherwise})\end{cases}$ (81)
[…]
$\mathrm{find}\ \bm{P}^{\star}\in\mathop{\mathrm{argmin}}\limits_{\bm{P}\in\mathbb{R}^{N\times N},\bm{0}\preceq\bm{P}\preceq\bm{I},{\rm Tr}(\bm{P})=K}{\rm Tr}(\bm{P}^{\mathsf{T}}\bm{L})+\lambda\left\lVert\bm{P}\right\rVert_{1},$ (82)
where $\{\bm{U}\bm{U}^{\mathsf{T}}\in\mathbb{R}^{N\times N}\mid\bm{U}\in{\rm St}(K,N)\}$ is relaxed to the convex set $\{\bm{P}\in\mathbb{R}^{N\times N}\mid\bm{0}\preceq\bm{P}\preceq\bm{I},{\rm Tr}(\bm{P})=K\}$. The problem (82) can be solved by classical convex optimization algorithms, e.g., ADMM (see, e.g., [8]). In the SSC[relax] [5], $\bm{U}^{\star}$ in the third step of Algorithm 3 is obtained as first $K$ eigenvectors corresponding to the largest $K$ eigenvalues of $\bm{P}^{\star}$ in (82).
To evaluate clustering performance, we computed two standard criteria: the Normalized Mutual Information (NMI) [83] and the Adjusted Rand Index (ARI) [84]. These scores closer to one indicate better clustering performance.
[…]
Table 2: Averaged NMI and ARI of each algorithm.
iris shuttle segmentation breast cancer glass wine seeds
NMI SC 0.778 0.435 0.501 0.417 0.321 0.433 0.662
SSC[relax] 0.785 0.485 0.503 0.433 0.322 0.433 0.671
[…]
SSC+MCP 0.794 0.496 0.507 0.514 0.331 0.432 0.698
ARI SC 0.745 0.231 0.341 0.419 0.174 0.363 0.659
SSC[relax] 0.786 0.315 0.343 0.462 0.174 0.363 0.675
[…]
SSC+MCP 0.794 0.397 0.352 0.595 0.181 0.388 0.709
5.2 Sparse Principal Component Analysis (SPCA)
We consider the SPCA (see, e.g., [20, 24, 25]) formulated as
$\mathrm{find}\ \bm{U}^{\star}\in\mathop{\mathrm{argmin}}\limits_{\bm{U}\in{\rm St}(p,N)}\underbrace{-{\rm Tr}(\bm{U}^{\mathsf{T}}\bm{\Xi}^{\mathsf{T}}\bm{\Xi}\bm{U})}_{\eqqcolon h(\bm{U})}+\underbrace{\lambda\left\lVert\bm{U}\right\rVert_{1}}_{\eqqcolon g(\bm{U})}$ (83)
where $\bm{\Xi}\in\mathbb{R}^{\mathcal{I}\times N}$ is a data matrix, $\mathcal{I}\in\mathbb{N}$ is the number of data, $N$ is the dimension of each data, and $\lambda>0$ is a predetermined weight. The SPCA in (83) is a special instance of Problem 1.1 with $C\coloneqq{\rm St}(p,N)$ and $\mathfrak{S}=\mathrm{Id}$. By Corollary 5.1, the SPCA in (83) can be reformulated as Problem 1.2 with $C\coloneqq{\rm St}(p,N)$ and $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$ in (72):
$\mathrm{find}\ \bm{V}^{\star}\in\mathop{\mathrm{argmin}}\limits_{\bm{V}\in Q_{N,p}}(h+g)\circ\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V})$ (84)
In this experiment, we applied Algorithm 1 (VSmooth) to the problem in (84) under the setting (79) with $\gamma_{n}$ given by Algorithm 2.
We compared the numerical performance of VSmooth with that of three Riemannian nonsmooth optimization algorithms applied to the problem in (83): (i) the Riemannian smoothing gradient algorithm [24, Section 4.1] (RSmooth), (ii) the Riemannian subgradient method [21, (2.6)] (RSub), and (iii) the Riemannian proximal gradient method [20, Algorithm 2] (ManPGAda [^18] We used codes in. https://github.com/chenshixiang/ManPG. ). For each Riemannian optimization algorithm, the update from $\bm{U}_{n}\in{\rm St}(p,N)$ to $\bm{U}_{n+1}\in{\rm St}(p,N)$ is expressed as
$\bm{U}_{n+1}\coloneqq R_{\bm{U}_{n}}(\gamma_{n}\bm{\mathcal{D}}_{n})$ (85)
with a stepsize $\gamma_{n}>0$ and a search direction $\bm{\mathcal{D}}_{n}\in T_{{\rm St}(p,N)}(\bm{U}_{n})\coloneqq\{\bm{\mathcal{D}}\in\mathbb{R}^{N\times p}\mid\bm{U}_{n}^{\mathsf{T}}\bm{\mathcal{D}}+\bm{\mathcal{D}}^{\mathsf{T}}\bm{U}_{n}=\bm{0}\}$ in the tangent space $T_{{\rm St}(p,N)}(\bm{U}_{n})$ (see Example 2.3), where $R_{\bm{U}_{n}}:T_{{\rm St}(p,N)}(\bm{U}_{n})\to{\rm St}(p,N)$ is a retraction satisfying $R_{\bm{U}_{n}}(\bm{0})=\bm{U}_{n}$ and $\mathrm{D}R_{\bm{U}_{n}}(\bm{0})[\bm{\mathcal{D}}]=\bm{\mathcal{D}}\ (\bm{\mathcal{D}}\in T_{{\rm St}(p,N)}(\bm{U}_{n}))$ (see, e.g., [30, 31, 32]). For these algorithms, we employed the polar decomposition-based retraction [30] $R_{\bm{U}}^{\rm polar}(\bm{\mathcal{D}})\coloneqq(\bm{U}+\bm{\mathcal{D}})(\bm{I}_{p}+\bm{\mathcal{D}}^{\mathsf{T}}\bm{\mathcal{D}})^{-\frac{1}{2}}\ (\bm{U}\in{\rm St}(p,N),\bm{\mathcal{D}}\in T_{{\rm St}(p,N)}(\bm{U}))$ as the retraction $R_{\bm{U}}$ because $R_{\bm{U}}^{\rm polar}$ is the retraction that was used in [20, 21, 24]. Every search direction $\bm{\mathcal{D}}_{n}$ in (85) is given respectively by:
$\bm{\mathcal{D}}_{n}^{\rm RSmooth}$ $\coloneqq-P_{T_{{\rm St}(p,N)}(\bm{U}_{n})}(\nabla(h+{}^{\mu_{n}}g)(\bm{U}_{n}))\ \mathrm{with}\ \mu_{n}\searrow 0;$ (86)
$\bm{\mathcal{D}}_{n}^{\rm RSub}$ $\coloneqq-P_{T_{{\rm St}(p,N)}(\bm{U}_{n})}(\nabla h(\bm{U}_{n})+\bm{\mathcal{G}})\ \mathrm{with}\ \bm{\mathcal{G}}\in\partial_{\rm L}g(\bm{U}_{n});$ (87)
$\bm{\mathcal{D}}_{n}^{\rm ManPGAda}$ $\coloneqq\mathop{\mathrm{argmin}}\limits_{\bm{\mathcal{D}}\in T_{{\rm St}(p,N)}(\bm{U}_{n})}{\left\langle\nabla h(\bm{U}_{n}),\bm{\mathcal{D}}\right\rangle}+\frac{1}{2t_{n}}\left\lVert\bm{\mathcal{D}}\right\rVert_{\bm{\varphi}}^{2}+g(\bm{U}_{n}+\bm{\mathcal{D}})$ (88)
$\hskip 30.00005pt\mathrm{with}\ t_{n}>0,$ (89)
where $P_{T_{{\rm St}(p,N)}(\bm{U}_{n})}:\mathbb{R}^{N\times p}\to T_{{\rm St}(p,N)}(\bm{U}_{n})$ is the orthogonal projection mapping onto $T_{{\rm St}(p,N)}(\bm{U}_{n})$ (see, e.g., [30]), and $t_{n}>0$ for ManPGAda [20] is adaptively designed to accelerate convergence speed. Clearly, ManPGAda requires some iterative solver for the subproblem (89) to find a search direction while the others do not require, i.e., the others are single-loop algorithms. For RSmooth, we employed $(\mu_{n})_{n=1}^{\infty}=((2\eta)^{-1}n^{-1/3})_{n=1}^{\infty}$ because RSmooth with this $(\mu_{n})_{n=1}^{\infty}$ achieves the best convergence rate [24, Remark 4.1]. In this experiment, to find a stepsize $\gamma_{n}$, we used backtracking algorithms for ManPGAda and RSmooth suggested in [20] and [24] respectively, and we employed $\gamma_{n}\coloneqq 0.99^{n}$ for RSub.
For the problem (83) with each $N\in\{200,500,1000\}$, $p\in\{1,N/100,N/10\}$ and $\lambda=0.1$, we tested performance of all algorithms with $10$ random $\bm{\Xi}\in\mathbb{R}^{\mathcal{I}\times N}$ with $\mathcal{I}=5000$ generated by the following procedure: (i) generate $\bm{\Xi}\in\mathbb{R}^{\mathcal{I}\times N}$ from the standard normal distribution; (ii) shift the columns of $\bm{\Xi}$ to have zero means; (iii) normalize $\bm{\Xi}$ such that $\left\lVert\bm{\Xi}\right\rVert_{F}=1$. For each trial, all algorithms started with the same initial guess $\bm{U}_{1}\in{\rm St}(p,N)$ generated by a MATLAB code “orth(randn(N,p))”. Each algorithm was terminated at the $n$th iteration when the elapsed CPU time at the end of $n$th iteration exceeded $t(>0)$ seconds (see Table 3 for the value of $t$).
Table 3: Time limits $t$ (s) used for termination
p=1 p=100/N p=10/N
N=200 0.5 0.5 1.5
N=500 1.5 1.5 3.0
N=1000 3.0 3.0 9.0
Figure 5.2 demonstrates the convergence histories for the problem (83), where the plots show CPU time on the horizontal axis versus the averaged cost function value on the vertical axis. The detailed results are demonstrated in Table 4, where for each final estimate $\bm{U}^{\diamond}\in{\rm St}(p,N)$, ’fval’ means the averaged value $f(\bm{U}^{\diamond})$, ’feasi’ means the averaged feasibility error $\left\lVert\bm{I}_{p}-\bm{U}^{\diamond\mathsf{T}}\bm{U}^{\diamond}\right\rVert_{F}$ of $\bm{U}^{\diamond}$ from the constraint set ${\rm St}(p,N)$, ’itr’ means the averaged number of iterations, ’time’ means the CPU time (s), and ’sparsity’ means the averaged ratio of entries in $\bm{U}^{\diamond}$ such that $\lvert[\bm{U}^{\diamond}]_{i,j}\rvert<10^{-4}$.
From Figure 5.2, we see that the proposed VSmooth converges faster than RSmooth and RSub for all problem sizes. Thus, VSmooth achieves the best performance among algorithms that do not require any iterative solver for subproblems. For $p\in\{1,N/100\}$, VSmooth seems to be competitive or inferior to ManPGAda. However, for $N\in\{500,1000\}$ and large $p=N/10$, the performance of ManPGAda severely deteriorates while no such deterioration is observed for VSmooth. Moreover, from Table 4, we also observe that ManPGAda updates for only a few iterations; and CPU time of ManPGAda significantly exceeds the time limitation shown by Table 3. These observations imply that ManPGAda took a long time to find an approximated solution of a subproblem in (89) at every iteration, and thus single-loop algorithms, namely, VSmooth, RSmooth and RSub, are more computationally reliable than ManPGAda for large problem sizes. From this experimental result, VSmooth has a comparable numerical performance compared to the fairly standard Riemannian nonsmooth optimization algorithms [20, 21, 24] even for Problem 1.1 with a simple case, i.e., $g$ is convex and $\mathfrak{S}=\mathrm{Id}$, where the proposed VSmooth can be applied to Problem 1.1, via Problem 1.2, with more general setting of the nonlinearity of $\mathfrak{S}$ as seen in Section 5.1.
Figure 1: Convergence histories of all algorithms for the problem (83) regarding the value $f(\bm{U})$ at CPU time for each problem size. Markers are put at every 100 iterations.
Table 4: Averaged performance of all algorithms for the problem (83).
Problem size Algorithm fval feasi itr time sparsity
$(N,p)=(200,1)$ RSmooth 9.50229e-02 7.216e-16 3673.8 0.50 9.950e-01
RSub 9.50082e-02 0.000e+$0$ 17357 0.50 9.950e-01
ManPGAda 9.50082e-02 2.220e-17 3927.5 0.50 9.950e-01
VSmooth 9.50728e-02 2.220e-16 5493 0.50 9.950e-01
$(N,p)=(200,2)$ RSmooth 1.90027e-01 9.578e-16 2292.4 0.50 9.950e-01
RSub 1.89986e-01 9.687e-106 11810.9 0.50 9.950e-01
ManPGAda 1.89973e-01 2.116e-16 1956.2 0.50 9.950e-01
VSmooth 1.90124e-01 7.477e-16 2994.9 0.50 9.950e-01
$(N,p)=(200,20)$ RSmooth 1.90032e+$0$ 5.081e-15 1646.5 1.50 9.950e-01
RSub 1.89997e+$0$ 3.513e-15 8918.2 1.50 9.950e-01
ManPGAda 1.89987e+$0$ 6.250e-15 1377.4 1.50 9.950e-01
VSmooth 1.90149e+$0$ 3.050e-15 2552.4 1.50 9.950e-01
$(N,p)=(500,1)$ RSmooth 9.80240e-02 7.661e-16 1943 1.50 9.980e-01
RSub 9.80058e-02 0.000e+$0$ 7027.6 1.50 9.980e-01
ManPGAda 9.80058e-02 2.220e-17 1805.2 1.50 9.980e-01
VSmooth 9.80899e-02 1.554e-16 2513.7 1.50 9.980e-01
$(N,p)=(500,5)$ RSmooth 4.90112e-01 3.119e-15 1133.5 1.50 9.980e-01
RSub 4.89997e-01 9.807e-16 4875.8 1.50 9.980e-01
ManPGAda 4.90014e-01 2.370e-15 784.8 1.50 9.980e-01
VSmooth 4.90525e-01 2.025e-15 1299.5 1.50 9.980e-01
$(N,p)=(500,50)$ RSmooth 4.98856e+$0$ 1.732e-14 774.8 4.51 9.977e-01
RSub 4.90002e+$0$ 9.091e-15 2145.3 4.50 9.980e-01
ManPGAda 1.02775e+01 1.992e-14 5 5.57 8.736e-01
VSmooth 4.90486e+$0$ 6.629e-15 1150.1 4.50 9.980e-01
$(N,p)=(1000,1)$ RSmooth 9.90208e-02 8.882e-16 1076.8 3.00 9.990e-01
RSub 9.89987e-02 0.000e+$0$ 3238 3.00 9.990e-01
ManPGAda 9.89987e-02 4.441e-17 967.9 3.00 9.990e-01
VSmooth 9.91019e-02 3.775e-16 1338.7 3.00 9.990e-01
$(N,p)=(1000,10)$ RSmooth 1.02418e+$0$ 6.645e-15 570.5 3.01 9.988e-01
RSub 9.90010e-01 2.970e-14 1700.3 3.00 9.990e-01
ManPGAda 9.90017e-01 4.901e-15 328.8 3.00 9.990e-01
VSmooth 9.91335e-01 3.006e-15 612.5 3.00 9.990e-01
$(N,p)=(1000,100)$ RSmooth 1.12218e+01 3.053e-14 376.2 9.02 9.973e-01
RSub 1.02637e+01 3.397e-14 723.6 9.01 9.051e-01
ManPGAda 4.77660e+01 3.306e-14 3 18.35 3.618e-01
VSmooth 1.01238e+01 9.032e-15 502.4 9.01 9.989e-01
6 Conclusions
We addressed a manifold constrained nonsmooth optimization problem, where the cost function consists of a smooth function and the composite function of a weakly convex function and a smooth (nonlinear) mapping. For the target problem, we proposed a variable smoothing algorithm with a parametrization of a nonconvex constraint set. We also presented an asymptotic convergence analysis and a convergence rate analysis of the proposed algorithm in terms of stationary points. The numerical experiments demonstrate that (i) the proposed algorithm improves the numerical performance of sparsity-aware application by using a weakly convex regularizer; (ii) the proposed algorithm has a comparable numerical performance compared with the existing Riemannian nonsmooth optimization algorithms based on a retraction.
[…]
Acknowledgements This work was supported partially by JSPS Grants-in-Aid (19H04134, 22KJ1270, 24K23885) and by JST SICORP (JPMJSC20C6). We thank the anonymous reviewers for their thoughtful comments, which greatly improved the clarity and quality of the paper.
Appendix A Proof of Lemma 2.4
Although Lemma 2.4 can be checked by the subdifferential calculus rules [57, Thm. 10.6, Cor. 10.9], we present a short proof of Lemma 2.4 with some necessary facts in order to increase the readability for researchers in the area of applications. The subdifferential calculus rules heavily rely on the horizontal subdifferential, which gives constraint qualifications, and the subdifferential regularity of a function (see Definition 2.2 (b) and Example A.2).
Definition A.1 (Horizontal subdifferential [57, Def. 8.3 and Thm. 8.9]).
Let $J:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ be a proper lower semicontinuous function. Then, $\partial^{\infty}J(\widebar{\bm{x}})\coloneqq\{\bm{v}\in\mathcal{X}\mid(\bm{v},0)\in N_{\mathrm{epi}(J)}(\widebar{\bm{x}},J(\widebar{\bm{x}}))\}$ is called the horizontal subdifferential of $J$ at $\widebar{\bm{x}}\in\mathrm{dom}(J)$.
Example A.2 (Subdifferentially regular functions).
Subdifferentially regular functions include, e.g., continuously differentiable functions [57, Exm. 7.28], proper lower semicontinuous (weakly) convex functions [57, Exm. 7.27] [86, Prop. 4.4.15], and indicator functions of Clarke regular subsets [57, Exm. 7.28].
Fact A.3 presents useful properties, regarding $\partial^{\infty}J$, that will be used for our analyses.
Fact A.3 (Horizontal subdifferential being zero [57, Thm. 9.13]).
Let $J:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ be a proper lower semicontinuous function. Then, for $\widebar{\bm{x}}\in\mathrm{dom}(J)$, the following conditions are equivalent:
  1. (a)
    $\partial^{\infty}J(\widebar{\bm{x}})=\{\bm{0}\}$.
  2. (b)
    $J$ is locally Lipschitz continuous at $\widebar{\bm{x}}$.
  3. (c)
    $\partial_{\rm L}J$ is locally bounded at $\widebar{\bm{x}}$, i.e., there exists an open neighborhood $\mathcal{N}_{\widebar{\bm{x}}}\subset\mathcal{X}$ of $\widebar{\bm{x}}$ such that $\partial_{\rm L}J(\mathcal{N}_{\widebar{\bm{x}}})=\bigcup_{\bm{x}\in\mathcal{N}_{\widebar{\bm{x}}}}\partial_{\rm L}J(\bm{x})$ is bounded.
Moreover, if these conditions hold at $\widebar{\bm{x}}\in\mathrm{dom}(J)$, then $\partial_{\rm L}J(\widebar{\bm{x}})\neq\emptyset$.
Fact A.4 (Calculus rules for subdifferential).
Let $\mathcal{X}$ and $\mathcal{Y}$ be Euclidean spaces. Let $J_{1},J_{2}:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ be proper lower semicontinuous and subdifferentially regular functions, and $\mathcal{F}:\mathcal{Y}\to\mathcal{X}$ be a continuously differentiable mapping. Then, the following hold:
  1. (a)
    (Chain rule [57, Thm. 10.6]) If the constraint qualification $\partial^{\infty}J_{1}(\mathcal{F}(\widebar{\bm{y}}))\cap\mathrm{Ker}((\mathrm{D}\mathcal{F}(\widebar{\bm{y}}))^{*})=\{\bm{0}\}$ holds at $\widebar{\bm{y}}\in\mathrm{dom}(J_{1}\circ\mathcal{F})$, then $J_{1}\circ\mathcal{F}$ is subdifferentially regular at $\widebar{\bm{y}}$ and
    $\partial_{\rm L}(J_{1}\circ\mathcal{F})(\widebar{\bm{y}})$ $=(\mathrm{D}\mathcal{F}(\widebar{\bm{y}}))^{*}[\partial_{\rm L}J_{1}(\mathcal{F}(\widebar{\bm{y}}))];$ (90)
    $\partial^{\infty}(J_{1}\circ\mathcal{F})(\widebar{\bm{y}})$ $=(\mathrm{D}\mathcal{F}(\widebar{\bm{y}}))^{*}[\partial^{\infty}J_{1}(\mathcal{F}(\widebar{\bm{y}}))].$ (91)
  2. (b)
    (Sum rule [57, Cor. 10.9, Exe. 8.8]) If the constraint qualification $\partial^{\infty}J_{1}(\widebar{\bm{x}})\cap(-\partial^{\infty}J_{2}(\widebar{\bm{x}}))=\{\bm{0}\}$ holds at $\widebar{\bm{x}}\in\mathrm{dom}(J_{1}+J_{2})$, then $J_{1}+J_{2}$ is subdifferentially regular at $\widebar{\bm{x}}$, and
    $\partial_{\rm L}(J_{1}+J_{2})(\widebar{\bm{x}})$ $=\partial_{\rm L}J_{1}(\widebar{\bm{x}})+\partial_{\rm L}J_{2}(\widebar{\bm{x}}).$ (92)
    Moreover, if $J_{2}$ is continuously differentiable, then $J_{1}+J_{2}$ is subdifferentially regular at $\widebar{\bm{x}}\in\mathrm{dom}(J_{1})$, and
    $\partial_{\rm L}(J_{1}+J_{2})(\widebar{\bm{x}})$ $=\partial_{\rm L}J_{1}(\widebar{\bm{x}})+\nabla J_{2}(\widebar{\bm{x}});$ (93)
    $\partial^{\infty}(J_{1}+J_{2})(\widebar{\bm{x}})$ $=\partial^{\infty}J_{1}(\widebar{\bm{x}}).$ (94)
Proof of Lemma 2.4.
[…]
(Step 2) Recall that $\iota_{C}$ is subdifferentially regular (see Example A.2 and Example 2.3), and lower semicontinuous due to the closedness of $C$. Then, by letting $J_{1}\coloneqq F$, $J_{2}\coloneqq\iota_{C}$, and $\mathcal{F}\coloneqq\bm{\varphi}$, Fact A.4 yields that (i) the subdifferential regularities of $F\circ\bm{\varphi}$ and $F+\iota_{C}$; (ii) the equalities (17) and (18) from (92) and (90), where $J_{1}$ is subdifferentially regular and $\partial^{\infty}J_{1}(\widebar{\bm{x}})=\{\bm{0}\}\ (\widebar{\bm{x}}\in\mathcal{X})$ from (Step 1). ∎
Appendix B Proof of Proposition 4.2
To prove Proposition 4.2, we use the following lemma for two times.
Lemma B.1 (Lipschitz continuity of gradient of composite function).
Let $\mathcal{H},\mathcal{K}$ be Euclidean spaces. Let $\mathcal{F}:\mathcal{K}\to\mathcal{H}$ and $J:\mathcal{H}\to\mathbb{R}$ be continuously differentiable, and let $(\emptyset\neq)E\subset\mathcal{K}$ be a subset of $\mathcal{K}$. Assume that
  1. (i)
    The operator norm of $(\mathrm{D}\mathcal{F}(\cdot))^{*}$ is bounded above by $\kappa_{\mathcal{F}}>0$ over $\mathrm{Conv}(E)$, and thus $\mathcal{F}$ is Lipschitz continuous with a Lipschitz constant $\kappa_{\mathcal{F}}$ over $\mathrm{Conv}(E)$ by [57, Thm. 9.2 and Thm. 9.7].
  2. (ii)
    $(\mathrm{D}\mathcal{F}(\cdot))^{*}$ is Lipschitz continuous with a Lipschitz constant $L_{\mathrm{D}\mathcal{F}}>0$ over $E$.
  3. (iii)
    The norm of $\nabla J$ is bounded above by $\kappa_{J}>0$ over $\mathcal{F}(E)$.
  4. (iv)
    $\nabla J$ is Lipschitz continuous with a Lipschitz constant $L_{\nabla J}>0$ over $\mathcal{F}(E)$.
Then, $\nabla(J\circ\mathcal{F})$ is Lipschitz continuous with a Lipschitz constant $\kappa_{\mathcal{F}}^{2}L_{\nabla J}+\kappa_{J}L_{\mathrm{D}\mathcal{F}}$ over $E$.
Proof.
For $\bm{x}_{1},\bm{x}_{2}\in E$, we have
$\left\lVert\nabla(J\circ\mathcal{F})(\bm{x}_{1})-\nabla(J\circ\mathcal{F})(\bm{x}_{2})\right\rVert=\left\lVert(\mathrm{D}\mathcal{F}(\bm{x}_{1}))^{*}[\nabla J(\mathcal{F}(\bm{x}_{1}))]-(\mathrm{D}\mathcal{F}(\bm{x}_{2}))^{*}[\nabla J(\mathcal{F}(\bm{x}_{2}))]\right\rVert$ (95)
$\leq\left\lVert(\mathrm{D}\mathcal{F}(\bm{x}_{1}))^{*}[\nabla J(\mathcal{F}(\bm{x}_{1}))-\nabla J(\mathcal{F}(\bm{x}_{2}))]\right\rVert+\left\lVert(\mathrm{D}\mathcal{F}(\bm{x}_{1})-\mathrm{D}\mathcal{F}(\bm{x}_{2}))^{*}[\nabla J(\mathcal{F}(\bm{x}_{2}))]\right\rVert$ (96)
[…]
Proof of Proposition 4.2.
From Fact 2.7 (b), recall that $\nabla{}^{\mu}g$ is Lipschitz continuous with the Lipschitz constant $L_{\nabla{}^{\mu}g}\coloneqq\max\left\{\mu^{-1},\frac{\eta}{1-\eta\mu}\right\}=\mu^{-1}$ by $\mu\in(0,2^{-1}\eta^{-1}]$.
(a) Since $\nabla{}^{\mu}g$ is Lipschitz continuous with the Lipschitz constant $L_{\nabla{}^{\mu}g}$ and $\left\lVert\nabla{}^{\mu}g(\bm{z})\right\rVert\leq L_{g}\ (\bm{z}\in\mathcal{Z})$ by (22), Lemma B.1 with setting $(\mathcal{H},\mathcal{K},\mathcal{F},J,E)\coloneqq(\mathcal{Z},\mathcal{X},\mathfrak{S},{}^{\mu}g,C)$ ensures the Lipschitz continuity of $\nabla({}^{\mu}g\circ\mathfrak{S})$ with a Lipschitz constant $L_{\nabla({}^{\mu}g\circ\mathfrak{S})}\coloneqq\kappa_{\mathfrak{S}}^{2}L_{\nabla{}^{\mu}g}+L_{g}L_{\mathrm{D}\mathfrak{S}}=\kappa_{\mathfrak{S}}^{2}\mu^{-1}+L_{g}L_{\mathrm{D}\mathfrak{S}}$. Thus, $\nabla(h+{}^{\mu}g\circ\mathfrak{S})$ is Lipschitz continuous over $C$ with a Lipschitz constant $L_{\nabla(h+{}^{\mu}g\circ\mathfrak{S})}=L_{\nabla h}+L_{g}L_{\mathrm{D}\mathfrak{S}}+\kappa_{\mathfrak{S}}^{2}\mu^{-1}$, where $L_{\nabla h}$ is the Lipschitz constant of $\nabla h$ by the setting of Problem 1.2.
(b) To apply Lemma B.1, we derive an upper bound $\kappa_{h+{}^{\mu}g\circ\mathfrak{S}}\coloneqq\kappa_{h}+\kappa_{\mathfrak{S}}L_{g}>0$ of $\sup_{\bm{x}\in C}\left\lVert\nabla(h+{}^{\mu}g\circ\mathfrak{S})(\bm{x})\right\rVert$ as:
$\sup_{\bm{x}\in C}\left\lVert\nabla(h+{}^{\mu}g\circ\mathfrak{S})(\bm{x})\right\rVert\leq\sup_{\bm{x}\in C}\left\lVert\nabla h(\bm{x})\right\rVert+\sup_{\bm{x}\in C}\left\lVert\nabla({}^{\mu}g\circ\mathfrak{S})(\bm{x})\right\rVert\leq\kappa_{h}+\kappa_{\mathfrak{S}}L_{g},$ (100)
[…]
Appendix C Proof of Theorem 4.4
To prove Theorem 4.4, we need the following lemma.
Lemma C.1 (Subdifferential limit).
Let $\mathcal{Y}$ and $\mathcal{Z}$ be Euclidean spaces. Let $J:\mathcal{Z}\to\mathbb{R}$ be locally Lipschitz continuous at $\widebar{\bm{z}}\in\mathcal{Z}$, and $A:\mathcal{Z}\to\mathcal{Y}$ a linear operator. Suppose that (i) an arbitrarily given $(\widebar{\bm{z}}_{n})_{n=1}^{\infty}\subset\mathcal{Z}$ satisfies $\lim_{n\to\infty}\widebar{\bm{z}}_{n}=\widebar{\bm{z}}$; (ii) an arbitrarily given $(A_{n})_{n=1}^{\infty}$ with linear operators $A_{n}:\mathcal{Z}\to\mathcal{Y}\ (n\in\mathbb{N})$ satisfies $\lim_{n\to\infty}A_{n}=A$. Then,
$\mathop{\mathrm{Lim~sup}}_{n\to\infty}A_{n}(\partial_{\rm L}J(\widebar{\bm{z}}_{n}))\subset A\left(\mathop{\mathrm{Lim~sup}}_{n\to\infty}\partial_{\rm L}J(\widebar{\bm{z}}_{n})\right)\subset A(\partial_{\rm L}J(\widebar{\bm{z}})).$ (101)
Proof.
The last inclusion in (101) can be checked by
[…]
where the last equality follows by the outer semicontinuity of the limiting subdifferential [57, Prop. 8.7] with the continuity of $J$ at $\widebar{\bm{z}}$.
To show the first inclusion in (101), let $\bm{v}\in\mathop{\mathrm{Lim~sup}}_{n\to\infty}A_{n}(\partial_{\rm L}J(\widebar{\bm{z}}_{n}))$, i.e., there exists a sequence $(\bm{v}_{n})_{n=1}^{\infty}\subset\mathcal{Y}$ with $\bm{v}_{n}\in A_{n}(\partial_{\rm L}J(\widebar{\bm{z}}_{n}))$ such that its subsequence $(\bm{v}_{m(l)})_{l=1}^{\infty}$ converges to $\bm{v}$, where $m:\mathbb{N}\to\mathbb{N}$ is monotonically increasing. Then, there exists a sequence $(\bm{u}_{n})_{n=1}^{\infty}\subset\mathcal{Z}$ with $\bm{u}_{n}\in\partial_{\rm L}J(\widebar{\bm{z}}_{n})$ satisfying $\bm{v}_{n}=A_{n}(\bm{u}_{n})$. From Fact A.3, the local Lipschitz continuity of $J$ at $\widebar{\bm{z}}$ ensures the boundedness of $\bigcup_{n=n_{0}}^{\infty}\partial_{\rm L}J(\widebar{\bm{z}}_{n})$ with a sufficiently large $n_{0}\in\mathbb{N}$. This implies the boundedness of $(\bm{u}_{n})_{n=1}^{\infty}$, thereby its subsequence $(\bm{u}_{m(l)})_{l=1}^{\infty}$ converges to some $\bm{u}\in\mathcal{Z}$ (by passing to further subsequence if necessary), i.e., $\bm{u}\in\mathop{\mathrm{Lim~sup}}_{n\to\infty}\partial_{\rm L}J(\widebar{\bm{z}}_{n})$. We get $\bm{v}=\lim_{l\to\infty}A\bm{u}_{m(l)}$ by
$\left\lVert\bm{v}-A\bm{u}_{m(l)}\right\rVert$ $\leq\left\lVert\bm{v}-\bm{v}_{m(l)}\right\rVert+\left\lVert\bm{v}_{m(l)}-A_{m(l)}\bm{u}_{m(l)}\right\rVert+\left\lVert A_{m(l)}\bm{u}_{m(l)}-A\bm{u}_{m(l)}\right\rVert$ (103)
$=\left\lVert\bm{v}-\bm{v}_{m(l)}\right\rVert+\left\lVert A_{m(l)}\bm{u}_{m(l)}-A\bm{u}_{m(l)}\right\rVert\quad(\because\bm{v}_{m(l)}=A_{m(l)}\bm{u}_{m(l)})$ (104)
$\leq\left\lVert\bm{v}-\bm{v}_{m(l)}\right\rVert+\left\lVert A_{m(l)}-A\right\rVert_{\rm op}\left\lVert\bm{u}_{m(l)}\right\rVert\overset{l\to\infty}{\to}0$ (105)
$\hskip 100.00015pt(\because\bm{v}_{m(l)}\overset{l\to\infty}{\to}\bm{v},\ A_{m(l)}\overset{l\to\infty}{\to}A,\ \bm{u}_{m(l)}\overset{l\to\infty}{\to}\bm{u}).$ (106)
[…]
Proof of Theorem 4.4.
(a) Let $\bm{v}\in\mathop{\mathrm{Lim~sup}}\limits_{\mathbb{R}_{++}\ni\mu\searrow 0,\ \bm{y}\to\widebar{\bm{y}}}\nabla({}^{\mu}J\circ\mathcal{F})(\bm{y})$. Then, there exist (i) $(\bm{y}_{n})_{n=1}^{\infty}\subset\mathcal{Y}$ with $\lim_{n\to\infty}\bm{y}_{n}=\widebar{\bm{y}}$; (ii) $(\mu_{n})_{n=1}^{\infty}\subset(0,\eta^{-1})$ with $\mu_{n}\searrow 0$; and (iii) $\bm{v}_{n}\coloneqq\nabla({}^{\mu_{n}}J\circ\mathcal{F})(\bm{y}_{n})=(\mathrm{D}\mathcal{F}(\bm{y}_{n}))^{*}[\nabla{}^{\mu_{n}}J(\mathcal{F}(\bm{y}_{n}))]$ with $\lim_{n\to\infty}\bm{v}_{n}=\bm{v}$. From Fact 2.7 (c), we have $\nabla{}^{\mu_{n}}J(\mathcal{F}(\bm{y}_{n}))\in\partial_{\rm L}J(\mathrm{prox}_{\mu_{n}J}(\mathcal{F}(\bm{y}_{n})))\ (n\in\mathbb{N})$. This implies $\bm{v}_{n}\in(\mathrm{D}\mathcal{F}(\bm{y}_{n}))^{*}[\partial_{\rm L}J(\mathrm{prox}_{\mu_{n}J}(\mathcal{F}(\bm{y}_{n})))]\ (n\in\mathbb{N})$, and thus
$\bm{v}\in\mathop{\mathrm{Lim~sup}}_{n\to\infty}(\mathrm{D}\mathcal{F}(\bm{y}_{n}))^{*}[\partial_{\rm L}J(\mathrm{prox}_{\mu_{n}J}(\mathcal{F}(\bm{y}_{n})))].$ (107)
For $n\in\mathbb{N}$, let $\widebar{\bm{z}}_{n}\coloneqq\mathrm{prox}_{\mu_{n}J}(\mathcal{F}(\bm{y}_{n}))$ and $A_{n}\coloneqq(\mathrm{D}\mathcal{F}(\bm{y}_{n}))^{*}$. Then, we have $\lim_{n\to\infty}A_{n}=(\mathrm{D}\mathcal{F}(\widebar{\bm{y}}))^{*}\eqqcolon A$ from the continuous differentiability of $\mathcal{F}$. Recall that $J$ is locally Lipschitz continuous [^19] Since every finite convex function is locally Lipschitz continuous [57, Exm. 9.14], $J+\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2}$ and $\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2}$ are locally Lipschitz continuous due to the weak convexity of $J$. Then, $J=(J+\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2})+(-\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2})$ is locally Lipschitz continuous because the sum of locally Lipschitz continuous functions is locally Lipschitz continuous [57, Exe. 9.8]. , and thus $\partial_{\rm L}J(\bm{z})\neq\emptyset$ for all $\bm{z}\in\mathcal{Z}$, by Fact A.3, i.e., $\mathrm{dom}(\partial_{\rm L}J)=\mathcal{Z}$. From $\mathrm{dom}(\partial_{\rm L}J)=\mathcal{Z}$, $\lim_{n\to\infty}\mathcal{F}(\bm{y}_{n})=\mathcal{F}(\widebar{\bm{y}})$, and $\mu_{n}\searrow 0$, Fact 2.7 (a) implies $\lim_{n\to\infty}\widebar{\bm{z}}_{n}=\lim_{n\to\infty}\mathrm{prox}_{\mu_{n}J}(\mathcal{F}(\bm{y}_{n}))=P_{\overline{\mathrm{dom}(\partial_{\rm L}J)}}(\mathcal{F}(\widebar{\bm{y}}))=\mathcal{F}(\widebar{\bm{y}})\eqqcolon\widebar{\bm{z}}$. From the local Lipschitz continuity of $J$, Lemma C.1 with $\lim_{n\to\infty}\widebar{\bm{z}}_{n}=\widebar{\bm{z}}$ and $\lim_{n\to\infty}A_{n}=A$ yields
[…]
from which $\bm{v}\in\partial_{\rm L}(J\circ\mathcal{F})(\widebar{\bm{y}})$ holds by (107).
(b) Theorem 4.4 (a) with $J\coloneqq g$ and $\mathcal{F}\coloneqq\mathfrak{S}\circ\bm{\varphi}$ yields $\partial_{\rm L}(g\circ\mathfrak{S}\circ\bm{\varphi})(\widebar{\bm{y}})\supset\mathop{\mathrm{Lim~sup}}\limits_{\mathbb{R}_{++}\ni\mu\searrow 0,\ \mathcal{Y}\ni\bm{y}\to\widebar{\bm{y}}}\nabla({}^{\mu}g\circ\mathfrak{S}\circ\bm{\varphi})(\bm{y})$. The inclusion in (36) is obtained by
[…]
$\supset\nabla(h\circ\bm{\varphi})(\widebar{\bm{y}})+\mathop{\mathrm{Lim~sup}}_{\mathbb{R}_{++}\ni\mu\searrow 0,\ \mathcal{Y}\ni\bm{y}\to\widebar{\bm{y}}}\nabla({}^{\mu}g\circ\mathfrak{S}\circ\bm{\varphi})(\bm{y})=\mathop{\mathrm{Lim~sup}}_{\mathbb{R}_{++}\ni\mu\searrow 0,\ \mathcal{Y}\ni\bm{y}\to\widebar{\bm{y}}}\nabla((h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi})(\bm{y}),\hskip-60.00009pt$ (111)
where the last equality follows by the continuity [^20] Indeed, the continuity of $\nabla(h\circ\bm{\varphi})$ yields the following relation: $\bm{v}\in\nabla(h\circ\bm{\varphi})(\widebar{\bm{y}})+\mathop{\mathrm{Lim~sup}}_{\mathbb{R}_{++}\ni\mu\searrow 0,\ \mathcal{Y}\ni\bm{y}\to\widebar{\bm{y}}}\nabla({}^{\mu}g\circ\mathfrak{S}\circ\bm{\varphi})(\bm{y})$ (112) $\Leftrightarrow\exists(\bm{y}_{n})_{n=1}^{\infty}\subset\mathcal{Y},\ \exists(\mu_{n})_{n=1}^{\infty}\subset(0,\eta^{-1})\ \mathrm{such\ that}\ \widebar{\bm{y}}=\lim_{n\to\infty}\bm{y}_{n},\ \mu_{n}\searrow 0,$ (113) $\hskip 20.00003pt\mathrm{and}\ \bm{v}=\nabla(h\circ\bm{\varphi})(\widebar{\bm{y}})+\lim_{n\to\infty}\nabla({}^{\mu_{n}}g\circ\mathfrak{S}\circ\bm{\varphi})(\bm{y}_{n})=\lim_{n\to\infty}\left(\nabla(h\circ\bm{\varphi})(\bm{y}_{n})+\nabla({}^{\mu_{n}}g\circ\mathfrak{S}\circ\bm{\varphi})(\bm{y}_{n})\right)$ (114) $\Leftrightarrow\bm{v}\in\mathop{\mathrm{Lim~sup}}_{\mathbb{R}_{++}\ni\mu\searrow 0,\ \mathcal{Y}\ni\bm{y}\to\widebar{\bm{y}}}(\nabla(h\circ\bm{\varphi})(\bm{y})+\nabla({}^{\mu}g\circ\mathfrak{S}\circ\bm{\varphi})(\bm{y}))=\mathop{\mathrm{Lim~sup}}_{\mathbb{R}_{++}\ni\mu\searrow 0,\ \mathcal{Y}\ni\bm{y}\to\widebar{\bm{y}}}\nabla((h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi})(\bm{y}).$ (115) of $\nabla(h\circ\bm{\varphi})$. Moreover, we have
[…]
Then, the inequality in (37) is obtained by
[…]
$=\liminf_{n\to\infty}d(\bm{0},\nabla((h+{}^{\mu_{n}}g\circ\mathfrak{S})\circ\bm{\varphi})(\bm{y}_{n}))=\liminf_{n\to\infty}\left\lVert\nabla((h+{}^{\mu_{n}}g\circ\mathfrak{S})\circ\bm{\varphi})(\bm{y}_{n})\right\rVert,$ (118)
where the first equality follows by the fact $\liminf_{n\to\infty}d(\bm{v},E_{n})=d\left(\bm{v},\mathop{\mathrm{Lim~sup}}_{n\to\infty}E_{n}\right)$ for $\bm{v}\in\mathcal{Y}$ and a given sequence $(E_{n})_{n=1}^{\infty}$ of subsets $E_{n}\subset\mathcal{Y}$ (see [57, Exe. 4.8]). ∎
Appendix D Proof of Corollary 5.1
To prove Corollary 5.1, we need the following fact and lemma.
Fact D.1 ([54, Lemma A.4 and proof in Proposition 2.9]).
  1. (a)
    For $\bm{A}\in\mathbb{R}^{l\times m}$ and $\bm{B}\in\mathbb{R}^{m\times n}$, $\|\bm{A}\bm{B}\|_{F}\leq\|\bm{A}\|_{2}\|\bm{B}\|_{F}$ and $\|\bm{A}\bm{B}\|_{F}\leq\|\bm{A}\|_{F}\|\bm{B}\|_{2}$ hold, where $\left\lVert\cdot\right\rVert_{F}$ and $\left\lVert\cdot\right\rVert_{2}$ denote the Frobenius norm and the spectral norm.
  2. (b)
    For $\bm{V}\in Q_{N,p}$, we have $\|(\bm{I}+\bm{V})^{-1}\|_{2}\leq 1$.
  3. (c)
    For $\bm{V}_{1},\bm{V}_{2}\in Q_{N,p}$, $\left\|(\bm{I}+\bm{V}_{1})^{-1}-(\bm{I}+\bm{V}_{2})^{-1}\right\|_{F}\leq\|\bm{V}_{1}-\bm{V}_{2}\|_{F}$.
  4. (d)
    For $\bm{V},\bm{D}\in Q_{N,p}$, $\mathrm{D}\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V})[\bm{D}]=-2\bm{S}(\bm{I}+\bm{V})^{-1}\bm{D}(\bm{I}+\bm{V})^{-1}\bm{I}_{N\times p}$
Lemma D.2 (Boundedness and Lipschitz continuity of $\mathrm{D}\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V})$).
For $\bm{\varphi}_{\bm{S}}^{\rm Cay}$ in (72) with $\bm{S}\in{\rm O}(N)$, we have
$(\bm{V}\in Q_{N,p})\quad$ $\left\lVert\mathrm{D}\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V})\right\rVert_{\rm op}\leq 2;$ (119)
$(\bm{V}_{1},\bm{V}_{2}\in Q_{N,p})\quad$ $\left\lVert\mathrm{D}\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}_{1})-\mathrm{D}\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}_{2})\right\rVert_{\rm op}\leq 4\left\lVert\bm{V}_{1}-\bm{V}_{2}\right\rVert_{F}.$ (120)
Proof.
(Proof of (119)) Let $\bm{V},\bm{D}\in Q_{N,p}$ such that $\left\lVert\bm{D}\right\rVert_{F}\leq 1$. By Fact D.1 (d), we have
$\left\lVert\mathrm{D}\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V})[\bm{D}]\right\rVert_{F}=2\left\lVert\bm{S}(\bm{I}+\bm{V})^{-1}\bm{D}(\bm{I}+\bm{V})^{-1}\bm{I}_{N\times p}\right\rVert_{F}$ (121)
[…]
implying thus (119), where we used $\left\lVert\bm{S}\right\rVert_{2}=\left\lVert\bm{I}_{N\times p}\right\rVert_{2}=1$.
(Proof of (120)) Let $\bm{V}_{1},\bm{V}_{2},\bm{D}\in Q_{N,p}$ such that $\left\lVert\bm{D}\right\rVert_{F}\leq 1$. By Fact D.1 (d), we have
$\left\lVert\mathrm{D}\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}_{1})[\bm{D}]-\mathrm{D}\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}_{2})[\bm{D}]\right\rVert_{F}$ (123)
$=2\left\lVert\bm{S}\left((\bm{I}+\bm{V}_{1})^{-1}\bm{D}(\bm{I}+\bm{V}_{1})^{-1}-(\bm{I}+\bm{V}_{2})^{-1}\bm{D}(\bm{I}+\bm{V}_{2})^{-1}\right)\bm{I}_{N\times p}\right\rVert_{F}$ (124)
[…]
$\leq 2\left\lVert(\bm{I}+\bm{V}_{1})^{-1}\bm{D}\left((\bm{I}+\bm{V}_{1})^{-1}-(\bm{I}+\bm{V}_{2})^{-1}\right)\right\rVert_{F}$ (126)
$\quad+2\left\lVert\left((\bm{I}+\bm{V}_{1})^{-1}-(\bm{I}+\bm{V}_{2})^{-1}\right)\bm{D}(\bm{I}+\bm{V}_{2})^{-1}\right\rVert_{F}.\quad(\because\left\lVert\bm{S}\right\rVert_{2}=\left\lVert\bm{I}_{N\times p}\right\rVert_{2}=1)$ (127)
The first term in (127) can be evaluated further as
$\left\lVert(\bm{I}+\bm{V}_{1})^{-1}\bm{D}\left((\bm{I}+\bm{V}_{1})^{-1}-(\bm{I}+\bm{V}_{2})^{-1}\right)\right\rVert_{F}$ (128)
[…]
Since the second term in (127) can be evaluated similarly to (131), we have $\left\lVert\mathrm{D}\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}_{1})[\bm{D}]-\mathrm{D}\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}_{2})[\bm{D}]\right\rVert_{F}\leq 4\left\lVert\bm{V}_{1}-\bm{V}_{2}\right\rVert_{F}$, implying thus (120). ∎
Proof of Corollary 5.1.
(a) By noting that $\bm{\varphi}_{\bm{S}}^{\rm Cay}:Q_{N,p}\to{\rm St}(p,N)\setminus E_{N,p}(\bm{S})$ is a diffeomorphism [54, Prop. 2.2], Corollary 3.4 with $C\coloneqq{\rm St}(p,N)$ and $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$ completes the proof [^21] Although $\bm{\varphi}_{\bm{S}}^{\rm Cay}$ is not surjective onto ${\rm St}(p,N)$, the same statement in Corollary 3.4 follows (see Remark 3.5). (see also the footnote7 in Corollary 3.4).
(b) To invoke Proposition 4.2, we check below the conditions (A1)-(A5) in Proposition 4.2 with $C\coloneqq{\rm St}(p,N)$ and $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$. Since ${\rm St}(p,N)$ is compact and $\mathfrak{S}$ is twice continuously differentiable, (A1)-(A3) are automatically satisfied (see Remark 4.3 (a)). For (A4)-(A5), see Lemma D.2. Therefore, by Proposition 4.2 (b), Assumption 4.7 (a) is satisfied with $C\coloneqq{\rm St}(p,N)$ and $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$. ∎