### v4

Abstract
In this paper, we address a manifold constrained nonsmooth optimization problem involving the composition of a weakly convex function and a smooth mapping under the availability of a parametrization of the manifold. To find a stationary point of the target problem, we propose a variable smoothing-type algorithm by combining the ideas of (i) translating the constrained problem into a Euclidean optimization problem with a parametrization of the constraint set; (ii) exploiting a sequence of smoothed surrogate functions, of the cost function, given with the Moreau envelope of a weakly convex function. The proposed algorithm produces a vector sequence by the gradient descent update of a smoothed surrogate function at each iteration. In a case where the proximity operator of the weakly convex function is available, the proposed algorithm does not require any iterative solver for subproblems therein. By leveraging tools in the variational analysis, we show the so-called gradient consistency property, which is a key ingredient for smoothing-type algorithms, of the smoothed surrogate function used in this paper. Based on the gradient consistency property, we also establish an asymptotic convergence analysis for the proposed algorithm regarding a stationary point. Numerical experiments demonstrate the efficacy of the proposed algorithm.
keywords:
manifold constrained nonsmooth optimization, weakly convex composite, variable smoothing, gradient consistency, Moreau envelope, parametrization
pacs:
MSC Classification]49J52, 49K99
1 Introduction
In this paper, we consider the following nonsmooth optimization problem with a manifold constraint.
Problem 1.1.
Let $\mathcal{X}$ and $\mathcal{Z}$ be Euclidean spaces, i.e., finite-dimensional real Hilbert spaces, and $(\emptyset\neq)C\subset\mathcal{X}$ a connected closed embedded submanifold of $\mathcal{X}$ (see Example 2.3). Then,
$\mathrm{find}\ \bm{x}^{\star}\in\mathop{\mathrm{argmin}}\limits_{\bm{x}\in C}(\underbrace{h+g\circ\mathfrak{S}}_{\eqqcolon F})(\bm{x})=\mathop{\mathrm{argmin}}\limits_{\bm{x}\in\mathcal{X}}(h+g\circ\mathfrak{S}+\iota_{C})(\bm{x}),$ (1)
where $\iota_{C}$ is the indicator function of $C$ defined by
$\iota_{C}:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}:\bm{x}\mapsto\left\{\begin{array}[]{cl}0,&\mathrm{if}\ \bm{x}\in C;\\ +\infty,&\mathrm{if}\ \bm{x}\notin C,\end{array}\right$. (2)
and $h$, $\mathfrak{S}$, and $g$ satisfy the following:
  1. (i)
    $h:\mathcal{X}\to\mathbb{R}$ is continuously differentiable and its gradient $\nabla h:\mathcal{X}\to\mathcal{X}$ is Lipschitz continuous with a Lipschitz constant $L_{\nabla h}>0$ over $C$, i.e.,
    $(\bm{x}_{1},\bm{x}_{2}\in C)\quad\left\lVert\nabla h(\bm{x}_{1})-\nabla h(\bm{x}_{2})\right\rVert\leq L_{\nabla h}\left\lVert\bm{x}_{1}-\bm{x}_{2}\right\rVert;$ (3)
  2. (ii)
    $\mathfrak{S}:\mathcal{X}\to\mathcal{Z}$ is a continuously differentiable (possibly nonlinear) mapping;
  3. (iii)
    $g:\mathcal{Z}\to\mathbb{R}$ is a (possibly nonsmooth) Lipschitz continuous with a Lipschitz constant $L_{g}>0$, i.e.,
    $(\bm{z}_{1},\bm{z}_{2}\in\mathcal{Z})\quad\lvert g(\bm{z}_{1})-g(\bm{z}_{2})\rvert\leq L_{g}\left\lVert\bm{z}_{1}-\bm{z}_{2}\right\rVert,$ (4)
    and $\eta$-weakly convex function with $\eta>0$, i.e., $g+\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2}$ is convex over $\mathcal{Z}$. Moreover, $g$ is prox-friendly, i.e., the closed-form expression of the proximity operator (see (20)) of $g$ is available as a computable tool;
  4. (iv)
    $F$ is bounded below over $C$, i.e., $\inf\{F(\bm{x})\mid\bm{x}\in C\}>$-$\infty$.
Problem 1.1 has served as key models, e.g., for finding sparse solutions, in wide range of signal processing and machine learning. Such applications include, e.g., sparse Principal Component Analysis (PCA) [11], sparse variable PCA [75], robust sparse PCA [17], orthogonal dictionary learning/robust subspace recovering [24], and sparse spectral clustering [61].
One of simple cases of Problem 1.1 is a convex case where $h$ and $g$ are convex, $\mathfrak{S}$ is linear, and $C\coloneqq\mathcal{X}$. For the convex case, the so-called proximal splitting algorithms have been extensively applied as computationally reliable algorithms (see, e.g., [8, 28]). The algorithms consist of simple steps involving the gradient of $h$ and the so-called proximity operator $\mathrm{prox}_{\mu g}:\mathcal{Z}\to\mathcal{Z}:\widebar{\bm{z}}\mapsto\mathop{\mathrm{argmin}}\limits_{\bm{z}\in\mathcal{Z}}\left(g(\bm{z})+\frac{1}{2\mu}\left\lVert\bm{z}-\widebar{\bm{z}}\right\rVert^{2}\right)$ of $g$ with index $\mu>0$. Although the computation of $\mathrm{prox}_{\mu g}$ requires, in general, solving a convex optimization problem, the proximity operators for many commonly-used functions, e.g., $\ell_{1}$-norm and nuclear norm, have closed-form expressions (see, e.g., [8, 27]).
In Problem 1.1, restricting $g$ to be convex is not always sufficient [23, 71, 85, 1, 86, 52] in order to enhance estimation performance especially for sparsity-aware applications. For example, $\ell_{1}$-norm has been widely used as a best convex approximation of the naive sparsity promoting function, i.e., $\ell_{0}$-pseudonorm, which counts the number of nonzero entries of a vector. However, the use of $\ell_{1}$-norm as $g$ leads to some underestimation of the desired sparse target because $\ell_{1}$-norm is a coercive function (see, e.g., [89]), i.e., $\ell_{1}$-norm excessively penalizes large magnitude of target. To overcome the underestimation effect caused mainly by the sparse promoting convex function, the utilization of weakly convex functions has been attracting great attention [23, 85]. Recent advances in proximal splitting techniques [71, 1, 86, 52] encourage us to find a global minimizer of Problem 1.1 under the convexity of $h+g\circ\mathfrak{S}$ with a linear operator and even with a weakly convex $g$.
Table 1: Summary of existing algorithms for Problem 1.1 and the proposed algorithm, where “cvx” denotes convex function, “w-cvx” denotes weakly convex function, ”lin” denotes linear operator, $\bf C^{1}$” denotes continuously differentiable mapping, $\mathrm{Id}$ denotes the identity operator, and “man” denotes manifold. “Complexity” stands for the first-order oracle complexity for finding an $\epsilon$-approximate stationary point, where $\mathcal{O}$ stands for Landau’s big O-notation.
Assumption Convergence result
Category Method Ref. $g$ $\mathfrak{S}$ $C(\subset\mathcal{X})$ Asymptotic guarantee ? Complexity
Double-loop Proximal gradient [25] cvx lin man# No $\mathcal{O}(\epsilon^{-2})\flat$
Proximal linear [32] cvx $\bf{C^{1}}$ $\mathcal{X}$ No $\mathcal{O}(\epsilon^{-3}\log(\epsilon^{-1}))$
[57] w-cvx $\bf{C^{1}}$ $\mathcal{X}$ Yes -
[80] cvx $\bf{C^{1}}$ man# Yes $\mathcal{O}(\epsilon^{-2})\flat$
[…]
[83] cvx $\bf{C^{1}}$ man No $\mathcal{O}(\epsilon^{-3})$
Single-loop Subgradient [59] w-cvx $\mathrm{Id}$ man# No $\mathcal{O}(\epsilon^{-4})$
Variable smoothing [15] cvx lin $\mathcal{X}$ Yes $\mathcal{O}(\epsilon^{-1})$
[13] w-cvx lin $\mathcal{X}$ No $\mathcal{O}(\epsilon^{-3})$
[10] cvx lin man# Yes $\mathcal{O}(\epsilon^{-3})$
[66] w-cvx lin man# No $\mathcal{O}(\epsilon^{-3})$
Alg. 1 w-cvx $\bf{C^{1}}$ man Yes $\mathcal{O}(\epsilon^{-3})$
(this paper)
    The compactness of $C$ is assumed in [25, 80, 10, 66, 59].
  • $\flat$
    The oracle complexities in [80, 25] are derived in a case where oracle returns the exact solutions of subproblems.
  • $\dagger$}
    The augmented Lagrangian method in [90] can also deal with an additional inequality constraint.
Beyond the convex cases, nonconvex optimization problems with nonconvex constraint have a great potential for flexible formulation to align with real-world tasks. Indeed, the minimization of nonconvex $g\circ\mathfrak{S}$ with a nonlinear smooth mapping $\mathfrak{S}$ is known to have extremely wide applications including, e.g., robust phase retrieval [34], nonnegative matrix factorization [38], robust matrix recovery [20], and robust blind deconvolution [21]. In particular, manifold constrained optimization has been continuingly attracting great attention mainly by its remarkable expressive ability of the target conditions expected to be achieved by a solution (see, e.g., [3, 70, 16, 2]).
Recently, unconstrained nonsmooth optimization algorithms, e.g., [15, 13, 8, 57, 32], have been extended to manifold constrained nonsmooth optimization algorithms [25, 80, 90, 83, 59, 10, 66] (see Table 1). For Problem 1.1, these extended algorithms can be categorized as (i) double-loop algorithms that require an iterative solver to solve a certain subproblem at each iteration (e.g., a proximal gradient method [25], a proximal linear method [80] and augmented Lagrangian methods [90, 83]); (ii) single-loop algorithms without requiring such an iterative solver at each iteration (e.g., a subgradient method [59] and variable smoothing algorithms [10, 66]). Double-loop algorithms [80, 90, 83] can deal with a nonlinear continuously differentiable mapping $\mathfrak{S}$ but with a convex function $g$ in Problem 1.1. An asymptotic convergence guarantee to a stationary point has been shown for [80, 90], while the oracle complexity i.e., the number of oracle calls, for finding an $\epsilon$-approximate stationary point has been shown for [80, 83] (and for [25] under the linear case of $\mathfrak{S}$). However, each update of double-loop algorithms heavily relies on iterative solvers for solving subproblems. Indeed, convergence analyses of double-loop algorithms have been established typically with an exact (or approximate) solution to the subproblem, although such an exact solution is not available in general within a finite number of iterations of an iterative solver. Moreover, an iterative solver for subproblems sometimes takes a lot of time to obtain even an approximate solution to a subproblem (as will be shown in numerical experiments in Section 5.2).
Single-loop algorithms have also been attracting great attention, e.g., [59, 10, 66], for Problem 1.1 because such algorithms do not require solving any subproblem at every iteration. Under the linearity of $\mathfrak{S}$, the oracle complexity for finding an $\epsilon$-approximate stationary point has been shown for [59, 10, 66], where an asymptotic convergence guarantee to a stationary point has been shown for [10] under additional assumptions such as the convexity of $g$ and the compactness of $C$. However, to the best of the authors’ knowledge, for Problem 1.1 with a nonlinear continuously differentiable mapping $\mathfrak{S}$ and a weakly convex function $g$, no single-loop algorithm has been established with a convergence guarantee, in particular an asymptotic convergence guarantee. These situations raise the following research question:
For Problem 1.1 with a manifold constraint $C$, can we design a single-loop algorithm that (i) admits a nonlinear continuously differentiable mapping $\mathfrak{S}$ in the weakly convex composite term $g\circ\mathfrak{S}$, and (ii) converges asymptotically to a stationary point of Problem 1.1 ?
In this paper, we answer this question affirmatively under the availability of a parametrization of $C$ (see Assumption 1.2 below). This assumption is satisfied by a broad class of manifolds encountered in applications (see Example 1.3).
Assumption 1.2 (Parametrization).
The manifold $C$ in Problem 1.1 can be parameterized by a continuously differentiable mapping $\bm{\varphi}:\mathcal{Y}\to C(\subset\mathcal{X})$ defined over a Euclidean space $\mathcal{Y}$, i.e., $\bm{\varphi}(\mathcal{Y})\coloneqq\{\bm{\varphi}(\bm{y})\in\mathcal{X}\mid\bm{y}\in\mathcal{Y}\}=C$. We call $\bm{\varphi}$ a parametrization of $C$ (see Example 1.3 below for pairs of $C$ and $\bm{\varphi}$).
Example 1.3 (On parametrization; see Section 5 for other examples).
  1. (a)
    Let $C$ be a linear subspace of $\mathcal{X}$. Consider a linear operator $L:\mathcal{Y}\to\mathcal{X}$ satisfying $L(\mathcal{Y})\coloneqq\{L\bm{y}\in\mathcal{X}\mid\bm{y}\in\mathcal{Y}\}=C$. Then, $\bm{\varphi}\coloneqq L$ is a parametrization of $C$. In particular, $\bm{\varphi}\coloneqq\mathrm{Id}$ is a parametrization of $C$ if $C\coloneqq\mathcal{X}$.
  2. (b)
    For $\mathcal{X}\coloneqq\mathbb{R}^{2}$, let $C\coloneqq\{\bm{x}\in\mathbb{R}^{2}\mid\left\lVert\bm{x}\right\rVert=1\}$ be the unit circle. Then, $\bm{\varphi}:\mathbb{R}\to\mathbb{R}^{2}:\theta\mapsto[\cos(\theta)\ \sin(\theta)]^{\mathsf{T}}$ is a parametrization of $C$.
  3. (c)
    Consider $C\subset\mathcal{X}$ in Problem 1.1, i.e., $C$ is an embedded submanifold of $\mathcal{X}$, endowed with the Riemannian metric induced from $\mathcal{X}$, such that $C$ is connected and closed in $\mathcal{X}$. Let $T_{C}(\bm{x})\subset\mathcal{X}$ be a tangent (vector) space to $C$ at $\bm{x}\in C$ (see (16)). In this case, for any $\bm{x}\in C$, $\mathrm{Exp}_{\bm{x}}:T_{C}(\bm{x})\to C$ with the exponential map [^1] By [16, Defs. 10.6 and 10.7, Thm. 10.8, Exe. 10.12] and [55, Thm. 4.27], the exponential map is uniquely defined over the tangent bundle $TC\coloneqq\{(\bm{x},\bm{v})\mid\bm{x}\in C,\bm{v}\in T_{C}(\bm{x})\}$ as $\mathrm{Exp}:TC\to C:(\bm{x},\bm{v})\mapsto\mathrm{Exp}_{\bm{x}}(\bm{v})\coloneqq\gamma_{\bm{v}}(1)$ with the geodesics $\gamma_{\bm{v}}:I(\subset\mathbb{R})\to C$ (see [16, Def. 5.38][55, (4.16)]) such that $\gamma_{\bm{v}}(0)=\bm{x}$, $\gamma_{\bm{v}}^{\prime}(0)=\bm{v}$ and the geodesic equation $\gamma_{\bm{v}}^{\prime\prime}(t)=0$ is satisfied for all $t\in I$, where $I\subset\mathbb{R}$ is an interval containing $[0,1]$. $\mathrm{Exp}$ is continuously differentiable by [55, Prop. 5.19], and surjective onto $C$ by [55, Cor. 6.21] together with the completeness (closedness in $\mathcal{X}$) of $C$. Hence, $\mathrm{Exp}_{\bm{x}}$ is a parametrization of $C$. See [55, p.136-p.142] for examples of the exponential map of some manifolds.
With a parametrization $\bm{\varphi}$ of $C$, Problem 1.1 can be translated into the following unconstrained optimization Problem 1.4.
Problem 1.4.
Consider $F=h+g\circ\mathfrak{S}$ and $C\subset\mathcal{X}$ in Problem 1.1 under Assumption 1.2. Then, with a parametrization $\bm{\varphi}$ in Assumption 1.2,
$\mathrm{find}\ \bm{y}^{\star}\in\mathop{\mathrm{argmin}}\limits_{\bm{y}\in\mathcal{Y}}(h+g\circ\mathfrak{S})\circ\bm{\varphi}(\bm{y})(=F\circ\bm{\varphi}(\bm{y})).$ (5)
In Section 3, we discuss relations between Problem 1.1 and Problem 1.4. More precisely, we present sufficient conditions, e.g., $N_{C}(\bm{x}^{\star})\supset\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*})$ (see (14) for $N_{C}$), or the submersion condition of $\bm{\varphi}$, for the equivalence between Problem 1.1 and Problem 1.4 in view of stationary points (see Theorem 3.2 and Corollary 3.4).
For finding a stationary point of Problem 1.4, we propose a variable smoothing algorithm as an extension of [13] proposed originally for Problem 1.1 in a special case where $C\coloneqq\mathcal{X}$ and $\mathfrak{S}$ is a linear operator (equivalently, Problem 1.4 in a special case where $\bm{\varphi}\coloneqq\mathrm{Id}$ and $\mathfrak{S}$ is a linear operator). We note that the existing variable smoothing algorithm [13] cannot be applied to Problem 1.4 due to the nonlinearity of $\mathfrak{S}\circ\bm{\varphi}$. The proposed algorithm performs a gradient descent update of a smoothed surrogate function $F^{\langle\mu_{n}\rangle}\circ\bm{\varphi}\coloneqq(h+{}^{\mu_{n}}g\circ\mathfrak{S})\circ\bm{\varphi}\ (\mu_{n}\searrow 0)$, of $F\circ\bm{\varphi}$, designed partially with the Moreau envelope $^{\mu}g$ of $g$ (see (21)), where $F^{\langle\mu\rangle}\circ\bm{\varphi}$ is continuously differentiable with $\lim_{\mu\searrow 0}F^{\langle\mu\rangle}\circ\bm{\varphi}(\bm{y})=F\circ\bm{\varphi}(\bm{y})\ (\bm{y}\in\mathcal{Y})$ (see also Lemma 4.1). In a case where $\mathrm{prox}_{\mu g}$ is computable, the proposed algorithm is a single-loop algorithm. To establish an asymptotic convergence analysis of the proposed algorithm, we show the so-called gradient consistency property of $F^{\langle\mu\rangle}\circ\bm{\varphi}$ for $F\circ\bm{\varphi}$ in Theorem 4.4, which makes it possible to express the stationarity condition of Problem 1.4 in terms of a gradient sequence $F^{\langle\mu\rangle}\circ\bm{\varphi}$ with $\mu\searrow 0$. By exploiting the gradient consistency property, we present an asymptotic convergence analysis of the proposed algorithm (see Theorem 4.9). Finally, numerical experiments, in Section 5, demonstrate the efficacy of the proposed algorithm in two scenarios of the sparse spectral clustering and the sparse principal component analysis.
A preliminary short version of this paper was partially presented in [51].
Contributions and Novelties of This Paper
  1. (I)
    A framework of a parametrization strategy for Problem 1.1 via Problem 1.4 is presented in Section 3 where relations between stationarities of Problems 1.1 and Problem 1.4 are presented in Theorem 3.2 and Corollary 3.4. Parametrization strategies have been originally developed for smooth optimization problems (see, e.g., [84, 36, 43, 41, 58, 47, 48, 56]). In contrast, our framework enables us to apply a parametrization strategy even to nonsmooth optimization problems.
  2. (II)
    The gradient consistency property of our smoothed surrogate function $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ for $F\circ\bm{\varphi}$ in Problem 1.4 is presented in Section 4.1 (see Theorem 4.4 and Remark 4.5 (a)). The gradient consistency property plays a key role in establishing an asymptotic convergence property of smoothing-type algorithms [26] (see also Appendix A.1). In a simple case where $\bm{\varphi}\coloneqq\mathrm{Id}$ and $\mathfrak{S}$ is a linear operator, our smoothed surrogate function $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ has been used in variable smoothing algorithms [15, 13, 10, 66]; nevertheless, the gradient consistency property of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ has not been reported yet to the best of the authors’ knowledge. Hence, Theorem 4.4 bridges smoothing-type algorithms, e.g., [26], and variable smoothing algorithms [15, 13, 10, 66].
  3. (III)
    In Section 4.2, we propose a single-loop variable smoothing algorithm for Problem 1.1 via Problem 1.4. For the proposed algorithm, we present an asymptotic convergence guarantee to a stationary point $\bm{y}^{\star}\in\mathcal{Y}$ of Problem 1.4 (Theorem 4.9) by leveraging the gradient consistency property in Theorem 4.4. Moreover, $\bm{\varphi}(\bm{y}^{\star})\in C$ is a stationary point of Problem 1.1 under a certain condition of $\bm{\varphi}$ (see Theorem 3.2 and Corollary 3.4). Even without this condition on $\bm{\varphi}$, $\bm{\varphi}(\bm{y}^{\star})$ satisfies a necessary condition for the local optimality of Problem 1.1 (see Remark 3.6). To the best of the authors’ knowledge, for Problem 1.1 with a nonlinear continuously differentiable mapping $\mathfrak{S}$ and a weakly convex function $g$, the proposed algorithm is the first single-loop algorithm with an asymptotic convergence guarantee to a stationary point of Problem 1.4, and hence to a stationary point of Problem 1.1 under the above condition on $\bm{\varphi}$. In particular, the proposed algorithm reproduces the variable smoothing algorithm [13] in the unconstrained case $C=\mathcal{X}$ (equivalently, $\bm{\varphi}=\mathrm{Id}$ in Problem 1.4) and linear $\mathfrak{S}$, while the proposed algorithm with an asymptotic convergence guarantee covers the case of nonlinear $\mathfrak{S}$ (Note: such an asymptotic convergence guarantee has not been established in [13]; see also Table 1).
  4. (IV)
    As a supplementary result, Corollary 4.12 establishes a first-order oracle complexity $\mathcal{O}(\epsilon^{-3})$ of the proposed algorithm for finding an $\epsilon$-approximate stationary point of Problem 1.4. We note that our approximate stationarity notion in Definition 4.10 (b) is not directly comparable with the notions used in, e.g., [66, 10, 83] and [25, 80], for Problem 1.1 because the underlying stationarity notions differ. Instead, our notion can be viewed as a generalization of the $\epsilon$-approximate stationarity used in the variable smoothing algorithm [13] (see Remark 4.11). More precisely, in a special case $\bm{\varphi}=\mathrm{Id}$ and linear $\mathfrak{S}$, Definition 4.10 (b) reduces to the approximate stationarity notion in [13], and hence our analysis in Corollary 4.12 recovers the oracle complexity $\mathcal{O}(\epsilon^{-3})$ of [13] under that special case.
Notation $\mathbb{N}$, $\mathbb{R}$, $\mathbb{R}_{+}$, and $\mathbb{R}_{++}$ denote respectively the sets of all positive integers, all real numbers, all nonnegative real numbers, and all positive real numbers. ${\left\langle\cdot,\cdot\right\rangle}_{\mathcal{X}}$ and $\left\lVert\cdot\right\rVert_{\mathcal{X}}$ stand respectively for the standard inner product and its induced norm defined in the Euclidean space $\mathcal{X}$, i.e., $\left\lVert\bm{x}\right\rVert_{\mathcal{X}}=\sqrt{{\left\langle\bm{x},\bm{x}\right\rangle}_{\mathcal{X}}}\ (\bm{x}\in\mathcal{X})$. We simply use $\left\langle\cdot,\cdot\right\rangle$} and $\left\lVert\cdot\right\rVert$ without any subscript if their domains are clear. The symbol $\mathrm{Id}$ stands for the identity operator. For a linear operator $A:\mathcal{X}\to\mathcal{Y}$, its adjoint operator $A^{*}:\mathcal{Y}\to\mathcal{X}$ is defined as the linear operator satisfying ${\left\langle\bm{x},A^{*}\bm{y}\right\rangle}_{\mathcal{X}}={\left\langle A\bm{x},\bm{y}\right\rangle}_{\mathcal{Y}}\ (\forall\bm{x}\in\mathcal{X},\bm{y}\in\mathcal{Y})$, $\mathrm{Ker}(A)\coloneqq\{\bm{x}\in\mathcal{X}\mid A\bm{x}=\bm{0}\}$ and $\mathrm{Ran}(A)\coloneqq\{A\bm{x}\in\mathcal{Y}\mid\bm{x}\in\mathcal{X}\}$ are respectively the kernel (null) space and the range space of $A$, and $\left\lVert A\right\rVert_{\rm op}\coloneqq\sup_{\left\lVert\bm{x}\right\rVert_{\mathcal{X}}\leq 1,\bm{x}\in\mathcal{X}}\left\lVert A\bm{x}\right\rVert_{\mathcal{Y}}$ denotes the operator norm of $A$. The symbol $(\cdot)^{\perp}$ denotes the orthogonal complement subspace of a given set. For a matrix $\bm{X}\in\mathbb{R}^{m\times n}$, $\bm{X}]_{i,j}$ denotes the $(i,j)$ entry of $\bm{X}$, and $\bm{X}^{\mathsf{T}}$ denotes the transpose of $\bm{X}$. For a given point $\widebar{\bm{x}}\in\mathcal{X}$ and a given set $E\subset\mathcal{X}$, $d:\mathcal{X}\times 2^{\mathcal{X}}\to\mathbb{R}_{+}\cup\{+\infty\}$ stands for the distance function, i.e.,
$d(\widebar{\bm{x}},E)\coloneqq\left\{\begin{array}[]{cl}\inf\{\left\lVert\bm{v}-\widebar{\bm{x}}\right\rVert\mid\bm{v}\in E\},&\mathrm{if}\ E\neq\emptyset;\\ +\infty,&\mathrm{if}\ E=\emptyset.\end{array}\right$. (6)
For a given $(a_{n})_{n=1}^{\infty}\subset\mathbb{R}$ and a given $\widebar{a}\in\mathbb{R}$, $a_{n}\searrow\widebar{a}$ and $(a_{n})_{n=1}^{\infty}\searrow\widebar{a}$ mean in this paper that $(a_{n})_{n=1}^{\infty}$ is monotonically nonincreasing together with $\lim_{n\to\infty}a_{n}=\widebar{a}$.
For a differentiable mapping $\mathcal{F}:\mathcal{Y}\to\mathcal{X}$, its Gâteaux derivative at $\bm{y}\in\mathcal{Y}$ is the linear operator $\mathrm{D}\mathcal{F}(\bm{y}):\mathcal{Y}\to\mathcal{X}$ defined by
$(\bm{v}\in\mathcal{Y})\quad\mathrm{D}\mathcal{F}(\bm{y})[\bm{v}]=\lim_{\mathbb{R}\setminus\{0\}\ni t\to 0}\frac{\mathcal{F}(\bm{y}+t\bm{v})-\mathcal{F}(\bm{y})}{t}.$ (7)
For a differentiable function $J:\mathcal{X}\to\mathbb{R}$, $\nabla J(\bm{x})\in\mathcal{X}$ is the gradient of $J$ at $\bm{x}\in\mathcal{X}$ if $\mathrm{D}J(\bm{x})[\bm{v}]={\left\langle\nabla J(\bm{x}),\bm{v}\right\rangle}$ for all $\bm{v}\in\mathcal{X}$. Note that for a differentiable mapping $\mathcal{F}:\mathcal{Y}\to\mathcal{X}$, we have the expression [^2] It can be checked by $(\forall\bm{v}\in\mathcal{Y})\ {\left\langle\nabla(J\circ\mathcal{F})(\bm{x}),\bm{v}\right\rangle}_{\mathcal{Y}}=\mathrm{D}(J\circ\mathcal{F})(\bm{x})[\bm{v}]=\mathrm{D}J(\mathcal{F}(\bm{x}))[\mathrm{D}\mathcal{F}(\bm{x})[\bm{v}]]={\left\langle\nabla J(\mathcal{F}(\bm{x})),\mathrm{D}\mathcal{F}(\bm{x})[\bm{v}]\right\rangle}_{\mathcal{X}}={\left\langle(\mathrm{D}\mathcal{F}(\bm{x}))^{*}[\nabla J(\mathcal{F}(\bm{x}))],\bm{v}\right\rangle}_{\mathcal{Y}}$. $\nabla(J\circ\mathcal{F})(\bm{y})=(\mathrm{D}\mathcal{F}(\bm{y}))^{*}[\nabla J(\mathcal{F}(\bm{y}))]\ (\bm{y}\in\mathcal{Y})$. For a function $J:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$, $\mathrm{dom}(J)\coloneqq\{\bm{x}\in\mathcal{X}\mid J(\bm{x})<+\infty\}\subset\mathcal{X}$ and $\mathrm{epi}(J):=\{(\bm{x},a)\in\mathcal{X}\times\mathbb{R}\mid J(\bm{x})\leq a\}\subset\mathcal{X}\times\mathbb{R}$ denote respectively the domain and the epigraph of $J$. Moreover, $J$ is said to be (i) proper if $\mathrm{dom}(J)\neq\emptyset$; (ii) lower semicontinuous if $\mathrm{epi}(J)$ is closed in $\mathcal{X}\times\mathbb{R}$; (iii) locally Lipschitz continuous at $\widebar{\bm{x}}\in\mathcal{X}$ if there exist an open neighborhood $\mathcal{N}_{\widebar{\bm{x}}}\subset\mathcal{X}$ of $\widebar{\bm{x}}$ and $L>0$ such that $\lvert J(\bm{x}_{1})-J(\bm{x}_{2})\rvert\leq L\left\lVert\bm{x}_{1}-\bm{x}_{2}\right\rVert\ (\bm{x}_{1},\bm{x}_{2}\in\mathcal{N}_{\widebar{\bm{x}}})$. For a function $J:\mathcal{X}\to\mathbb{R}\cup\{-\infty,+\infty\}$, the limit inferior [69, Definition 1.5] and the limit superior [69, p.13] of $J$ at $\widebar{\bm{x}}\in\mathcal{X}$ are defined respectively by
$\liminf_{\mathcal{X}\ni\bm{x}\to\widebar{\bm{x}}}J(\bm{x})$ $\coloneqq\sup_{\epsilon>0}\left(\inf_{\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert<\epsilon}J(\bm{x})\right)\coloneqq\sup\limits_{\epsilon>0}\left(\inf\left\{J(\bm{x})\in\mathbb{R}\cup\{-\infty,+\infty\}\mid\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert<\epsilon\right\}\right);$ (8)
$\limsup_{\mathcal{X}\ni\bm{x}\to\widebar{\bm{x}}}J(\bm{x})$ $\coloneqq\inf_{\epsilon>0}\left(\sup_{\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert<\epsilon}J(\bm{x})\right)\coloneqq\inf\limits_{\epsilon>0}\left(\sup\left\{J(\bm{x})\in\mathbb{R}\cup\{-\infty,+\infty\}\mid\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert<\epsilon\right\}\right).$ (9)
For a set sequence $(E_{n})_{n=1}^{\infty}$ with subsets $E_{n}\subset\mathcal{X}$, the outer limit [69, Definition 4.1] of $(E_{n})_{n=1}^{\infty}$ is defined by
$\mathop{\mathrm{Lim~sup}}_{n\to\infty}E_{n}\coloneqq\left\{\bm{v}\in\mathcal{X}\mid\exists\mathcal{N}\subset\mathbb{N},\exists\bm{v}_{n}\in E_{n}\ (n\in\mathcal{N})\ {\rm with}\ \lvert\mathcal{N}\rvert=\aleph_{0}\ {\rm and}\ \bm{v}=\lim_{\mathcal{N}\ni n\to\infty}\bm{v}_{n}\right\},$ (10)
where $\aleph_{0}$ stands for the cardinality of $\mathbb{N}$. With the outer limit of a set sequence, the outer limit [69, p.152] of a set-valued mapping $S:\mathcal{X}\rightrightarrows\mathcal{X}$ at $\widebar{\bm{x}}\in\mathcal{X}$ is defined as
$\mathop{\mathrm{Lim~sup}}_{\mathcal{X}\ni\bm{x}\to\widebar{\bm{x}}}S(\bm{x})\coloneqq\bigcup_{\mathcal{X}\ni\bm{x}_{n}\to\widebar{\bm{x}}}\mathop{\mathrm{Lim~sup}}_{n\to\infty}S(\bm{x}_{n})$ (11)
$=\left\{\bm{v}\in\mathcal{X}\mid\exists(\bm{x}_{n})_{n=1}^{\infty}\subset\mathcal{X},\ \exists\bm{v}_{n}\in S(\bm{x}_{n})\ \mathrm{with}\ \widebar{\bm{x}}=\lim_{n\to\infty}\bm{x}_{n}\ \mathrm{and}\ \bm{v}=\lim_{n\to\infty}\bm{v}_{n}\right\},$ (12)
where we follow a clear notation $\mathop{\mathrm{Lim~sup}}$” (used, e.g., in [18]). In this paper, even for the outer limit of a single-valued mapping $S:\mathcal{X}\to\mathcal{X}$, we use $\mathop{\mathrm{Lim~sup}}_{\mathcal{X}\ni\bm{x}\to\widebar{\bm{x}}}S(\bm{x})\coloneqq\mathop{\mathrm{Lim~sup}}_{\mathcal{X}\ni\bm{x}\to\widebar{\bm{x}}}\{S(\bm{x})\}$ in order to distinguish the standard notation for the limit superior $\limsup$” in (9).
2 Preliminaries
2.1 Subdifferential and Normal Cone
We review some necessary notions, namely subdifferentials and normal cones, in nonsmooth analysis. We refer to the seminal book [69] for a thorough review.
Definition 2.1 (Subdifferential [69, Def. 8.3]).
For a possibly nonconvex function $J:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$, the limiting (or general) subdifferential of $J$ at $\widebar{\bm{x}}\in\mathcal{X}$ is defined by
$(\widebar{\bm{x}}\in\mathrm{dom}(J))\quad\partial_{\rm L}J(\widebar{\bm{x}})\coloneqq\mathop{\mathrm{Lim~sup}}_{\mathcal{X}\ni\bm{x}\to\widebar{\bm{x}},\ J(\bm{x})\to J(\widebar{\bm{x}})}\partial_{\rm F}J(\bm{x}),$ (13)
where $\partial_{\rm F}J(\widebar{\bm{x}})\coloneqq\left\{\bm{v}\in\mathcal{X}\mid\liminf_{\mathcal{X}\setminus\{\widebar{\bm{x}}\}\ni\bm{x}\to\widebar{\bm{x}}}\frac{J(\bm{x})-J(\widebar{\bm{x}})-{\left\langle\bm{v},\bm{x}-\widebar{\bm{x}}\right\rangle}}{\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert}\geq 0\right\}$ is called the Fréchet (or regular) subdifferential of $J$ at $\widebar{\bm{x}}\in\mathrm{dom}(J)$, and $\partial_{\rm L}J(\widebar{\bm{x}})$ and $\partial_{\rm F}J(\widebar{\bm{x}})$ at $\widebar{\bm{x}}\notin\mathrm{dom}(J)$ are understood as $\emptyset$ (Note: for a convex function $J$, $\partial_{\rm L}J$ and $\partial_{\rm F}J$ are identical to the standard convex subdifferential of $J$ [69, Prop. 8.12]).
By using the subdifferentials, the general normal cone $N_{C}(\widebar{\bm{x}})$ and the regular normal cone $\widehat{N}_{C}(\widebar{\bm{x}})$ to a nonempty subset $C\subset\mathcal{X}$ at $\widebar{\bm{x}}\in C$ are defined (see [69, Def. 6.3, Prop. 6.5, and Ex. 8.14]) respectively by
$N_{C}(\widebar{\bm{x}})$ $\coloneqq\partial_{\rm L}\iota_{C}(\widebar{\bm{x}})=\mathop{\mathrm{Lim~sup}}_{C\ni\bm{x}\to\widebar{\bm{x}}}\partial_{\rm F}\iota_{C}(\widebar{\bm{x}})=\mathop{\mathrm{Lim~sup}}_{C\ni\bm{x}\to\widebar{\bm{x}}}\widehat{N}_{C}(\widebar{\bm{x}})$ (14)
$\widehat{N}_{C}(\widebar{\bm{x}})$ $\coloneqq\partial_{\rm F}\iota_{C}(\widebar{\bm{x}})=\left\{\bm{v}\in\mathcal{X}\mid\limsup_{C\setminus\{\widebar{\bm{x}}\}\ni\bm{x}\to\widebar{\bm{x}}}\frac{{\left\langle\bm{v},\bm{x}-\widebar{\bm{x}}\right\rangle}}{\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert}\leq 0\right\},$ (15)
where $N_{C}(\widebar{\bm{x}})$ and $\widehat{N}_{C}(\widebar{\bm{x}})$ at $\widebar{\bm{x}}\notin C$ are understood as $\emptyset$.
With notions of normal cone operators, we define regularities of a subset and functions as follows. Such regularities ensure the chain and sum rules for the limiting subdifferential as shown in Lemma 2.4 below (see also Fact B.4 in Appendix B).
Definition 2.2 (Regularity).
  1. (a)
    (Clarke regularity [69, Def. 6.4]) A nonempty set $C\subset\mathcal{X}$ is said to be Clarke regular at $\widebar{\bm{x}}\in C$ if (i) $N_{C}(\widebar{\bm{x}})=\widehat{N}_{C}(\widebar{\bm{x}})$ (Note: the inclusion $N_{C}(\widebar{\bm{x}})\supset\widehat{N}_{C}(\widebar{\bm{x}})$ always holds), and (ii) $C\cap\overline{\mathcal{N}}_{\widebar{\bm{x}}}$ is closed for some closed neighborhood [^3] In other words, $\overline{\mathcal{N}}_{\widebar{\bm{x}}}$ is a closed set that contains $\widebar{\bm{x}}$ as an interior point of $\overline{\mathcal{N}}_{\widebar{\bm{x}}}$. $\overline{\mathcal{N}}_{\widebar{\bm{x}}}\subset\mathcal{X}$ of $\widebar{\bm{x}}$. Moreover, $C$ is said to be Clarke regular if it is Clarke regular at every $\widebar{\bm{x}}\in C$.
  2. (b)
    (Subdifferential regularity [69, Def. 7.25]) A proper lower semicontinuous function $J:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ is said to be subdifferentially regular at $\widebar{\bm{x}}\in\mathrm{dom}(J)$ if $\mathrm{epi}(J)\subset\mathcal{X}\times\mathbb{R}$ is Clarke regular at $(\widebar{\bm{x}},J(\widebar{\bm{x}}))$. Moreover, $J$ is said to be subdifferentially regular if it is subdifferentially regular at every $\widebar{\bm{x}}\in\mathrm{dom}(J)$. We note that all functions in this paper are shown to be subdifferentially regular (see, e.g., Lemma 2.4 below and also Example B.2).
A typical Clarke regular set is a closed convex set [69, Thm. 6.9]. Another important example of Clarke regular sets is an embedded submanifold of the Euclidean space $\mathcal{X}$ [69, Exm. 6.8] as follows.
Example 2.3 (Embedded submanifold of $\mathcal{X}$ [69, Exm. 6.8]).
Let $(\emptyset\neq)C\subset\mathcal{X}$ be a $d(\leq\dim(\mathcal{X}))$-dimensional embedded submanifold of $\mathcal{X}$, i.e., for each $\widebar{\bm{x}}\in C$, there exist (i) an open neighborhood $\mathcal{N}_{\widebar{\bm{x}}}\subset\mathcal{X}$ of $\widebar{\bm{x}}$ in $\mathcal{X}$ and (ii) a continuously differentiable mapping $H_{\widebar{\bm{x}}}:\mathcal{N}_{\widebar{\bm{x}}}\to\mathbb{R}^{\dim(\mathcal{X})-d}$ such that $C\cap\mathcal{N}_{\widebar{\bm{x}}}=H_{\widebar{\bm{x}}}^{-1}(\bm{0})$ and $\mathrm{Rank}(\mathrm{D}H_{\widebar{\bm{x}}}(\widebar{\bm{x}}))=\dim(\mathcal{X})-d$, where $H_{\widebar{\bm{x}}}$ is called a local defining map [16, Def. 3.10] (Note: by Whitney’s embedding theorem (see, e.g., [54, Thm. 6.15]), every $d$-dimensional smooth manifold can be seen as an embedded submanifold in $\mathbb{R}^{2d+1}$). Then, by [69, Def. 6.4 and Exm. 6.8], the embedded submanifold $C$ is Clarke regular, and we have the following expression:
$(\widebar{\bm{x}}\in C)\quad N_{C}(\widebar{\bm{x}})=\widehat{N}_{C}(\widebar{\bm{x}})=\mathrm{Ran}((\mathrm{D}H_{\widebar{\bm{x}}}(\widebar{\bm{x}}))^{*})\subset\mathcal{X},\quad T_{C}(\widebar{\bm{x}})=\left(N_{C}(\widebar{\bm{x}})\right)^{\perp}\subset\mathcal{X},$ (16)
where the orthogonal complement subspace $T_{C}(\widebar{\bm{x}})$ of $N_{C}(\widebar{\bm{x}})$ is called the tangent cone to $C$ at $\widebar{\bm{x}}$ (Note: $N_{C}(\widebar{\bm{x}})$ and $T_{C}(\widebar{\bm{x}})$ are also called respectively the normal space and the tangent space to the manifold $C$ at $\widebar{\bm{x}}\in C$ [16, Thm. 3.15, Def. 3.16, and p.107]).
Lemma 2.4 (Subdifferential calculus for Problems 1.1 and 1.4).
Consider $F=h+g\circ\mathfrak{S}$, $C\subset\mathcal{X}$ and $\bm{\varphi}$ in Problems 1.1 and 1.4. Then, $F$, $F+\iota_{C}$ and $F\circ\bm{\varphi}$ are subdifferentially regular, and we have the following expressions:
$(\widebar{\bm{x}}\in C)\quad$ $\partial_{\rm L}(F+\iota_{C})(\widebar{\bm{x}})=\partial_{\rm L}F(\widebar{\bm{x}})+N_{C}(\widebar{\bm{x}});$ (17)
$(\widebar{\bm{y}}\in\mathcal{Y})\quad$ $\partial_{\rm L}(F\circ\bm{\varphi})(\widebar{\bm{y}})=(\mathrm{D}\bm{\varphi}(\widebar{\bm{y}}))^{*}[\partial_{\rm L}F(\bm{\varphi}(\widebar{\bm{y}}))].$ (18)
Proof.
See Appendix B. ∎
The following states a relation between the normal cone to the parameterized set, and the parametrization $\bm{\varphi}$.
Fact 2.5 (Normal cone to image sets [69, Thm. 6.43]).
Let $\mathcal{X},\mathcal{Y}$ be Euclidean spaces, and $\bm{\varphi}:\mathcal{Y}\to\mathcal{X}$ a continuously differentiable mapping. Then, for every $\widebar{\bm{x}}\in\bm{\varphi}(\mathcal{Y})$,
$\widehat{N}_{\bm{\varphi}(\mathcal{Y})}(\widebar{\bm{x}})\subset\bigcap_{\widebar{\bm{y}}\in\bm{\varphi}^{-1}(\widebar{\bm{x}})}\left\{\bm{v}\in\mathcal{X}\mid(\mathrm{D}\bm{\varphi}(\widebar{\bm{y}}))^{*}[\bm{v}]\in\widehat{N}_{\mathcal{Y}}(\widebar{\bm{y}})=\{\bm{0}\}\right\}=\bigcap_{\widebar{\bm{y}}\in\bm{\varphi}^{-1}(\widebar{\bm{x}})}\mathrm{Ker}\left((\mathrm{D}\bm{\varphi}(\widebar{\bm{y}}))^{*}\right),$ (19)
where $\bm{\varphi}^{-1}(\widebar{\bm{x}})\coloneqq\{\widebar{\bm{y}}\in\mathcal{Y}\mid\bm{\varphi}(\widebar{\bm{y}})=\widebar{\bm{x}}\}\subset\mathcal{Y}$ is the preimage of $\widebar{\bm{x}}\in\mathcal{X}$ under $\bm{\varphi}$.
Proof.
The first inclusion is verified by [69, Thm. 6.43]. The equality follows by the relation $(\mathrm{D}\bm{\varphi}(\widebar{\bm{y}}))^{*}[\bm{v}]\in\{\bm{0}\}\Leftrightarrow\bm{v}\in\mathrm{Ker}\left((\mathrm{D}\bm{\varphi}(\widebar{\bm{y}}))^{*}\right)\ (\widebar{\bm{y}}\in\bm{\varphi}^{-1}(\widebar{\bm{x}}))$. ∎
2.2 Proximity Operator and Moreau Envelope
The proximity operator and the Moreau envelope have been serving as key ingredients for minimization of nonsmooth convex functions (see, e.g., [8, 28]). Even for weakly convex functions, they are useful in design of nonsmooth optimization algorithms. We note that closed-form expressions of the proximity operators for commonly-used weakly convex functions have been found, e.g., $\ell_{1}$-norm [8], Minimax Concave Penalty (MCP) [89], Smoothly Clipped Absolute Deviation (SCAD) [35], log-sum penalty [68], piece-wise exponential function [60], and arctangent penalty [40].
Definition 2.6 (Proximity operator and Moreau envelope [69, Def. 1.22]).
Let $g:\mathcal{Z}\to\mathbb{R}\cup\{+\infty\}$ be proper lower semicontinuous, and $\eta$-weakly convex, i.e., $g+\eta/2\left\lVert\cdot\right\rVert^{2}$ is convex with $\eta>0$. For $\mu\in(0,\eta^{-1})$, the proximity operator and the Moreau envelope of $g$ of index $\mu$ are defined respectively as
$(\widebar{\bm{z}}\in\mathcal{Z})\$$\mathrm{prox}_{\mu g}(\widebar{\bm{z}})\coloneqq\mathop{\mathrm{argmin}}\limits_{\bm{z}\in\mathcal{Z}}\left(g(\bm{z})+\frac{1}{2\mu}\left\lVert\bm{z}-\widebar{\bm{z}}\right\rVert^{2}\right);$ (20)
$(\widebar{\bm{z}}\in\mathcal{Z})\$${}^{\mu}g(\widebar{\bm{z}})\coloneqq\inf_{\bm{z}\in\mathcal{Z}}\left(g(\bm{z})+\frac{1}{2\mu}\left\lVert\bm{z}-\widebar{\bm{z}}\right\rVert^{2}\right)=g(\mathrm{prox}_{\mu g}(\widebar{\bm{z}}))+\frac{1}{2\mu}\left\lVert\mathrm{prox}_{\mu g}(\widebar{\bm{z}})-\widebar{\bm{z}}\right\rVert^{2},$ (21)
where $\mathrm{prox}_{\mu g}(\widebar{\bm{z}})$ is single-valued because the unique existence of the minimizer is guaranteed by the strong convexity of $g(\cdot)+\frac{1}{2\mu}\left\lVert\cdot-\widebar{\bm{z}}\right\rVert^{2}$.
Fact 2.7 (Properties of Moreau envelope of weakly convex function).
Let $g:\mathcal{Z}\to\mathbb{R}\cup\{+\infty\}$ be proper lower semicontinuous, and $\eta$-weakly convex with some $\eta>0$. Then, the following hold:
  1. (a)
    For any $\widebar{\bm{z}}\in\mathcal{Z}$ and $(\bm{z}_{n})_{n=1}^{\infty}\subset\mathcal{Z}$ satisfying $\lim_{n\to\infty}\bm{z}_{n}=\widebar{\bm{z}}$, we have $\lim_{n\to\infty}\mathrm{prox}_{\mu_{n}g}(\bm{z}_{n})=P_{\overline{\mathrm{dom}(\partial_{\rm L}g)}}(\widebar{\bm{z}})$ if $(\mu_{n})_{n=1}^{\infty}\searrow 0$ with $\mu_{n}\in(0,\eta^{-1})$ [7, Lemma 2.3], where $\mathrm{dom}(\partial_{\rm L}g)\coloneqq\{\bm{z}\in\mathcal{Z}\mid\partial_{\rm L}g(\bm{z})\neq\emptyset\}\subset\mathcal{Z}$, $\overline{(\cdot)}$ denotes the closure of a given set, and $P_{\overline{\mathrm{dom}(\partial_{\rm L}g)}}=\mathrm{prox}_{\iota_{\overline{\mathrm{dom}(\partial_{\rm L}g)}}}$ denotes the projection mapping onto the closed convex set $\overline{\mathrm{dom}(\partial_{\rm L}g)}\subset\mathcal{Z}$.
  2. (b)
    For any $\mu\in(0,\eta^{-1})$, (i) $^{\mu}g:\mathcal{Z}\to\mathbb{R}$ is continuously differentiable with $\nabla{}^{\mu}g=\frac{\mathrm{Id}-\mathrm{prox}_{\mu g}}{\mu}$, and (ii) $\nabla{}^{\mu}g$ is Lipschitz continuous with a Lipschitz constant $\max\left\{\mu^{-1},\frac{\eta}{1-\eta\mu}\right\}$ [42, Cor. 3.4].
  3. (c)
    $(\widebar{\bm{z}}\in\mathcal{Z},\mu\in(0,\eta^{-1}))\quad\nabla{}^{\mu}g(\widebar{\bm{z}})\in\partial_{\rm L}g(\mathrm{prox}_{\mu g}(\widebar{\bm{z}}))$ [13, Lemma 3.2].
  4. (d)
    Assume that $g$ is Lipschitz continuous with a Lipschitz constant $L_{g}>0$. By [13, Lemma 3.3] and [67, Prop. 3.4] [^4] In [67, Prop. 3.4], the inequality (23) is proved under the convexity of $g$. However, the proof in [67, Prop. 3.4] can be applied to the weakly convex case as well. , the following hold:
    $(\widebar{\bm{z}}\in\mathcal{Z},\mu\in(0,\eta^{-1}))\quad$ $\left\lVert\nabla{}^{\mu}g(\widebar{\bm{z}})\right\rVert\leq L_{g};$ (22)
    $(\widebar{\bm{z}}\in\mathcal{Z},\mu\in(0,\eta^{-1}))\quad$ ${}^{\mu}g(\widebar{\bm{z}})\leq g(\widebar{\bm{z}})\leq{}^{\mu}g(\widebar{\bm{z}})+\frac{1}{2}\mu L_{g}^{2}.$ (23)
    Moreover, for $0<\mu_{2}<\mu_{1}<\eta^{-1}$, we have [69, Thm. 1.25], [13, Lemma 4.1]
    $(\widebar{\bm{z}}\in\mathcal{Z})\quad{}^{\mu_{1}}g(\widebar{\bm{z}})\leq{}^{\mu_{2}}g(\widebar{\bm{z}})\leq{}^{\mu_{1}}g(\widebar{\bm{z}})+\frac{1}{2}\frac{\mu_{1}-\mu_{2}}{\mu_{2}}\mu_{1}L_{g}^{2}.$ (24)
3 Relation between First-order Optimality Conditions of Problems 1.1 and 1.4
This section presents a relation between the original Problem 1.1 and the parameterized Problem 1.4. We begin with the notions of optimality.
Definition 3.1 (Optimality).
Let $J:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ be proper.
  1. (a)
    (Local minimizer) $\bm{x}^{\star\star}\in\mathrm{dom}(J)$ is said to be a local minimizer of $J$ if there exists an open neighborhood $\mathcal{N}_{\bm{x}^{\star\star}}\subset\mathcal{X}$ of $\bm{x}^{\star\star}$ such that $J(\bm{x}^{\star\star})\leq J(\bm{x})\ (\bm{x}\in\mathcal{N}_{\bm{x}^{\star\star}})$.
  2. (b)
    (Stationary point) $\bm{x}^{\star}\in\mathcal{X}$ is said to be a stationary point [^5] A stationary point $\bm{x}^{\star}\in\mathcal{X}$ is also called a limiting stationary point of $J$ in order to distinguish from a Fréchet stationary point where $\partial_{\rm F}J(\bm{x}^{\star})\ni\bm{0}$ holds. However, these stationarities are identical in a case where $J$ is subdifferentially regular in Definition 2.2 (b) (see [69, Cor. 8.11]). Since all functions in this paper are subdifferentially regular, we do not distinguish these stationarities in this paper. of $J$ if $\partial_{\rm L}J(\bm{x}^{\star})\ni\bm{0}$.
By Fermat’s rule [69, Thm. 10.1], for a proper function $J:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$, $\bm{x}^{\star\star}\in\mathcal{X}$ is a stationary point of $J$ if $\bm{x}^{\star\star}$ is a local minimizer of $J$. Thus, the condition $\partial_{\rm L}J(\bm{x}^{\star})\ni\bm{0}$ at $\bm{x}^{\star}\in\mathcal{X}$ seems to be the most versatile expressions of the first-order optimality condition for minimization of $J$ at $\bm{x}^{\star}$ (see, e.g., [69, p.207]). For this reason, we focus on finding such stationary points of Problem 1.1 and Problem 1.4 throughout this paper. By Lemma 2.4, the first-order optimality conditions can be expressed respectively with $F=h+g\circ\mathfrak{S}$ as:
[…]
for $\bm{x}^{\star}\in C$ to be a stationary point of Problem 1.1, and
[…]
${\rm or\ equivalently},\ d\left(\bm{0},\partial_{\rm L}(F\circ\bm{\varphi})(\bm{y}^{\star})\right)=0$ (27)
for $\bm{y}^{\star}\in\mathcal{Y}$ to be a stationary point of Problem 1.4.
Theorem 3.2 below presents a relation between (25) and (26). Based on Theorem 3.2, the conditions (25) and (26) are equivalent with $\bm{x}^{\star}=\bm{\varphi}(\bm{y}^{\star})$ under a certain condition (see Corollary 3.4). In a special case where the cost function $F$ is continuously differentiable, a similar relation between the first-order optimality conditions is found in [56, Thm. 2.4]. We note that, even if $C$ is not an embedded submanifold of $\mathcal{X}$ but is just a Clarke regular subset of $\mathcal{X}$, the same statements in Theorem 3.2 hold.
Theorem 3.2 (Point-wise relation between the first-order optimality conditions (25) and (26)).
Consider Problems 1.1 and 1.4. Let $\bm{x}^{\star}\coloneqq\bm{\varphi}(\bm{y}^{\star})\in C$ with some $\bm{y}^{\star}\in\mathcal{Y}$. Then, the following hold:
  1. (a)
    If $\bm{x}^{\star}$ satisfies (25), then $\bm{y}^{\star}$ satisfies (26).
  2. (b)
    Assume $N_{C}(\bm{x}^{\star})\supset\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*})$. If $\bm{y}^{\star}$ satisfies (26), then $\bm{x}^{\star}$ satisfies (25).
To prove Theorem 3.2, we use the following lemma.
Lemma 3.3.
Consider Problem 1.1. Let $\bm{\varphi}:\mathcal{Y}\to C$ be continuously differentiable, and let $\widebar{\bm{x}}\coloneqq\bm{\varphi}(\widebar{\bm{y}})\in C$ with some $\widebar{\bm{y}}\in\mathcal{Y}$. Then, $N_{C}(\widebar{\bm{x}})\subset\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\widebar{\bm{y}}))^{*})$.
Proof.
Recall $\widehat{N}_{\bm{\varphi}(\mathcal{Y})}(\widebar{\bm{x}})\subset\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\widebar{\bm{y}}))^{*})$ from Fact 2.5. Moreover, we have $\widehat{N}_{\bm{\varphi}(\mathcal{Y})}(\widebar{\bm{x}})\supset\widehat{N}_{C}(\widebar{\bm{x}})=N_{C}(\widebar{\bm{x}})$, where $\supset$” follows [^6] We can check the inclusion by showing $\limsup_{\bm{\varphi}(\mathcal{Y})\setminus\{\widebar{\bm{x}}\}\ni\bm{x}\to\widebar{\bm{x}}}\frac{{\left\langle\bm{v},\bm{x}-\widebar{\bm{x}}\right\rangle}}{\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert}\leq\limsup_{C\setminus\{\widebar{\bm{x}}\}\ni\bm{x}\to\widebar{\bm{x}}}\frac{{\left\langle\bm{v},\bm{x}-\widebar{\bm{x}}\right\rangle}}{\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert}$ from (15). The inequality can be verified with (9) from $\inf\limits_{\epsilon>0}\left(\sup\limits_{\begin{subarray}{c}0<\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert<\epsilon\\ \bm{x}\in\bm{\varphi}(\mathcal{Y})\end{subarray}}\frac{{\left\langle\bm{v},\bm{x}-\widebar{\bm{x}}\right\rangle}}{\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert}\right)\overset{\bm{\varphi}(\mathcal{Y})\subset C}{\leq}\inf\limits_{\epsilon>0}\left(\sup\limits_{\begin{subarray}{c}0<\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert<\epsilon\\ \bm{x}\in C\end{subarray}}\frac{{\left\langle\bm{v},\bm{x}-\widebar{\bm{x}}\right\rangle}}{\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert}\right)$. by $C\supset\bm{\varphi}(\mathcal{Y})$ (Note: $\supset$” can be replaced by “==” if $\bm{\varphi}(\mathcal{Y})=C$), and the equality follows by the Clarke regularity of $C$ (see (16) in Example 2.3). By combining these inclusions, we complete the proof. ∎
Proof of Theorem 3.2.
For simplicity, let $F\coloneqq h+g\circ\mathfrak{S}$.
(a) Let $\bm{x}^{\star}$ satisfy (25), i.e., there exists $\bm{v}\in\partial_{\rm L}F(\bm{x}^{\star})$ such that $\bm{v}\in N_{C}(\bm{x}^{\star})$. Then, Lemma 3.3 implies $\bm{v}\in N_{C}(\bm{x}^{\star})\subset\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*})$. By combining this inclusion with the linearity of $\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*})$, we have $\bm{v}\in\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*})$. Thus, $(\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*}\left[\partial_{\rm L}F\left(\bm{\varphi}(\bm{y}^{\star})\right)\right]\ni\bm{0}$, i.e., $\bm{y}^{\star}$ satisfies (26).
(b) Let $\bm{y}^{\star}\in\mathcal{Y}$ satisfy (26), i.e., there exists $\bm{v}\in\partial_{\rm L}F(\bm{\varphi}(\bm{y}^{\star}))$ such that $(\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*}[\bm{v}]=\bm{0}$. The linearity of $\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*})$ implies $\bm{v}\in\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*})$, and then the assumption $N_{C}(\bm{x}^{\star})\supset\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*})$ yields $\bm{v}\in N_{C}(\bm{\varphi}(\bm{y}^{\star}))$. Thus, we have $\bm{0}\in\partial_{\rm L}F(\bm{\varphi}(\bm{y}^{\star}))+N_{C}(\bm{\varphi}(\bm{y}^{\star}))=\partial_{\rm L}F(\bm{x}^{\star})+N_{C}(\bm{x}^{\star})$, i.e., $\bm{x}^{\star}$ satisfies (25). ∎
Corollary 3.4 (Special case for (25) and (26) to be equivalent).
[…]
Proof.
[…]
Remark 3.5 (Case for $\bm{\varphi}(\mathcal{Y})\subsetneq C$).
Even in a case where $\bm{\varphi}:\mathcal{Y}\to C$ is not surjective onto $C$, the same conclusions as in Theorem 3.2 (a) and (b), and in Corollary 3.4 can be achieved because the surjectivity of $\bm{\varphi}$ is not used in their proofs at all.
Remark 3.6 (On submersion condition).
A parametrization $\bm{\varphi}$ of $C$ in Problem 1.1 does not automatically satisfy the submersion condition $\mathrm{Ran}(\mathrm{D}\bm{\varphi}(\bm{y}))=T_{C}(\bm{\varphi}(\bm{y}))$ at each point $\bm{y}\in\mathcal{Y}$ (see, e.g., [47] and footnote21 in Appendix A.2, for a parametrization with the submersion condition). However, even in the absence of the submersion condition [^9] For a manifold $C$ in Problem 1.1, there exists a parametrization satisfying the submersion condition locally. Indeed, for $\bm{x}\in C$, the exponential map $\mathrm{Exp}_{\bm{x}}:T_{C}(\bm{x})\to C$ (see Example 1.3 (c)) is a local diffeomorphism at $\bm{0}\in T_{C}(\bm{x})$ [55, Lemma 6.16], hence submersion over an open neighborhood of $\bm{0}$. By using a family of parametrizations satisfying the submersion condition locally, an adaptive parametrization strategy has also been proposed [58, 29, 50] for the smooth case of the cost function. of $\bm{\varphi}$, Theorem 3.2 (a) ensures that the first-order optimality condition (26) of Problem 1.4 at $\bm{y}^{\star}\in\mathcal{Y}$ is a necessary condition for the first-order optimality condition (25) of Problem 1.1 at $\bm{x}^{\star}\coloneqq\bm{\varphi}(\bm{y}^{\star})\in C$, hence is a necessary condition for the local optimality of Problem 1.1 at $\bm{x}^{\star}$.
Theorem 3.2 and Corollary 3.4 encourage us to focus on finding a stationary point of Problem 1.4 for our goal of finding a stationary point of Problem 1.1. Since Problem 1.4 is an unconstrained optimization problem thanks to the parametrization $\bm{\varphi}$, Problem 1.4 seems to be more tractable than Problem 1.1 in terms of designing iterative algorithms. Indeed, iterative algorithms with parametrization have been proposed in a case where the cost function $F=h+g\circ\mathfrak{S}$ is smooth [84, 36, 43, 41, 58, 39, 47, 48, 29, 50]. In Section 4, along such parametrization strategies, we propose an iterative algorithm applicable to Problem 1.4 even with a nonsmooth cost function $F$.
We remark that, as an alternative strategy to the parametrization strategy for translating Problem 1.1 into an unconstrained problem, penalty-based strategies have been proposed, e.g., [65, Ch. 12][82], in a case where $C$ is defined as equality constraints. See Appendix A.2 for a further discussion of the penalty-based and parametrization strategies.
4 Variable Smoothing Algorithm for Problem 1.4
4.1 Properties of smoothed surrogate function for Problem 1.4
In this section, we investigate properties of our smoothed surrogate function $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}\ (\mu\in(0,\eta^{-1}))$ of the cost function $(h+g\circ\mathfrak{S})\circ\bm{\varphi}$ in Problem 1.4, where $^{\mu}g$ is the Moreau envelope of $g$ in (21). The goal of this section is to present an asymptotic property, called gradient (sub-)consistency, of the gradient of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}\ (\mu\searrow 0)$ to approximate subgradients of $(h+g\circ\mathfrak{S})\circ\bm{\varphi}$ (see Theorem 4.4 and Remark 4.5). By exploiting the gradient (sub-)consistency of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$, we will establish an asymptotic convergence analysis for the proposed variable smoothing algorithm in Section 4.2. We summarize basic properties of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$, with $\mu\in(0,\eta^{-1})$ fixed, which follow from direct applications of Fact 2.7 (d).
Lemma 4.1.
Let $F^{\langle\mu\rangle}\circ\bm{\varphi}\coloneqq(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}\ (\mu\in(0,\eta^{-1}))$ be a smoothed surrogate function of $F\circ\bm{\varphi}=(h+g\circ\mathfrak{S})\circ\bm{\varphi}$ in Problem 1.4. Then, for $\widebar{\bm{y}}\in\mathcal{Y}$ and $\mu\in(0,\eta^{-1})$, we have the following expressions by (21) and Fact 2.7 (b):
$\begin{cases}F^{\langle\mu\rangle}\circ\bm{\varphi}(\widebar{\bm{y}})=h(\bm{\varphi}(\widebar{\bm{y}}))+g\left(\mathrm{prox}_{\mu g}\left(\mathfrak{S}(\bm{\varphi}(\widebar{\bm{y}}))\right)\right)+\frac{1}{2\mu}\left\lVert\mathrm{prox}_{\mu g}\left(\mathfrak{S}\left(\bm{\varphi}(\widebar{\bm{y}})\right)\right)-\mathfrak{S}\left(\bm{\varphi}(\widebar{\bm{y}})\right)\right\rVert^{2},\\ \nabla(F^{\langle\mu\rangle}\circ\bm{\varphi})(\widebar{\bm{y}})=(\mathrm{D}\bm{\varphi}(\bm{y}))^{*}\left[\nabla h(\bm{\varphi}(\widebar{\bm{y}}))+\left(\mathrm{D}\mathfrak{S}\left(\bm{\varphi}(\widebar{\bm{y}})\right)\right)^{*}\left[\frac{\mathfrak{S}(\bm{\varphi}(\widebar{\bm{y}}))-\mathrm{prox}_{\mu g}\left(\mathfrak{S}(\bm{\varphi}(\widebar{\bm{y}}))\right)}{\mu}\right]\right].\end{cases}$ (28)
With a Lipschitz constant $L_{g}>0$ of $g$, we have
$(\widebar{\bm{y}}\in\mathcal{Y},\mu\in(0,\eta^{-1}))\quad$ $F^{\langle\mu\rangle}\circ\bm{\varphi}(\widebar{\bm{y}})\leq F\circ\bm{\varphi}(\widebar{\bm{y}})\leq F^{\langle\mu\rangle}\circ\bm{\varphi}(\widebar{\bm{y}})+\frac{1}{2}\mu L_{g}^{2},$ (29)
$(\mu\in(0,\eta^{-1}))\quad$ $-\infty<\inf_{\bm{y}\in\mathcal{Y}}F^{\langle\mu\rangle}\circ\bm{\varphi}(\bm{y}).$ (30)
Moreover, with $0<\mu_{2}<\mu_{1}<\eta^{-1}$, we have
$(\widebar{\bm{y}}\in\mathcal{Y})\quad F^{\langle\mu_{1}\rangle}\circ\bm{\varphi}(\widebar{\bm{y}})\leq F^{\langle\mu_{2}\rangle}\circ\bm{\varphi}(\widebar{\bm{y}})\leq F^{\langle\mu_{1}\rangle}\circ\bm{\varphi}(\widebar{\bm{y}})+L_{g}^{2}(\mu_{1}-\mu_{2}).$ (31)
Proof.
The expressions (28) follow from (21) and Fact 2.7 (b). The inequality (29) follows from (23). The inequality (30) is verified by (29) and $\inf_{\bm{y}\in\mathcal{Y}}F\circ\bm{\varphi}(\bm{y})>$-$\infty$ from the condition (iv) in Problem 1.1. To derive (31), it suffices to show
$(0<\mu_{2}<\mu_{1}<\eta^{-1},\ \widebar{\bm{z}}\in\mathcal{Z})\quad{}^{\mu_{1}}g(\widebar{\bm{z}})\leq{}^{\mu_{2}}g(\widebar{\bm{z}})\leq{}^{\mu_{1}}g(\widebar{\bm{z}})+L_{g}^{2}(\mu_{1}-\mu_{2}).$ (32)
For the case $\frac{\mu_{1}}{\mu_{2}}\leq 2$, the inequality (32) follows from (24). Now suppose $\frac{\mu_{1}}{\mu_{2}}>2$, and define $\widetilde{\mu}_{t}\coloneqq 2^{-t}\mu_{1}\ (t=0,1,2,\ldots,l-1)$ and $\widetilde{\mu}_{l}\coloneqq\mu_{2}(\geq 2^{-l}\mu_{1})$ with $l\coloneqq\lceil\log_{2}(\frac{\mu_{1}}{\mu_{2}})\rceil\geq 2$, where $\lceil\cdot\rceil$ is the ceiling function. Then, each pair $(\widetilde{\mu}_{t},\widetilde{\mu}_{t+1})$ satisfies $\frac{\widetilde{\mu}_{t}}{\widetilde{\mu}_{t+1}}\leq 2$, and thus the inequality (32) with $(\mu_{1},\mu_{2})\coloneqq(\widetilde{\mu}_{t},\widetilde{\mu}_{t+1})$ (already proved for the case $\frac{\mu_{1}}{\mu_{2}}\leq 2$) yields $0\leq{}^{\widetilde{\mu}_{t+1}}g(\widebar{\bm{z}})-{}^{\widetilde{\mu}_{t}}g(\widebar{\bm{z}})\leq L_{g}^{2}(\widetilde{\mu}_{t}-\widetilde{\mu}_{t+1})\ (\widebar{\bm{z}}\in\mathcal{Z})$. Hence, we get $0\leq{}^{\mu_{2}}g(\widebar{\bm{z}})-{}^{\mu_{1}}g(\widebar{\bm{z}})=\sum_{t=0}^{l-1}({}^{\widetilde{\mu}_{t+1}}g(\widebar{\bm{z}})-{}^{\widetilde{\mu}_{t}}g(\widebar{\bm{z}}))\leq L_{g}^{2}\sum_{t=0}^{l-1}(\widetilde{\mu}_{t}-\widetilde{\mu}_{t+1})=L_{g}^{2}(\mu_{1}-\mu_{2})$, implying the inequality (32) also follows for the case $\frac{\mu_{1}}{\mu_{2}}>2$. ∎
The Lipschitz continuity of the gradient of a cost function is one of important ingredients for optimization to ensure the descent lemma (see, e.g., [9, Lemma 5.7]). The following presents a sufficient condition for the Lipschitz constant of the gradient of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ (see also Remark 4.3 for its further sufficient conditions).
Proposition 4.2 (Lipschitzian of $\nabla((h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi})$).
Consider a smoothed surrogate function $F^{\langle\mu\rangle}\circ\bm{\varphi}\coloneqq(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}\ (\mu\in(0,\eta^{-1}))$ of $F\circ\bm{\varphi}=(h+g\circ\mathfrak{S})\circ\bm{\varphi}$ in Problem 1.4, and the following conditions:
  1. (A1)
    The operator norm of $(\mathrm{D}\mathfrak{S}(\cdot))^{*}$ is bounded above by $\kappa_{\mathfrak{S}}>0$ over the convex hull $\mathrm{Conv}(C)\coloneqq\{t\bm{x}_{1}+(1-t)\bm{x}_{2}\in\mathcal{X}\mid\bm{x}_{1},\bm{x}_{2}\in C,t\in[0,1]\}\subset\mathcal{X}$ of $C$, i.e.,
    $(\exists\kappa_{\mathfrak{S}}>0,\forall\bm{x}\in\mathrm{Conv}(C))\quad\left\lVert(\mathrm{D}\mathfrak{S}(\bm{x}))^{*}\right\rVert_{\mathrm{op}}\leq\kappa_{\mathfrak{S}};$ (33)
  2. (A2)
    $(\mathrm{D}\mathfrak{S}(\cdot))^{*}$ is Lipschitz continuous with a Lipschitz constant $L_{\mathrm{D}\mathfrak{S}}>0$ over $C$, i.e.,
    $(\exists L_{\mathrm{D}\mathfrak{S}}>0,\forall\bm{x}_{1},\bm{x}_{2}\in C)\ \left\lVert(\mathrm{D}\mathfrak{S}(\bm{x}_{1}))^{*}-(\mathrm{D}\mathfrak{S}(\bm{x}_{2}))^{*}\right\rVert_{\mathrm{op}}\leq L_{\mathrm{D}\mathfrak{S}}\left\lVert\bm{x}_{1}-\bm{x}_{2}\right\rVert;$ (34)
  3. (A3)
    The norm of $\nabla h$ is bounded above by $\kappa_{h}>0$ over $C$, i.e.,
    $(\exists\kappa_{h}>0,\forall\bm{x}\in C)\ \left\lVert\nabla h(\bm{x})\right\rVert\leq\kappa_{h};$ (35)
  4. (A4)
    The operator norm of $(\mathrm{D}\bm{\varphi}(\cdot))^{*}$ is bounded above by $\kappa_{\bm{\varphi}}>0$ over $\mathcal{Y}$;
  5. (A5)
    $(\mathrm{D}\bm{\varphi}(\cdot))^{*}$ is Lipschitz continuous with a Lipschitz constant $L_{\mathrm{D}\bm{\varphi}}>0$ over $\mathcal{Y}$.
Then, for $\mu\in(0,2^{-1}\eta^{-1}]$, $F^{\langle\mu\rangle}\circ\bm{\varphi}$ is continuously differentiable, and the following hold:
  1. (a)
    If the conditions (A1) and (A2) hold, then $\nabla F^{\langle\mu\rangle}$ is Lipschitz continuous with a Lipschitz constant $L_{\nabla F^{\langle\mu\rangle}}\coloneqq L_{\nabla h}+L_{g}L_{\mathrm{D}\mathfrak{S}}+\kappa_{\mathfrak{S}}^{2}\mu^{-1}>0$ over $C$.
  2. (b)
    If the conditions (A1)-(A5) hold, then $\nabla(F^{\langle\mu\rangle}\circ\bm{\varphi})$ is Lipschitz continuous with a Lipschitz constant $L_{\nabla(F^{\langle\mu\rangle}\circ\bm{\varphi})}\coloneqq\varpi_{1}+\varpi_{2}\mu^{-1}>0$ over $\mathcal{Y}$, where constants $\varpi_{1}\coloneqq\kappa_{\bm{\varphi}}^{2}(L_{\nabla h}+L_{g}L_{\mathrm{D}\mathfrak{S}})+L_{\mathrm{D}\bm{\varphi}}(\kappa_{h}+\kappa_{\mathfrak{S}}L_{g})$ and $\varpi_{2}\coloneqq\kappa_{\bm{\varphi}}^{2}\kappa_{\mathfrak{S}}^{2}$ do not depend on $\mu$.
Proof.
See Appendix C. ∎
Remark 4.3 (Sufficient conditions for assumptions in Proposition 4.2).
  1. (a)
    (Sufficient conditions regarding $\mathfrak{S}$) In a case where $\mathfrak{S}$ is a linear operator, the conditions (A1) and (A2) in Proposition 4.2 are automatically satisfied. In another case where $C$ is compact and $\mathfrak{S}$ is twice continuously differentiable, the conditions (A1)(A2) and (A3) are automatically satisfied.
  2. (b)
    (Sufficient conditions regarding $\bm{\varphi}$) The conditions (A4) and (A5) are automatically satisfied in a case where $C$ is a linear subspace of $\mathcal{X}$ and $\bm{\varphi}$ is a linear operator. Even for the other cases, we will present an example of $(C,\bm{\varphi})$, in Corollary 5.1 (see also Lemma E.2), achieving the conditions (A4) and (A5).
From (27) and Theorem 3.2, we aim to find a stationary point $\bm{y}^{\star}\in\mathcal{Y}$ of Problem 1.4 such that $d\left(\bm{0},\partial_{\rm L}(F\circ\bm{\varphi})(\bm{y}^{\star})\right)=0$ with $F=h+g\circ\mathfrak{S}$. Theorem 4.4 below presents an upper bound of this distance with the gradient of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ in (38) by showing the gradient sub-consistency of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ for $F$ (see also Remark 4.5).
Theorem 4.4 (Gradient sub-consistency of smoothed surrogate functions).
  1. (a)
    Let $\mathcal{Y}$ and $\mathcal{Z}$ be Euclidean spaces. Let $J:\mathcal{Z}\to\mathbb{R}$ be $\eta$-weakly convex, and $\mathcal{F}:\mathcal{Y}\to\mathcal{Z}$ be continuously differentiable. Then, $^{\mu}J\circ\mathcal{F}$ enjoys the gradient sub-consistency property (see Remark 4.5) for $J\circ\mathcal{F}$, i.e.,
    $(\widebar{\bm{y}}\in\mathcal{Y})\quad\mathop{\mathrm{Lim~sup}}_{\mathbb{R}_{++}\ni\mu\searrow 0,\ \mathcal{Y}\ni\bm{y}\to\widebar{\bm{y}}}\nabla({}^{\mu}J\circ\mathcal{F})(\bm{y})\subset\partial_{\rm L}(J\circ\mathcal{F})(\widebar{\bm{y}}).$ (36)
  2. (b)
    Consider a smoothed surrogate function $F^{\langle\mu\rangle}\circ\bm{\varphi}\coloneqq(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}\ (\mu\in(0,\eta^{-1}))$ of $F\circ\bm{\varphi}=(h+g\circ\mathfrak{S})\circ\bm{\varphi}$ in Problem 1.4 (see also Problem 1.1). Then, we have
    $\mathop{\mathrm{Lim~sup}}_{\mathbb{R}_{++}\ni\mu\searrow 0,\ \mathcal{Y}\ni\bm{y}\to\widebar{\bm{y}}}\nabla(F^{\langle\mu\rangle}\circ\bm{\varphi})(\bm{y})\subset\partial_{\rm L}(F\circ\bm{\varphi})(\widebar{\bm{y}}).$ (37)
    Moreover, if a given sequence $(\bm{y}_{n})_{n=1}^{\infty}\subset\mathcal{Y}$ converges to some $\widebar{\bm{y}}\in\mathcal{Y}$, and $(\mu_{n})_{n=1}^{\infty}\searrow 0$ with $\mu_{n}\in(0,\eta^{-1})$, then
    $d(\bm{0},\partial_{\rm L}(F\circ\bm{\varphi})(\widebar{\bm{y}}))\leq\liminf_{n\to\infty}\left\lVert\nabla(F^{\langle\mu_{n}\rangle}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert.$ (38)
Proof.
See Appendix D. ∎
Remark 4.5 (Gradient (sub-)consistency).
  1. (a)
    Although the converse inclusion of (36) is not required in this paper, the equality
    $(\widebar{\bm{y}}\in\mathcal{Y})\quad\mathop{\mathrm{Lim~sup}}_{\mathbb{R}_{++}\ni\mu\searrow 0,\ \mathcal{Y}\ni\bm{y}\to\widebar{\bm{y}}}\nabla({}^{\mu}J\circ\mathcal{F})(\bm{y})=\partial_{\rm L}(J\circ\mathcal{F})(\widebar{\bm{y}}),$ (39)
    called the gradient consistency property [26] of $^{\mu}J\circ\mathcal{F}$ for $J\circ\mathcal{F}$, can also be shown (i) by checking that $^{\mu}J\circ\mathcal{F}$ epi-converges to $J\circ\mathcal{F}$ (see [69, Def. 7.1] for the definition of epi-convergence); and (ii) by applying [18, Lemma 3.4] to $^{\mu}J\circ\mathcal{F}$. Indeed, (i) can be verified by [69, Thm. 7.4(d)] and $^{\mu}J\circ\mathcal{F}\nearrow J\circ\mathcal{F}$ as $\mu\searrow 0$.
  2. (b)
    In a special case where $J$ is convex, we can check that $^{\mu}J\circ\mathcal{F}$ in Theorem 4.4 (a) has the gradient consistency property for $J\circ\mathcal{F}$ by using [18, Cor. 5.8]. However, the discussion in the proof of [18, Cor. 5.8] cannot apply to our case where $J$ is just weakly convex because the proof of [18, Cor. 5.8] relies on the structure of the convexity of $J$ (see Attouch’s theorem [69, Thm. 12.35]). In order to show the gradient consistency of $^{\mu}J\circ\mathcal{F}$ even with $J$ weakly convex, we take a different approach in the proof of Theorem 4.4. We note that our smoothed surrogate function $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ with a linear $\mathfrak{S}$ and $\bm{\varphi}\coloneqq\mathrm{Id}$ has been employed in the existing variable smoothing algorithms [13, 10, 66], although the gradient consistency property has not been discussed therein.
  3. (c)
    We can construct another smoothed surrogate function, with gradient consistency property, of $F\circ\bm{\varphi}$ via the integral convolution of $F\circ\bm{\varphi}$ and a kernel function, called mollifier (see [69, Exm. 7.19, Thm. 9.67] or [19, Thm. 4.7]), because $F\circ\bm{\varphi}$ is subdifferentially regular (see Lemma 2.4). However, it is questionable to obtain closed-form expressions of this type of smoothed surrogate functions for $F\circ\bm{\varphi}$ and its gradient, except for simple functions, e.g., the plus function $\mathbb{R}\ni x\mapsto\max\{x,0\}$ [22]. In contrast, the closed-form expressions of our smoothed surrogate function $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ and its gradient are available as computable tools in a case where $g$ is prox-friendly (see (28)).
4.2 Variable Smoothing Algorithm with Convergence Analysis
The inequality (38) in Theorem 4.4 (b) implies that finding a stationary point $\bm{y}^{\star}\in\mathcal{Y}$ of Problem 1.4 satisfying (27) can be reduced to the following problem:
$\begin{split}&{\rm find\ a\ convergent\ sequence}\ (\bm{y}_{n})_{n=1}^{\infty}\subset\mathcal{Y}\ {\rm such\ that}\\ &\quad\liminf_{n\to\infty}\left\lVert\nabla((h+{}^{\mu_{n}}g\circ\mathfrak{S})\circ\bm{\varphi})(\bm{y}_{n})\right\rVert=0\ {\rm with\ }(\mu_{n})_{n=1}^{\infty}(\subset(0,\eta^{-1}))\searrow 0.\end{split}$ (40)
In this section, for finding a stationary point of Problem 1.4 via the problem in (40), we present a variable smoothing algorithm in Algorithm 1 as an extension of the variable smoothing algorithm [13] that has been proposed for Problem 1.4 with $\bm{\varphi}=\mathrm{Id}$ and a linear operator $\mathfrak{S}$. We note that the algorithm [13] is not applicable to Problem 1.4 due to the nonlinearity of $\mathfrak{S}$ even if $\bm{\varphi}=\mathrm{Id}$. The proposed algorithm is designed with a time-varying gradient descent update for minimization of
$F_{n}\circ\bm{\varphi}\coloneqq(h+{}^{\mu_{n}}g\circ\mathfrak{S})\circ\bm{\varphi}\ \mathrm{with}\ F_{n}:\mathcal{X}\to\mathbb{R}:\widebar{\bm{x}}\mapsto(h+{}^{\mu_{n}}g\circ\mathfrak{S})(\widebar{\bm{x}})$ (41)
at $n$th update from $\bm{y}_{n}\in\mathcal{Y}$ to $\bm{y}_{n+1}\coloneqq\bm{y}_{n}-\gamma_{n}\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\in\mathcal{Y}$ with $\mu_{n}\searrow 0$ and a stepsize $\gamma_{n}>0$. In a case where $\mathrm{prox}_{\mu_{n}g}$ has a closed-form expression, the values of $F_{n}\circ\bm{\varphi}(\widebar{\bm{y}})$ and $\nabla(F_{n}\circ\bm{\varphi})(\widebar{\bm{y}})$ are available for every $\widebar{\bm{y}}\in\mathcal{Y}$ (see (28)).
The index $\mu_{n}\in(0,(2\eta)^{-1}]$ in Algorithm 1 is designed to establish a convergence analysis of Algorithm 1 as
$(\mu_{n})_{n=1}^{\infty}\subset(0,(2\eta)^{-1}]\ \mathrm{satisfies}\ \begin{cases}&{\rm(i)}\ \lim_{n\to\infty}\mu_{n}=0,{\rm(ii)}\ \sum_{n=1}^{\infty}\mu_{n}=+\infty,\\ &{\rm(iii)}\ \mu_{n+1}\leq\mu_{n}\ (n\in\mathbb{N}),\end{cases}$ (42)
where $\eta>0$ is the level of the weak convexity of $g$ (see Problem 1.1). We have many choices of $(\mu_{n})_{n=1}^{\infty}$ enjoying (42) as shown in Example 4.6.
Algorithm 1 Variable smoothing for Problem 1.4 (VSmooth)
[…]
3:while stopping criterion in not satisfied do
4:  Set $F_{n}\coloneqq(h+{}^{\mu_{n}}g\circ\mathfrak{S})$
5:  Set $\gamma_{n}>0$ satisfying (49) with $c$ (see Example 4.8)
6:  $\bm{y}_{n+1}\leftarrow\bm{y}_{n}-\gamma_{n}\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})$ (see (28) for the expression of $\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})$)
7:  $n\leftarrow n+1$
8:end while
[…]
Example 4.6 ($(\mu_{n})_{n=1}^{\infty}$ satisfying (42)).
  1. (a)
    For any $\alpha\geq 1$, $(\mu_{n})_{n=1}^{\infty}\coloneqq((2\eta)^{-1}n^{-1/\alpha})_{n=1}^{\infty}$ satisfies (42).
  2. (b)
    $(\mu_{n})_{n=1}^{\infty}\coloneqq\left(\frac{1}{2\eta(n+1)\log(n+1)}\right)_{n=1}^{\infty}$ satisfies (42).
Proof.
(a) Let $\alpha\geq 1$ be chosen arbitrarily. Clearly, $(\mu_{n})_{n=1}^{\infty}\subset(0,(2\eta)^{-1}]$ satisfies the conditions (i) and (iii) in (42). The condition (ii) can be checked by
$\sum_{n=1}^{K}n^{-1/\alpha}=\sum_{n=1}^{K}\int_{n}^{n+1}n^{-1/\alpha}dx\geq\int_{1}^{K+1}x^{-1/\alpha}dx=\begin{cases}\frac{(K+1)^{1-\alpha^{-1}}-1}{1-\alpha^{-1}}&(\alpha>1)\\ \log(K+1)&(\alpha=1)\end{cases}\overset{K\to\infty}{\to}+\infty.$ (43)
(b) Clearly, $(\mu_{n})_{n=1}^{\infty}\subset(0,(2\eta)^{-1}]$ satisfies the conditions (i) and (iii). The condition (ii) can be checked by
$\sum_{n=1}^{K}\frac{1}{(n+1)\log(n+1)}=\sum_{n=1}^{K}\int_{n}^{n+1}\frac{dx}{(n+1)\log(n+1)}\geq\int_{1}^{K+1}\frac{dx}{(x+1)\log(x+1)}$ (44)
$\overset{u=\log(x+1)}{=}\int_{\log(2)}^{\log(K+2)}\frac{du}{u}=\log(\log(K+2))-\log(\log(2))\overset{K\to\infty}{\to}+\infty.$ (45)
Every stepsize $\gamma_{n}>0$ in Algorithm 1 is chosen to enjoy the so-called Armijo condition (see, e.g., [5, Thm. 1.1]) with a predetermined constant $c\in(0,1)$:
$F_{n}\circ\bm{\varphi}(\bm{y}_{n}-\gamma_{n}\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n}))\leq F_{n}\circ\bm{\varphi}(\bm{y}_{n})-c\gamma_{n}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert^{2}.$ (46)
In a case where $\gamma_{n}$ satisfies the Armijo condition (46), the descent condition $F_{n}\circ\bm{\varphi}(\bm{y}_{n}-\gamma_{n}\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n}))<F_{n}\circ\bm{\varphi}(\bm{y}_{n})$ is achieved if $\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\neq\bm{0}$, and thus we update $\bm{y}_{n+1}\coloneqq\bm{y}_{n}-\gamma_{n}\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})$.
To use $\gamma_{n}$ satisfying (46) in Algorithm 1, we assume Assumption 4.7 below. Sufficient conditions for Assumption 4.7 (a) are found in Proposition 4.2 and Remark 4.3. Moreover, Example 4.8 presents typical choices of $\gamma_{n}$ achieving Assumption 4.7 (b) under Assumption 4.7 (a).
Assumption 4.7.
Consider Problem 1.4 and Alg. 1. For $F_{n}=h+{}^{\mu_{n}}g\circ\mathfrak{S}$, we assume:
  1. (a)
    (Gradient Lipschitz continuity condition)
    $\hskip-18.49988pt(\exists\varpi_{1}\in\mathbb{R}_{+},\exists\varpi_{2}\in\mathbb{R}_{++},\forall n\in\mathbb{N})$ (47)
    $\hskip-18.49988pt\nabla(F_{n}\circ\bm{\varphi})~\mathrm{is~Lipschitz~continuous}\mathrm{~with~a~Lipschitz~constant~}L_{\nabla(F_{n}\circ\bm{\varphi})}\coloneqq\varpi_{1}+\frac{\varpi_{2}}{\mu_{n}}.$ (48)
    (Note: under the condition (48), we have $L_{\nabla(F_{n+1}\circ\bm{\varphi})}\geq L_{\nabla(F_{n}\circ\bm{\varphi})}\geq\cdots\geq L_{\nabla(F_{1}\circ\bm{\varphi})}$ by $\mu_{n}\searrow 0$ [see the condition (iii) in (42)] for $(\mu_{n})_{n=1}^{\infty}$)
  2. (b)
    (Lower bound condition for stepsizes)
[…]
Example 4.8 (Stepsize achieving Assumption 4.7 (b)).
Under Assumption 4.7 (a), we have mainly two choices of $\gamma_{n}$ achieving (49) (see, e.g., [5, Thm. 1.1]):
  1. (a)
    With the Lipschitz constant $L_{\nabla(F_{n}\circ\bm{\varphi})}\in\mathbb{R}_{++}$,
    $(n\in\mathbb{N})\quad\gamma_{n}\coloneqq\beta L_{\nabla(F_{n}\circ\bm{\varphi})}^{-1}\coloneqq 2(1-c)L_{\nabla(F_{n}\circ\bm{\varphi})}^{-1}\in\mathbb{R}_{++}$ (50)
    satisfies the Armijo condition (46) with $c\in(0,1)$, and thus (49).
  2. (b)
    The so-called backtracking algorithm in Algorithm 2 returns $\gamma_{n}$ enjoying the Armijo condition (46) with $c\in(0,1)$, where $\gamma_{\rm init}\in\mathbb{R}_{++}$ is an initial guess for $\gamma_{n}$, and $\rho\in(0,1)$. In this case, $\gamma_{n}$ is known to have the following lower bound:
    $(n\in\mathbb{N})\quad\gamma_{n}\geq\min\left\{\gamma_{\rm init}L_{\nabla(F_{n}\circ\bm{\varphi})},2\rho(1-c)\right\}L_{\nabla(F_{n}\circ\bm{\varphi})}^{-1}\geq\beta L_{\nabla(F_{n}\circ\bm{\varphi})}^{-1}$ (51)
    with $\beta\coloneqq\min\left\{\gamma_{\rm init}L_{\nabla(F_{1}\circ\bm{\varphi})},2\rho(1-c)\right\}\in\mathbb{R}_{++}$, where the first inequality follows by [5, Thm. 1.1], and the second inequality follows [^10] The derivation of the second inequality is inspired by [10, the equality just after (4.12)]. by $L_{\nabla(F_{n}\circ\bm{\varphi})}\geq L_{\nabla(F_{1}\circ\bm{\varphi})}$ (see the sentence just after (48)). From (51), Algorithm 2 is guaranteed to terminate in at most $\max\left\{1,\left\lceil\log_{\rho}\left(\frac{2(1-c)}{\gamma_{\rm init}L_{\nabla(F_{n}\circ\bm{\varphi})}}\right)\right\rceil\right\}=\max\left\{1,\left\lceil\log_{\rho}\left(\frac{2\mu_{n}(1-c)}{\gamma_{\rm init}(\mu_{n}\varpi_{1}+\varpi_{2})}\right)\right\rceil\right\}$ updates, and $\gamma_{n}$ given by Algorithm 2 satisfies (49), where $\lceil\cdot\rceil$ stands for the ceiling function. We note that any knowledge on $L_{\nabla(F_{n}\circ\bm{\varphi})}=\varpi_{1}+\varpi_{2}\mu_{n}^{-1}$ with $\varpi_{1},\varpi_{2}\in\mathbb{R}_{+}$ is not required in Algorithm 2.
Algorithm 2 Backtracking algorithm to find $\gamma_{n}$ satisfying (46)
$c\in(0,1),\rho\in(0,1),\ \gamma_{\rm init}\in\mathbb{R}_{++}$
$\gamma_{n}\leftarrow\gamma_{\rm init}$
$f_{n}\leftarrow F_{n}\circ\bm{\varphi}(\bm{y}_{n})\in\mathbb{R}$, $\bm{d}_{n}\leftarrow\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\in\mathcal{Y}$
while $F_{n}\circ\bm{\varphi}(\bm{y}_{n}-\gamma_{n}\bm{d}_{n})>f_{n}-c\gamma_{n}\left\lVert\bm{d}_{n}\right\rVert^{2}$ do
  $\gamma_{n}\leftarrow\rho\gamma_{n}$
end while
$\gamma_{n}$
We present below a convergence analysis of Algorithm 1 under Assumption 4.7.
Theorem 4.9 (Asymptotic convergence analysis of Algorithm 1).
Consider Problem 1.4. Choose arbitrarily $\bm{y}_{1}\in\mathcal{Y}$, $c\in(0,1)$, and $(\mu_{n})_{n=1}^{\infty}\subset(0,(2\eta)^{-1}]$ satisfying (42). Suppose that $(\bm{y}_{n})_{n=1}^{\infty}\subset\mathcal{Y}$ is generated by Algorithm 1 under Assumption 4.7 with $F_{n}\coloneqq F^{\langle\mu_{n}\rangle}\coloneqq h+{}^{\mu_{n}}g\circ\mathfrak{S}\ (n\in\mathbb{N})$. Then, $F_{n}\circ\bm{\varphi}\ (n\in\mathbb{N})$ and $(\bm{y}_{n})_{n=1}^{\infty}$ enjoy the following:
  1. (a)
    For any pair $(\underline{k},\overline{k})\in\mathbb{N}^{2}$ such that $\underline{k}\leq\overline{k}$, we have
    $\min_{\underline{k}\leq n\leq\overline{k}}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert\leq\sqrt{\frac{\chi}{\sum_{n=\underline{k}}^{\overline{k}}\mu_{n}}}.$ (52)
    with $\chi\coloneqq\frac{(F_{1}\circ\bm{\varphi}(\bm{y}_{1})-\inf_{\bm{y}\in\mathcal{Y}}F_{1}\circ\bm{\varphi}(\bm{y})+L_{g}^{2}\mu_{1})(\varpi_{1}+\eta\varpi_{2})}{c\beta\eta}\in\mathbb{R}_{++}$, where (I) $L_{g},\eta\in\mathbb{R}_{++}$ are respectively the Lipschitz constant and the level of the weak convexity of $g$ (see the condition (iii) in Problem 1.1), and (II) $\varpi_{1}\in\mathbb{R}_{+}$ and $\varpi_{2},\beta\in\mathbb{R}_{++}$ are given in Assumption 4.7.
  2. (b)
    $\liminf_{n\to\infty}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert=0.$ (53)
  3. (c)
    Choose a subsequence $(\bm{y}_{m(l)})_{l=1}^{\infty}\subset\mathcal{Y}$ of $(\bm{y}_{n})_{n=1}^{\infty}$ such that $\lim_{l\to\infty}\left\lVert\nabla(F_{m(l)}\circ\bm{\varphi})(\bm{y}_{m(l)})\right\rVert=0$, where $m:\mathbb{N}\to\mathbb{N}$ is monotonically increasing [^11] From Theorem 4.9 (b), we can construct such a subsequence $(\bm{y}_{m(l)})_{l=1}^{\infty}$ by using, e.g., $m(l)\coloneqq\begin{cases}1&(l=1)\\ \min\{n\in\mathbb{N}\mid n>m(l-1),\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert\leq 2^{-l}\}&(l>1).\end{cases}$ (54) . Then, every cluster point $\bm{y}^{\star}\in\mathcal{Y}$ of $(\bm{y}_{m(l)})_{l=1}^{\infty}$ is a stationary point of Problem 1.4. Moreover, if $N_{C}(\bm{\varphi}(\bm{y}^{\star}))\supset\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*})$ holds, then $\bm{\varphi}(\bm{y}^{\star})\in C$ is a stationary point of Problem 1.1 by Theorem 3.2 (b) (see also Corollary 3.4 and Remark 3.6).
  4. (d)
    Consider $\mu_{n}\coloneqq\tau n^{-1/\alpha}\ (n\in\mathbb{N})$ with $\tau\in(0,(2\eta)^{-1}]$ and $\alpha>1$. Then, $(\mu_{n})_{n=1}^{\infty}$ satisfies (42) by Example 4.6 (a), and, for a given tolerance $\epsilon\in(0,1)$, the conditions $\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert<\epsilon$ and $\mu_{n}<\epsilon$ are jointly achieved within at most
    $\left\lceil\left(\chi\tau^{-1}(1-\alpha^{-1})+\tau^{\alpha-1}+2\right)^{\frac{\alpha}{\alpha-1}}\epsilon^{\min\left\{\frac{-2\alpha}{\alpha-1},-\alpha\right\}}\right\rceil\ \mathrm{iterations},$ (55)
    where $\chi\in\mathbb{R}_{++}$ is given in Theorem 4.9 (a), $\lceil\cdot\rceil$ stands for the ceiling function, and such $\bm{y}_{n}$ is an $\epsilon$-approximate stationary point defined in Definition 4.10 (b).
Proof.
The proof for (a) is inspired partially by that for [13, Thm. 4.1].
(a) Since $\gamma_{n}>0$ satisfies the Armijo condition (46) by Assumption 4.7 (b), we have
$(n\in\mathbb{N})\quad F_{n}\circ\bm{\varphi}(\bm{y}_{n+1})$ $=F_{n}\circ\bm{\varphi}(\bm{y}_{n}-\gamma_{n}\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n}))$ (56)
$\leq F_{n}\circ\bm{\varphi}(\bm{y}_{n})-c\gamma_{n}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert^{2}.$ (57)
By (31) in Lemma 4.1 with $0<\mu_{n+1}\leq\mu_{n}<\eta^{-1}$, we have
[…]
By combining (57) and (58), we obtain
[…]
and thus
$(n\in\mathbb{N})\quad F_{n+1}\circ\bm{\varphi}(\bm{y}_{n+1})\leq F_{n}\circ\bm{\varphi}(\bm{y}_{n})+L_{g}^{2}(\mu_{n}-\mu_{n+1}).$ (61)
By summing (60) up from $n=\underline{k}$ to $\overline{k}$, the following inequality holds:
$c\sum_{n=\underline{k}}^{\overline{k}}\gamma_{n}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert^{2}$ $\leq F_{\underline{k}}\circ\bm{\varphi}(\bm{y}_{\underline{k}})-F_{\overline{k}+1}\circ\bm{\varphi}(\bm{y}_{\overline{k}+1})+L_{g}^{2}(\mu_{\underline{k}}-\mu_{\overline{k}+1})$ (62)
$\leq F_{\underline{k}}\circ\bm{\varphi}(\bm{y}_{\underline{k}})-\inf_{\bm{y}\in\mathcal{Y}}F_{1}\circ\bm{\varphi}(\bm{y})+L_{g}^{2}(\mu_{\underline{k}}-\mu_{\overline{k}+1})$ (63)
$\leq F_{1}\circ\bm{\varphi}(\bm{y}_{1})-\inf_{\bm{y}\in\mathcal{Y}}F_{1}\circ\bm{\varphi}(\bm{y})+L_{g}^{2}(\mu_{1}-\mu_{\overline{k}+1})$ (64)
[…]
By Assumption 4.7, recall that (i) $\nabla(F_{n}\circ\bm{\varphi})$ is Lipschitz continuous with a Lipschitz constant $L_{\nabla(F_{n}\circ\bm{\varphi})}=\varpi_{1}+\varpi_{2}\mu_{n}^{-1}$ with some constants $\varpi_{1}\geq 0,\varpi_{2}>0$; (ii) every $\gamma_{n}$ is bounded below by $\beta L_{\nabla(F_{n}\circ\bm{\varphi})}^{-1}$ with some $\beta\in\mathbb{R}_{++}$. We have
$(n\in\mathbb{N})\quad\gamma_{n}\geq\beta L_{\nabla(F_{n}\circ\bm{\varphi})}^{-1}=\frac{\beta\mu_{n}}{\varpi_{1}\mu_{n}+\varpi_{2}}=\frac{\beta\eta\mu_{n}}{\varpi_{1}\eta\mu_{n}+\eta\varpi_{2}}\geq\frac{\beta\eta}{\varpi_{1}+\eta\varpi_{2}}\mu_{n},$ (66)
where the last inequality follows by $\eta\mu_{n}\in(0,2^{-1}]$. Then, the LHS in (65) can be further bounded below as
$c\sum_{n=\underline{k}}^{\overline{k}}\gamma_{n}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert^{2}$ $\geq\frac{c\beta\eta}{\varpi_{1}+\eta\varpi_{2}}\sum_{n=\underline{k}}^{\overline{k}}\mu_{n}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert^{2}$ (67)
$\geq\frac{c\beta\eta}{\varpi_{1}+\eta\varpi_{2}}\min_{\underline{k}\leq n\leq\overline{k}}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert^{2}\sum_{n=\underline{k}}^{\overline{k}}\mu_{n}.$ (68)
By (65) and (68), we get the desired inequality in (52).
(b) Let $a_{n}\coloneqq\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert$, and $b_{n}\coloneqq\inf\{a_{m}\mid m\geq n\}$. For every $\underline{k}\in\mathbb{N}$, the inequality (52) with $\overline{k}\to\infty$ implies $b_{\underline{k}}=0$ thanks to the nonsummability of $(\mu_{n})_{n=1}^{\infty}$ (see the condition (ii) in (42)). This implies $\liminf_{n\to\infty}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert=\liminf_{n\to\infty}a_{n}=\lim_{n\to\infty}b_{n}=0$.
(c) Let $\bm{y}^{\star}\in\mathcal{Y}$ be a cluster point of $(\bm{y}_{m(l)})_{l=1}^{\infty}$. By (38) and Theorem 4.9 (b), we have $d(\bm{0},\partial_{\rm L}(F\circ\bm{\varphi})(\bm{y}^{\star}))=0$ with $F=h+g\circ\mathfrak{S}$, implying $\bm{y}^{\star}$ satisfies (26).
(d) By $\mu_{n}=\tau n^{-1/\alpha}$, we have $\mu_{n}<\epsilon$ for all $n\geq\underline{k}\coloneqq\lceil\tau^{\alpha}\epsilon^{-\alpha}\rceil+1$. We derive $\overline{k}\in\mathbb{N}$ below such that $\overline{k}\geq\underline{k}$ and $\min_{\underline{k}\leq n\leq\overline{k}}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert<\epsilon$. By (52), we have
$\min_{\underline{k}\leq n\leq\overline{k}}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert\leq\sqrt{\frac{\chi}{\tau\sum_{n=\underline{k}}^{\overline{k}}n^{-1/\alpha}}}\leq\sqrt{\frac{\chi(1-\alpha^{-1})}{\tau\left((\overline{k}+1)^{1-\alpha^{-1}}-\underline{k}^{1-\alpha^{-1}}\right)}},$ (69)
where we used $\sum_{n=\underline{k}}^{\overline{k}}n^{-1/\alpha}\geq\frac{(\overline{k}+1)^{1-\alpha^{-1}}-\underline{k}^{1-\alpha^{-1}}}{1-\alpha^{-1}}$ by a discussion similar to (43). Hence, $\min_{\underline{k}\leq n\leq\overline{k}}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert<\epsilon$ holds if $\overline{k}$ satisfies $\sqrt{\frac{\chi(1-\alpha^{-1})}{\tau\left((\overline{k}+1)^{1-\alpha^{-1}}-\underline{k}^{1-\alpha^{-1}}\right)}}<\epsilon$, or equivalently
$\overline{k}>\left(\chi\tau^{-1}(1-\alpha^{-1})\epsilon^{-2}+\underline{k}^{1-\alpha^{-1}}\right)^{\frac{\alpha}{\alpha-1}}-1.$ (70)
[…]
$\left(\chi\tau^{-1}(1-\alpha^{-1})\epsilon^{-2}+\underline{k}^{1-\alpha^{-1}}\right)^{\frac{\alpha}{\alpha-1}}\leq\left(\left(\chi\tau^{-1}(1-\alpha^{-1})+\tau^{\alpha-1}+2\right)\epsilon^{\min\{-2,1-\alpha\}}\right)^{\frac{\alpha}{\alpha-1}}$ (71)
$\leq\left\lceil\left(\chi\tau^{-1}(1-\alpha^{-1})+\tau^{\alpha-1}+2\right)^{\frac{\alpha}{\alpha-1}}\epsilon^{\min\left\{\frac{-2\alpha}{\alpha-1},-\alpha\right\}}\right\rceil\eqqcolon n_{0}.$ (72)
Hence, (70) is achieved by $\overline{k}\coloneqq n_{0}$, and thus we get $\min_{\underline{k}\leq n\leq n_{0}}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert<\epsilon$, i.e., there exists $\widehat{n}\in[\underline{k},n_{0}]\cap\mathbb{N}$ such that $\left\lVert\nabla(F_{\widehat{n}}\circ\bm{\varphi})(\bm{y}_{\widehat{n}})\right\rVert<\epsilon$. Since $\mu_{n}<\epsilon$ holds for all $n\geq\underline{k}$, the conditions $\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert<\epsilon$ and $\mu_{n}<\epsilon$ are jointly achieved within at most $n_{0}$ iterations. ∎
To derive a first-order oracle complexity of Algorithm 1, we introduce the definitions of the first-order oracle and an $\epsilon$-approximate stationary point, e.g., [37, 12, 88], for smoothing-type algorithms.
Definition 4.10.
Consider Problem 1.4 with $F^{\langle\mu\rangle}=h+{}^{\mu}g\circ\mathfrak{S}$.
  1. (a)
    For given $\bm{x},\bm{u}\in\mathcal{X}$, $\bm{y}\in\mathcal{Y}$, and $\bm{z},\bm{v}\in\mathcal{Z}$, and $\mu\in(0,\eta^{-1})$, the first-order oracle returns $h(\bm{x})$, $\nabla h(\bm{x})$, $\mathfrak{S}(\bm{x})$, $(\mathrm{D}\mathfrak{S}(\bm{x}))^{*}[\bm{v}]$, $\bm{\varphi}(\bm{y})$ $(\mathrm{D}\bm{\varphi}(\bm{y}))^{*}[\bm{u}]$, $g(\bm{z})$, and $\mathrm{prox}_{\mu g}(\bm{z})$ (Note: one evaluation of $F^{\langle\widebar{\mu}\rangle}\circ\bm{\varphi}(\widebar{\bm{y}})$ and $\nabla(F^{\langle\widebar{\mu}\rangle}\circ\bm{\varphi})(\widebar{\bm{y}})$ for $\widebar{\bm{y}}\in\mathcal{Y}$ and $\widebar{\mu}\in(0,\eta^{-1})$ can be implemented with at most five oracle calls [^13] By noting that * denotes an arbitrary value, consider the following five oracle calls with (i) $(\bm{x},\bm{u},\bm{y},\bm{z},\bm{v},\mu)=(*,*,\widebar{\bm{y}},*,*,*)$ to obtain $\widebar{\bm{x}}\coloneqq\bm{\varphi}(\widebar{\bm{y}})\in\mathcal{X}$; (ii) $(\bm{x},\bm{u},\bm{y},\bm{z},\bm{v},\mu)=(\widebar{\bm{x}},*,*,*,*,*)$ to obtain $\widebar{\xi}_{\rm cost}\coloneqq h(\widebar{\bm{x}})\in\mathbb{R}$, $\widebar{\bm{\xi}}_{\rm grad}\coloneqq\nabla h(\widebar{\bm{x}})\in\mathcal{X}$, $\widebar{\bm{z}}\coloneqq\mathfrak{S}(\widebar{\bm{x}})\in\mathcal{Z}$; (iii) $(\bm{x},\bm{u},\bm{y},\bm{z},\bm{v},\mu)=(*,*,*,\widebar{\bm{z}},*,\widebar{\mu})$ to obtain $\widebar{\bm{p}}\coloneqq\mathrm{prox}_{\widebar{\mu}g}(\widebar{\bm{z}})\in\mathcal{Z}$; (iv) $(\bm{x},\bm{u},\bm{y},\bm{z},\bm{v},\mu)=(\widebar{\bm{x}},*,*,\widebar{\bm{p}},(\widebar{\bm{z}}-\widebar{\bm{p}})/\widebar{\mu},*)$ to obtain $\widebar{\zeta}_{\rm cost}\coloneqq g(\widebar{\bm{p}})\in\mathbb{R}$, $\widebar{\bm{\zeta}}_{\rm grad}\coloneqq\left(\mathrm{D}\mathfrak{S}(\widebar{\bm{x}})\right)^{*}[(\widebar{\bm{z}}-\widebar{\bm{p}})/\widebar{\mu}]\in\mathcal{X}$; (v) $(\bm{x},\bm{u},\bm{y},\bm{z},\bm{v},\mu)=(*,\widebar{\bm{\xi}}_{\rm grad}+\widebar{\bm{\zeta}}_{\rm grad},\widebar{\bm{y}},*,*,*)$ to obtain $\nabla(F^{\langle\widebar{\mu}\rangle}\circ\bm{\varphi})(\widebar{\bm{y}})=(\mathrm{D}\bm{\varphi}(\widebar{\bm{y}}))^{*}[\widebar{\bm{\xi}}_{\rm grad}+\widebar{\bm{\zeta}}_{\rm grad}]\in\mathcal{Y}$, where we can compute $F^{\langle\widebar{\mu}\rangle}\circ\bm{\varphi}(\widebar{\bm{y}})=\widebar{\xi}_{\rm cost}+\widebar{\zeta}_{\rm cost}+\frac{1}{2\widebar{\mu}}\left\lVert\widebar{\bm{p}}-\widebar{\bm{z}}\right\rVert^{2}\in\mathbb{R}$ (Compare with the expression (28)). ).
  2. (b)
    For a given $\epsilon>0$, a point $\widehat{\bm{y}}\in\mathcal{Y}$ is said to be an $\epsilon$-approximate stationary point of Problem 1.4 if $\left\lVert\nabla(F^{\langle\mu\rangle}\circ\bm{\varphi})(\widehat{\bm{y}})\right\rVert\leq\epsilon$ holds for some $\mu\in(0,\epsilon]$
Compared with the oracle in [83] proposed for Problem 1.1 with $g$ convex, our first-order oracle returns additionally the values $\bm{\varphi}(\bm{y})$ and its derivative $(\mathrm{D}\bm{\varphi}(\bm{y}))^{*}[\bm{u}]$ for $\bm{y}\in\mathcal{Y}$ and $\bm{u}\in\mathcal{X}$, where an evaluation of $\bm{\varphi}$ corresponds to that of a retraction used in [83] for mapping a tangent vector to a point in the manifold $C$ (see just after (90) for a retraction). Moreover, our oracle is essentially the same one considered in [13] for Problem 1.4 in a special case where $\bm{\varphi}=\mathrm{Id}$ and $\mathfrak{S}$ is linear.
Our approximate stationarity notion for Problem 1.4 in Definition 4.10 (b) is not directly comparable to the notions used in, [66, 10, 83] and [25, 80], for Problem 1.1. Instead, our approximate stationarity reduces to another notion of approximate stationarity used in [13] for Problem 1.4 with $\bm{\varphi}=\mathrm{Id}$ and $\mathfrak{S}$ linear (see Remark 4.11).
Remark 4.11 (Relation between approximate stationarities in Definition 4.10 (b) and in [13]).
For Problem 1.4, another type of an $\epsilon$-approximate stationary point can be defined as a point $\widehat{\bm{y}}\in\mathcal{Y}$ satisfying
$\exists\bm{z}\in\mathcal{Z}\ \mathrm{s.t.}\ d\left(-\nabla(h\circ\bm{\varphi})(\widehat{\bm{y}}),(\mathrm{D}(\mathfrak{S}\circ\bm{\varphi})(\widehat{\bm{y}}))^{*}\left[\partial_{\rm L}g(\bm{z})\right]\right)\leq\epsilon\ \mathrm{and}\ \left\lVert\bm{z}-\mathfrak{S}\circ\bm{\varphi}(\widehat{\bm{y}})\right\rVert\leq\epsilon.$ (73)
In a special case where $\bm{\varphi}\coloneqq\mathrm{Id}$ and $\mathfrak{S}$ is linear, a variable smoothing algorithm in [13] can find an $\epsilon$-approximate stationary point in the sense of (73) within $\mathcal{O}(\epsilon^{-3})$ oracle calls. In fact, our approximate stationary point in Definition 4.10 (b) implies an approximate stationary point in the sense of (73) as follows. Let $\widehat{\bm{y}}\in\mathcal{Y}$ be an $\epsilon$-approximate stationary point in Definition 4.10 (b), i.e., $\left\lVert\nabla(F^{\langle\mu\rangle}\circ\bm{\varphi})(\widehat{\bm{y}})\right\rVert\leq\epsilon$ with some $\mu\in(0,\epsilon]$. Then, we have
[…]
$=$ $\left\lVert\nabla(h\circ\bm{\varphi})(\widehat{\bm{y}})+(\mathrm{D}(\mathfrak{S}\circ\bm{\varphi})(\widehat{\bm{y}}))^{*}[\nabla{}^{\mu}g(\mathfrak{S}\circ\bm{\varphi}(\widehat{\bm{y}}))]\right\rVert=\left\lVert\nabla(F^{\langle\mu\rangle}\circ\bm{\varphi})(\widehat{\bm{y}})\right\rVert\leq\epsilon.$ (76)
Moreover, with the Lipschitz constant $L_{g}$ of $g$, $\left\lVert\mathrm{prox}_{\mu g}\left(\mathfrak{S}\circ\bm{\varphi}(\widehat{\bm{y}})\right)-\mathfrak{S}\circ\bm{\varphi}(\widehat{\bm{y}})\right\rVert\leq L_{g}\mu$ holds from $\left\lVert\mathrm{prox}_{\mu g}(\widebar{\bm{z}})-\widebar{\bm{z}}\right\rVert\leq L_{g}\mu\ (\widebar{\bm{z}}\in\mathcal{Z})$ (see Fact 2.7 (b) and (22)). Hence, $\widehat{\bm{y}}$ is a $(\max\{1,L_{g}\}\epsilon)$-approximate stationary point in the sense of (73).
In Corollary 4.12, we derive the first-order complexity, i.e., the number of oracle calls, $\mathcal{O}(\epsilon^{-3})$ of Algorithm 1 for finding an $\epsilon$-approximate stationary point of Problem 1.4. We note that, under a special case where $\bm{\varphi}=\mathrm{Id}$ and $\mathfrak{S}$ is linear, Corollary 4.12 recovers the oracle complexity $\mathcal{O}(\epsilon^{-3})$ of the existing variable smoothing algorithm [13] (see Remark 4.11 and Table 1 in Section 1), whereas the proposed algorithm admits the nonlinearity of $\mathfrak{S}$.
Corollary 4.12 (First-order oracle complexity of Algorithm 1).
Consider Problem 1.4. Choose arbitrarily $\bm{y}_{1}\in\mathcal{Y}$, $c\in(0,1)$, and $\mu_{n}\coloneqq\tau n^{-1/3}\ (n\in\mathbb{N})$ with $\tau\in(0,(2\eta)^{-1}]$. Then, under Assumption 4.7, for any given $\epsilon>0$, Algorithm 1 can find an $\epsilon$-approximate stationary point of Problem 1.4 (i) with at most $\mathcal{O}(\epsilon^{-3})$ first-order oracle calls if $\gamma_{n}\coloneqq 2(1-c)L_{\nabla(F_{n}\circ\bm{\varphi})}^{-1}\ (n\in\mathbb{N})$ (see (48) and Example 4.8 (a)), and (ii) with at most $\mathcal{O}(\epsilon^{-3}\log(\epsilon^{-1}))$ first-order oracle calls if $\gamma_{n}$ is given by Algorithm 2 with $\rho\in(0,1)$ and $\gamma_{\rm init}\in\mathbb{R}_{++}$ (see Example 4.8 (b)).
Proof.
Recall that Algorithm 1 can find an $\epsilon$-approximate stationary point within $\mathcal{O}(\epsilon^{-3})$ iterations by Theorem 4.9 (d) with $\alpha\coloneqq 3$. Hence, for the case (i) $\gamma_{n}\coloneqq 2(1-c)L_{\nabla(F_{n}\circ\bm{\varphi})}^{-1}\ (n\in\mathbb{N})$, Algorithm 1 can find an $\epsilon$-approximate stationary point with at most $\mathcal{O}(\epsilon^{-3})$ first-order oracle calls because $F_{n}\circ\bm{\varphi}$ and $\nabla(F_{n}\circ\bm{\varphi})$ are evaluated only once at each iteration respectively (see also Example 4.8 (a)).
For the case (ii), at $n$th iteration of Algorithm 1, Algorithm 2 for estimating stepsize $\gamma_{n}$ terminates within $\max\left\{1,\left\lceil\log_{\rho}\left(\frac{2\mu_{n}(1-c)}{\gamma_{\rm init}(\mu_{n}\varpi_{1}+\varpi_{2})}\right)\right\rceil\right\}$ updates (see Example 4.8 (b)). Since $F_{n}\circ\bm{\varphi}$ is computed once at each update of $\gamma_{n}$ in Algorithm 2, $F_{n}\circ\bm{\varphi}$ is evaluated for $M_{n}\coloneqq\max\left\{1,\left\lceil\log_{\rho}\left(\frac{2\mu_{n}(1-c)}{\gamma_{\rm init}(\mu_{n}\varpi_{1}+\varpi_{2})}\right)\right\rceil\right\}+1$ times at most in Algorithm 2 for estimating $\gamma_{n}$ at $n$th iteration of Algorithm 1. Hence, the first-order oracle complexity can be obtained by evaluating an upper bound of $\sum_{n=1}^{\mathcal{O}(\epsilon^{-3})}M_{n}$. Since $\log_{\rho}(\cdot)$ with $\rho\in(0,1)$ is monotonically decreasing and $\mu_{n}=\tau n^{-1/3}\leq\tau\ (n\in\mathbb{N})$, we have $\log_{\rho}(\frac{2\mu_{n}(1-c)}{\gamma_{\rm init}(\mu_{n}\varpi_{1}+\varpi_{2})})\leq\log_{\rho}(\frac{2\mu_{n}(1-c)}{\gamma_{\rm init}(\tau\varpi_{1}+\varpi_{2})})=\log_{\rho}(\frac{2\tau(1-c)n^{-1/3}}{\gamma_{\rm init}(\tau\varpi_{1}+\varpi_{2})})=-\frac{1}{3}\log_{\rho}(n)+\Delta=-\frac{\log(n)}{3\log(\rho)}+\Delta$ with some $\Delta\in\mathbb{R}$. From $\lceil t\rceil\leq t+1\leq\lvert t\rvert+1\ (t\in\mathbb{R})$, we get $M_{n}\leq\lvert\log_{\rho}(\frac{2\mu_{n}(1-c)}{\gamma_{\rm init}(\mu_{n}\varpi_{1}+\varpi_{2})})\rvert+1+1\leq\frac{\log(n)}{3\lvert\log(\rho)\rvert}+\lvert\Delta\rvert+2$. From $\sum_{n=1}^{k}\log(n)\leq\sum_{n=1}^{k}\log(k)=k\log(k)\ (k\in\mathbb{N})$, we get $\sum_{n=1}^{k}M_{n}\leq\frac{k\log(k)}{3\lvert\log(\rho)\rvert}+k(\lvert\Delta\rvert+2)$. By substituting $k=\mathcal{O}(\epsilon^{-3})$, we get the first-order oracle complexity $\mathcal{O}(\epsilon^{-3}\log(\epsilon^{-3}))=\mathcal{O}(\epsilon^{-3}\log(\epsilon^{-1}))$. ∎
5 Numerical experiments
We demonstrate the efficacy of Algorithm 1 with $\gamma_{n}$ given by Algorithm 2141414In our preliminary experiments, we examined that Algorithm 1 with $\gamma_{n}$ given by Algorithm 2 achieves faster convergence than Algorithm 1 with $\gamma_{n}$ in (50). Therefore, in the numerical experiments, we demonstrate the efficacy of Algorithm 1 with $\gamma_{n}$ given by Algorithm 2. by numerical experiments in scenarios of two applications: (i) the sparse spectral clustering (SSC) and (ii) the sparse principal component analysis (SPCA). In the SSC application (Section 5.1), we present a new formulation of SSC, via Problem 1.1 with a weakly convex and smooth composition, and demonstrate how this new formulation improves estimation performance of the existing SSC algorithm. In the SPCA application (Section 5.2), we demonstrate the efficient numerical performance of the proposed Algorithm 1 by comparing its empirical results, e.g., convergence speed, with those of fairly standard Riemmannian nonsmooth optimization algorithms [10, 59, 25] (Note: the proposed Algorithm 1 and the standard algorithms [10, 59, 25] are applicable to the SPCA problem). All experiments were performed by MATLAB on MacBookPro (Apple M3, 16GB).
Both applications are reduced to nonsmooth optimization problems over an embedded submanifold, say the Stiefel manifold ${\rm St}(p,N)\coloneqq\{\bm{U}\in\mathbb{R}^{N\times p}\mid\bm{U}^{\mathsf{T}}\bm{U}=\bm{I}_{p}\}\ (p\leq N)$, of $\mathbb{R}^{N\times p}$ [3, Section 3.3.2]. In order to apply Algorithm 1 to such problems, we employ the generalized inverse Cayley transform, denoted by $\Phi_{\bm{S}}^{-1}$ in [48], as a parametrization $\bm{\varphi}$ of ${\rm St}(p,N)$:
$\bm{\varphi}_{\bm{S}}^{\rm Cay}:(\mathcal{Y}\coloneqq)Q_{N,p}\to{\rm St}(p,N)\setminus E_{N,p}(\bm{S}):\bm{V}\mapsto\bm{S}(\bm{I}-\bm{V})(\bm{I}+\bm{V})^{-1}\bm{I}_{N\times p}$ (77)
with $\bm{S}\in{\rm O}(N)\coloneqq{\rm St}(N,N)$, where
$Q_{N,p}\coloneqq\left\{\left.\begin{bmatrix}\bm{A}&-\bm{B}^{\mathsf{T}}\\ \bm{B}&\bm{0}\end{bmatrix}\in\mathbb{R}^{N\times N}\;\right|\;\begin{subarray}{c}-\bm{A}^{\mathsf{T}}=\bm{A}\in\mathbb{R}^{p\times p},\\ \bm{B}\in\mathbb{R}^{(N-p)\times p}\end{subarray}\right\}\equiv\mathbb{R}^{Np-p(p+1)/2}$ (78)
is the set of skew-symmetric matrices, $E_{N,p}(\bm{S})\coloneqq\{\bm{U}\in{\rm St}(p,N)\mid\det(\bm{I}_{p}+\bm{I}_{N\times p}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{U})=0\}$ is called the singular-point set, and $\bm{I}_{N\times p}\in\mathbb{R}^{N\times p}$ is the first $p$ column vectors of the NN-by-NN identity matrix $\bm{I}$. Although the image $\bm{\varphi}_{\bm{S}}^{\rm Cay}(Q_{N,p})={\rm St}(p,N)\setminus E_{N,p}(\bm{S})$ is a proper subset of ${\rm St}(p,N)$, ${\rm St}(p,N)\setminus E_{N,p}(\bm{S})$ is open [^15] There exists an open subset $O\subset\mathbb{R}^{N\times p}$ in the Euclidean space $\mathbb{R}^{N\times p}$ such that ${\rm St}(p,N)\setminus E_{N,p}(\bm{S})=O\cap{\rm St}(p,N)$. and dense [^16] For any $\bm{U}\in{\rm St}(p,N)$, there exists $(\bm{U}_{n})_{n=1}^{\infty}\subset{\rm St}(p,N)\setminus E_{N,p}(\bm{S})$ such that $\lim_{n\to\infty}\bm{U}_{n}=\bm{U}$. in ${\rm St}(p,N)$ [48, Thm. 2.3 (b)], implying thus almost all points in ${\rm St}(p,N)$ can be parameterized in terms of the Euclidean space $Q_{N,p}$ with $\bm{\varphi}_{\bm{S}}^{\rm Cay}$.
Corollary 5.1 below implies that $\bm{\varphi}_{\bm{S}}^{\rm Cay}$ possesses well-suited properties for its application to Algorithm 1 as a parametrization $\bm{\varphi}$. More precisely, with $C\coloneqq{\rm St}(p,N)$ and $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$, (i) Corollary 5.1 (a) implies the equivalence of the first-order optimality conditions for Problem 1.1 and Problem 1.4; (ii) Corollary 5.1 (b) and Theorem 4.9 imply that Algorithm 1 is guaranteed to approximate a stationary point of Problem 1.4.
Corollary 5.1 (Properties of $\bm{\varphi}_{\bm{S}}^{\rm Cay}$ in (77) as parametrization of ${\rm St}(p,N)$).
Consider Problem 1.1 with $C\coloneqq{\rm St}(p,N)$. Let $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}:(\mathcal{Y}\coloneqq)Q_{N,p}\to{\rm St}(p,N)\setminus E_{N,p}(\bm{S})$ in (77) with arbitrarily chosen $\bm{S}\in{\rm O}(N)$. Then, the following hold:
  1. (a)
    For $\bm{V}^{\star}\in Q_{N,p}$ and $\bm{U}^{\star}\coloneqq\bm{\varphi}(\bm{V}^{\star})=\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}^{\star})\in{\rm St}(p,N)\setminus E_{N,p}(\bm{S})$,
    $\partial_{\rm L}(h+g\circ\mathfrak{S}+\iota_{C})(\bm{U}^{\star})\ni\bm{0}\Leftrightarrow\partial_{\rm L}((h+g\circ\mathfrak{S})\circ\bm{\varphi})(\bm{V}^{\star})\ni\bm{0}.$ (79)
  2. (b)
    If $\mathfrak{S}$ is twice continuously differentiable, then Assumption 4.7 (a) with $C\coloneqq{\rm St}(p,N)$ and $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$ is satisfied.
Proof.
See Appendix E. ∎
Remark 5.2 (Singular-point issue for minimizing $F\circ\bm{\varphi}_{\bm{S}}^{\rm Cay}$).
The procedure for minimization of $F\circ\bm{\varphi}_{\bm{S}}^{\rm Cay}$ with $\bm{S}\in{\rm O}(N)$ fixed may suffer from slow convergence of updating iterates [48] in a case where a minimizer $\bm{U}^{\star}\in{\rm St}(p,N)$ of $f:\mathbb{R}^{N\times p}\to\mathbb{R}$ over ${\rm St}(p,N)$ is close to the singular-point set $E_{N,p}(\bm{S})$. To suppress such a performance degradation called a singular-point issue, we recently proposed an adaptive parametrization strategy [49, 50] that updates $\bm{S}$ adaptively so that $E_{N,p}(\bm{S})$ is located far away from $\bm{U}^{\star}$, and its efficacy has been demonstrated numerically therein. This indicates that a similar adaptive parametrization strategy may improve the convergence speed of Algorithm 1, however this improvement is beyond the scope of this paper, and thus will be addressed in future work.
5.1 Sparse Spectral Clustering (SSC)
The spectral clustering (SC) [62] is one of powerful modern clustering algorithms, where the goal of clustering is to split given data $(\bm{\xi}_{i})_{i=1}^{N}\subset\mathbb{R}^{d}$ into $K\leq N$ groups without requiring labels of data. The SC has been used not only for clustering on graphs [31] in graph signal processing, but also for single-cell RNA sequence [80], remote sensing image analysis [73], and detecting clusters in networks [78].
Algorithm 3 (Sparse) Spectral clustering
1:Data $(\bm{\xi}_{i})_{i=1}^{N}\subset\mathbb{R}^{d}$.
2:Construct an affinity matrix $\bm{W}\in\mathbb{R}^{N\times N}$ whose entries $\bm{W}]_{i,j}\geq 0$ stands for the similarity between $\bm{\xi}_{i}$ and $\bm{\xi}_{j}$.
3:Compute the normalized Laplacian $\bm{L}\coloneqq\bm{I}-\bm{D}^{-1/2}\bm{W}\bm{D}^{-1/2}\in\mathbb{R}^{N\times N}$ by regarding $\bm{W}$ as the adjacency matrix of a certain graph $\mathcal{G}$, where $\bm{D}\in\mathbb{R}^{N\times N}$ is the degree (diagonal) matrix of $\bm{W}$, i.e., $\bm{D}]_{i,i}=\sum_{j=1}^{N}[\bm{W}]_{i,j}$.
4:Compute $\bm{U}^{\star}\in{\rm St}(K,N)$ by solving (80) ((82) in case of Sparse Spectral Clustering).
5:Form $\widehat{\bm{U}}^{\star}\in\mathbb{R}^{N\times K}$ by normalizing each row $\widehat{\bm{u}}^{\star}_{i}\in\mathbb{R}^{K}$ of $\bm{U}^{\star}$ to length 11.
6: Treat each row $\widehat{\bm{u}}^{\star}_{i}\in\mathbb{R}^{K}$ of $\widehat{\bm{U}}^{\star}$ as a feature vector of $\bm{\xi}_{i}$, and cluster $(\widehat{\bm{u}}_{i})_{i=1}^{N}$ into $K$ clusters by the k-means algorithm.
To split given data $(\bm{\xi}_{i})_{i=1}^{N}$ into $K$ groups, the SC is executed as in Algorithm 3 based on the graph theory (see, e.g., [77]). The SC consists of two parts: (i) constructing an affinity matrix $\bm{W}\in\mathbb{R}^{N\times N}$ and an affinity graph $\mathcal{G}$, e.g., k-nearest neighborhood graph [4], of $(\bm{\xi}_{i})_{i=1}^{N}$; (ii) decomposing $\mathcal{G}$ into $K$ connected subgraphs based on a certain spectral behavior of the normalized Laplacian $\bm{L}\coloneqq\bm{I}-\bm{D}^{-1/2}\bm{W}\bm{D}^{-1/2}\in\mathbb{R}^{N\times N}$ of $\mathcal{G}$, where $\bm{D}\in\mathbb{R}^{N\times N}$ is the degree matrix of $\mathcal{G}$. More precisely for (ii), the SC exploits the fact that if $\bm{L}$ has an eigenvalue 0 with the multiplicity $K$, then $\mathcal{G}$ can be decomposed into $K$ connected subgraphs by clustering the eigenvectors associated with the eigenvalue 0 (see, e.g., [77]). By leveraging this fact, the SC computes the $K$ smallest eigenvectors of $L$, in the third step of Algorithm 3, which can be characterized as a minimizer of the following optimization problem:
$\mathrm{find}\ \bm{U}^{\star}\in\mathop{\mathrm{argmin}}\limits_{\bm{U}\in{\rm St}(K,N)}{\rm Tr}(\bm{U}^{\mathsf{T}}\bm{L}\bm{U}).$ (80)
In the fourth step of Algorithm 3, we normalize each row vector, say $\widehat{\bm{u}}^{\star}_{i}\in\mathbb{R}^{K}$, of $\bm{U}^{\star}$ as $\left\lVert\widehat{\bm{u}}^{\star}_{i}\right\rVert_{2}=1$. Finally, we cluster $(\widehat{\bm{u}}^{\star}_{i})_{i=1}^{N}$ into $K$ groups by using the k-means algorithm.
In order to improve the numerical performance of the SC, the sparse SC (SSC) [61, 80] exploits a prior knowledge that the affinity matrix $\bm{W}$ and Laplacian $\bm{L}$ can be block diagonal, in the ideal case where there is no edge between nodes in different groups, by using a certain permutation $p:\{1,2,\ldots,N\}\to\{1,2,\ldots,N\}$ of indices of $(\bm{\xi}_{i})_{i=1}^{N}$ as
$(\underbrace{\bm{\xi}_{p(1)},\bm{\xi}_{p(2)},\ldots,\bm{\xi}_{p(\mathcal{I}_{1})}}_{1\mathrm{st\ group}},\underbrace{\bm{\xi}_{p(\mathcal{I}_{1}+1)},\ldots,\bm{\xi}_{p(\mathcal{I}_{1}+\mathcal{I}_{2})}}_{2\mathrm{nd\ group}},\ldots,\underbrace{\bm{\xi}_{p(\sum_{k=1}^{K-1}\mathcal{I}_{k}+1)},\ldots,\bm{\xi}_{p(N)}}_{K\mathrm{th\ group}}),$ (81)
where $\mathcal{I}_{k}\in\mathbb{N}$ denotes the size of $k$th group. The block diagonality of $\bm{L}$ indicates the block diagonality of $\bm{U}^{\star}\bm{U}^{\star\mathsf{T}}\in\mathbb{R}^{N\times N}$ with $\bm{U}^{\star}\in{\rm St}(K,N)$ in (80), thereby inducing the sparsity of $\bm{U}^{\star}\bm{U}^{\star\mathsf{T}}\in\mathbb{R}^{N\times N}$. Although the above permutation is not available in general, $\bm{U}^{\star}\bm{U}^{\star\mathsf{T}}$ is expected to be sparse even if any permutation is employed. Therefore, in the SSC, the sparsity of $\bm{U}^{\star}\bm{U}^{\star\mathsf{T}}$ is promoted via solving the following problem in place of (80):
$\mathrm{find}\ \bm{U}^{\star}\in\mathop{\mathrm{argmin}}\limits_{\bm{U}\in{\rm St}(K,N)}{\rm Tr}(\bm{U}^{\mathsf{T}}\bm{L}\bm{U})+\lambda\psi\circ\mathfrak{S}(\bm{U}),$ (82)
where $\lambda>0$, $\mathfrak{S}:\mathbb{R}^{N\times K}\to\mathbb{R}^{N\times N}:\bm{U}\mapsto\bm{U}\bm{U}^{\mathsf{T}}$, and $\psi:\mathbb{R}^{N\times N}\to\mathbb{R}$ is a sparsity promoting function, e.g., $\ell_{1}$-norm, MCP, and SCAD (see, e.g., [8, 27]).
By Corollary 5.1, the problem (82) can be reformulated as Problem 1.4 with $C\coloneqq{\rm St}(K,N)$, $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$ in (77), $h(\bm{U})\coloneqq{\rm Tr}(\bm{U}^{\mathsf{T}}\bm{L}\bm{U})$ and $g\coloneqq\lambda\psi$:
$\mathrm{find}\ \bm{V}^{\star}\in\mathop{\mathrm{argmin}}\limits_{\bm{V}\in Q_{N,K}}(\underbrace{h+g\circ\mathfrak{S}}_{\eqqcolon F})\circ\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}).$ (83)
By employing a Lipschitz continuous and weakly convex function as $\psi$, we can find a stationary point of (82) by Algorithm 1 (see Theorem 4.9 with Corollary 5.1). We note that (i) standard Riemannian nonsmooth optimization algorithms, e.g., [25, 10, 66], cannot be applied to the problem (82) because $\mathfrak{S}$ is nonlinear; (ii) [80] reported a prox-linear method for the problem (82) in a case where $\psi$ is the convex $\ell_{1}$-norm.
In this numerical experiment, we evaluated the numerical performance of SSC via the formulation in (83) by applying Algorithm 1 with $\gamma_{n}$ given by Algorithm 2 under the setting (for a choice of $(\mu_{n})_{n=1}^{\infty}$, we follow the setting in the oracle complexity result; see Corollary 4.12)
$(\mu_{n})_{n=1}^{\infty}=((2\eta)^{-1}n^{-1/3})_{n=1}^{\infty},c=2^{-13},\rho=0.5,\gamma_{\rm init}=\min\left\{1,\left\lVert\nabla(F_{1}\circ\bm{\varphi}_{\bm{S}}^{\rm Cay})(\bm{V}_{1})\right\rVert_{F}^{-1}\right\},$ (84)
where $\left\lVert\cdot\right\rVert_{F}$ is the Frobenius norm, $F_{1}$ is defined as in Algorithm 1, and $(\bm{S},\bm{V}_{1})\in{\rm O}(N)\times Q_{N,K}$ was chosen [^17] By following [48, Thm. 2.7], a pair $(\bm{S},\bm{V}_{1})\in{\rm O}(N)\times Q_{N,K}$ satisfying $\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}_{1})=\bm{U}_{1}$ can be chosen as $\bm{S}\coloneqq\mathrm{diag}(\bm{Q}_{1}\bm{Q}_{2}^{\mathsf{T}},\bm{I}_{N-K})$ and $\bm{V}_{1}\coloneqq\Phi_{\bm{S}}(\bm{U}_{1})$, where $\bm{Q}_{1},\bm{Q}_{2}\in{\rm O}(K)$ are obtained by a singular-value decomposition of the upper block matrix $\bm{U}_{0{\rm up}}=\bm{Q}_{1}\bm{\Sigma}\bm{Q}_{2}\in\mathbb{R}^{K\times K}$ of $\bm{U}_{1}$ with a nonnegative-valued diagonal matrix $\bm{\Sigma}\in\mathbb{R}^{K\times K}$, and $\Phi_{\bm{S}}:{\rm St}(K,N)\setminus E_{N,K}(\bm{S})\to Q_{N,K}$ is the generalized Cayley transform [48, (10)-(12)] (Note: $\Phi_{\bm{S}}\circ\bm{\varphi}_{\bm{S}}^{\rm Cay}=\mathrm{Id}$ and $\bm{\varphi}_{\bm{S}}^{\rm Cay}\circ\Phi_{\bm{S}}=\mathrm{Id}$ hold respectively on their domains). to satisfy $\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}_{1})=\bm{U}_{1}$ for an initial guess $\bm{U}_{1}\in{\rm St}(K,N)$. In order to examine how much weakly convex regularizers can enhance the performance, we compared (i) the proposed SSC (SSC+MCP) by employing MCP [89]
$(\bm{X}\in\mathbb{R}^{N\times N})\quad$ $\psi^{\rm MCP}_{\theta}(\bm{X})\coloneqq\sum_{1\leq i\leq N}\sum_{1\leq j\leq N}r^{\rm MCP}_{\theta}([\bm{X}]_{i,j})$ (85)
$\mathrm{with}\ r^{\rm MCP}_{\theta}([\bm{X}]_{i,j})\coloneqq\begin{cases}\lvert[\bm{X}]_{i,j}\rvert-\frac{[\bm{X}]_{i,j}^{2}}{2\theta}&(\mathrm{if}\ \lvert[\bm{X}]_{i,j}\rvert\leq\theta)\\ \frac{\theta}{2}&(\mathrm{otherwise})\end{cases}$ (86)
[…]
$\mathrm{find}\ \bm{P}^{\star}\in\mathop{\mathrm{argmin}}\limits_{\bm{P}\in\mathbb{R}^{N\times N},\bm{0}\preceq\bm{P}\preceq\bm{I},{\rm Tr}(\bm{P})=K}{\rm Tr}(\bm{P}^{\mathsf{T}}\bm{L})+\lambda\left\lVert\bm{P}\right\rVert_{1},$ (87)
where $\{\bm{U}\bm{U}^{\mathsf{T}}\in\mathbb{R}^{N\times N}\mid\bm{U}\in{\rm St}(K,N)\}$ is relaxed to the convex set $\{\bm{P}\in\mathbb{R}^{N\times N}\mid\bm{0}\preceq\bm{P}\preceq\bm{I},{\rm Tr}(\bm{P})=K\}$. The problem (87) can be solved by classical convex optimization algorithms, e.g., ADMM (see, e.g., [28]). In the SSC[relax] [61], $\bm{U}^{\star}$ in the third step of Algorithm 3 is obtained as first $K$ eigenvectors corresponding to the largest $K$ eigenvalues of $\bm{P}^{\star}$ in (87).
To evaluate clustering performance, we computed two standard criteria: the Normalized Mutual Information (NMI) [72] and the Adjusted Rand Index (ARI) [45]. These scores closer to one indicate better clustering performance.
[…]
Table 2: Averaged NMI and ARI of each algorithm.
iris shuttle segmentation breast cancer glass wine seeds
NMI SC 0.778 0.435 0.501 0.417 0.321 0.433 0.662
SSC[relax] 0.785 0.485 0.503 0.433 0.322 0.433 0.671
[…]
SSC+MCP 0.794 0.496 0.507 0.514 0.331 0.432 0.698
ARI SC 0.745 0.231 0.341 0.419 0.174 0.363 0.659
SSC[relax] 0.786 0.315 0.343 0.462 0.174 0.363 0.675
[…]
SSC+MCP 0.794 0.397 0.352 0.595 0.181 0.388 0.709
5.2 Sparse Principal Component Analysis (SPCA)
We consider the SPCA (see, e.g., [25, 10, 66]) formulated as
$\mathrm{find}\ \bm{U}^{\star}\in\mathop{\mathrm{argmin}}\limits_{\bm{U}\in{\rm St}(p,N)}\underbrace{-{\rm Tr}(\bm{U}^{\mathsf{T}}\bm{\Xi}^{\mathsf{T}}\bm{\Xi}\bm{U})}_{\eqqcolon h(\bm{U})}+\underbrace{\lambda\left\lVert\bm{U}\right\rVert_{1}}_{\eqqcolon g(\bm{U})}$ (88)
where $\bm{\Xi}\in\mathbb{R}^{\mathcal{I}\times N}$ is a data matrix, $\mathcal{I}\in\mathbb{N}$ is the number of data, $N$ is the dimension of each data, and $\lambda>0$ is a predetermined weight. The SPCA in (88) is a special instance of Problem 1.1 with $C\coloneqq{\rm St}(p,N)$ and $\mathfrak{S}=\mathrm{Id}$. By Corollary 5.1, the SPCA in (88) can be reformulated as Problem 1.4 with $C\coloneqq{\rm St}(p,N)$ and $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$ in (77):
$\mathrm{find}\ \bm{V}^{\star}\in\mathop{\mathrm{argmin}}\limits_{\bm{V}\in Q_{N,p}}(h+g)\circ\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V})$ (89)
In this experiment, we applied Algorithm 1 (VSmooth) to the problem in (89) under the setting (84) with $\gamma_{n}$ given by Algorithm 2.
We compared the numerical performance of VSmooth with that of three Riemannian nonsmooth optimization algorithms applied to the problem in (88): (i) the Riemannian smoothing gradient algorithm [10, Section 4.1] (RSmooth), (ii) the Riemannian subgradient method [59, (2.6)] (RSub), and (iii) the Riemannian proximal gradient method [25, Algorithm 2] (ManPGAda [^20] We used codes in. https://github.com/chenshixiang/ManPG. ). For each Riemannian optimization algorithm, the update from $\bm{U}_{n}\in{\rm St}(p,N)$ to $\bm{U}_{n+1}\in{\rm St}(p,N)$ is expressed as
$\bm{U}_{n+1}\coloneqq R_{\bm{U}_{n}}(\gamma_{n}\bm{\mathcal{D}}_{n})$ (90)
with a stepsize $\gamma_{n}>0$ and a search direction $\bm{\mathcal{D}}_{n}\in T_{{\rm St}(p,N)}(\bm{U}_{n})\coloneqq\{\bm{\mathcal{D}}\in\mathbb{R}^{N\times p}\mid\bm{U}_{n}^{\mathsf{T}}\bm{\mathcal{D}}+\bm{\mathcal{D}}^{\mathsf{T}}\bm{U}_{n}=\bm{0}\}$ in the tangent space $T_{{\rm St}(p,N)}(\bm{U}_{n})$ (see Example 2.3), where $R_{\bm{U}_{n}}:T_{{\rm St}(p,N)}(\bm{U}_{n})\to{\rm St}(p,N)$ is a retraction satisfying $R_{\bm{U}_{n}}(\bm{0})=\bm{U}_{n}$ and $\mathrm{D}R_{\bm{U}_{n}}(\bm{0})[\bm{\mathcal{D}}]=\bm{\mathcal{D}}\ (\bm{\mathcal{D}}\in T_{{\rm St}(p,N)}(\bm{U}_{n}))$ (see, e.g., [3, 70, 16]). For these algorithms, we employed the polar decomposition-based retraction [3] $R_{\bm{U}}^{\rm polar}(\bm{\mathcal{D}})\coloneqq(\bm{U}+\bm{\mathcal{D}})(\bm{I}_{p}+\bm{\mathcal{D}}^{\mathsf{T}}\bm{\mathcal{D}})^{-\frac{1}{2}}\ (\bm{U}\in{\rm St}(p,N),\bm{\mathcal{D}}\in T_{{\rm St}(p,N)}(\bm{U}))$ as the retraction $R_{\bm{U}}$ because $R_{\bm{U}}^{\rm polar}$ is the retraction that was used in [25, 59, 10]. Every search direction $\bm{\mathcal{D}}_{n}$ in (90) is given respectively by:
$\bm{\mathcal{D}}_{n}^{\rm RSmooth}$ $\coloneqq-P_{T_{{\rm St}(p,N)}(\bm{U}_{n})}(\nabla(h+{}^{\mu_{n}}g)(\bm{U}_{n}))\ \mathrm{with}\ \mu_{n}\searrow 0;$ (91)
$\bm{\mathcal{D}}_{n}^{\rm RSub}$ $\coloneqq-P_{T_{{\rm St}(p,N)}(\bm{U}_{n})}(\nabla h(\bm{U}_{n})+\bm{\mathcal{G}})\ \mathrm{with}\ \bm{\mathcal{G}}\in\partial_{\rm L}g(\bm{U}_{n});$ (92)
$\bm{\mathcal{D}}_{n}^{\rm ManPGAda}$ $\coloneqq\mathop{\mathrm{argmin}}\limits_{\bm{\mathcal{D}}\in T_{{\rm St}(p,N)}(\bm{U}_{n})}{\left\langle\nabla h(\bm{U}_{n}),\bm{\mathcal{D}}\right\rangle}+\frac{1}{2t_{n}}\left\lVert\bm{\mathcal{D}}\right\rVert_{\bm{\varphi}}^{2}+g(\bm{U}_{n}+\bm{\mathcal{D}})$ (93)
$\hskip 30.00005pt\mathrm{with}\ t_{n}>0,$ (94)
where $P_{T_{{\rm St}(p,N)}(\bm{U}_{n})}:\mathbb{R}^{N\times p}\to T_{{\rm St}(p,N)}(\bm{U}_{n})$ is the orthogonal projection mapping onto $T_{{\rm St}(p,N)}(\bm{U}_{n})$ (see, e.g., [3]), and $t_{n}>0$ for ManPGAda [25] is adaptively designed to accelerate convergence speed. Clearly, ManPGAda requires some iterative solver for the subproblem (94) to find a search direction while the others do not require, i.e., the others are single-loop algorithms. For RSmooth, we employed $(\mu_{n})_{n=1}^{\infty}=((2\eta)^{-1}n^{-1/3})_{n=1}^{\infty}$ because RSmooth with this $(\mu_{n})_{n=1}^{\infty}$ achieves the best convergence rate [10, Remark 4.1]. In this experiment, to find a stepsize $\gamma_{n}$, we used backtracking algorithms for ManPGAda and RSmooth suggested in [25] and [10] respectively, and we employed $\gamma_{n}\coloneqq 0.99^{n}$ for RSub.
For the problem (88) with each $N\in\{200,500,1000\}$, $p\in\{1,N/100,N/10\}$ and $\lambda=0.1$, we tested performance of all algorithms with $10$ random $\bm{\Xi}\in\mathbb{R}^{\mathcal{I}\times N}$ with $\mathcal{I}=5000$ generated by the following procedure: (i) generate $\bm{\Xi}\in\mathbb{R}^{\mathcal{I}\times N}$ from the standard normal distribution; (ii) shift the columns of $\bm{\Xi}$ to have zero means; (iii) normalize $\bm{\Xi}$ such that $\left\lVert\bm{\Xi}\right\rVert_{F}=1$. For each trial, all algorithms started with the same initial guess $\bm{U}_{1}\in{\rm St}(p,N)$ generated by a MATLAB code “orth(randn(N,p))”. Each algorithm was terminated at the $n$th iteration when the elapsed CPU time at the end of $n$th iteration exceeded $t(>0)$ seconds (see Table 3 for the value of $t$).
Table 3: Time limits $t$ (s) used for termination
p=1 p=100/N p=10/N
N=200 0.5 0.5 1.5
N=500 1.5 1.5 3.0
N=1000 3.0 3.0 9.0
Figure 5.2 demonstrates the convergence histories for the problem (88), where the plots show CPU time on the horizontal axis versus the averaged cost function value on the vertical axis. The detailed results are demonstrated in Table 4, where for each final estimate $\bm{U}^{\diamond}\in{\rm St}(p,N)$, ’fval’ means the averaged value $f(\bm{U}^{\diamond})$, ’feasi’ means the averaged feasibility error $\left\lVert\bm{I}_{p}-\bm{U}^{\diamond\mathsf{T}}\bm{U}^{\diamond}\right\rVert_{F}$ of $\bm{U}^{\diamond}$ from the constraint set ${\rm St}(p,N)$, ’itr’ means the averaged number of iterations, ’time’ means the CPU time (s), and ’sparsity’ means the averaged ratio of entries in $\bm{U}^{\diamond}$ such that $\lvert[\bm{U}^{\diamond}]_{i,j}\rvert<10^{-4}$.
From Figure 5.2, we see that the proposed VSmooth converges faster than RSmooth and RSub for all problem sizes. Thus, VSmooth achieves the best performance among algorithms that do not require any iterative solver for subproblems. For $p\in\{1,N/100\}$, VSmooth seems to be competitive or inferior to ManPGAda. However, for $N\in\{500,1000\}$ and large $p=N/10$, the performance of ManPGAda severely deteriorates while no such deterioration is observed for VSmooth. Moreover, from Table 4, we also observe that ManPGAda updates for only a few iterations; and CPU time of ManPGAda significantly exceeds the time limitation shown by Table 3. These observations imply that ManPGAda took a long time to find an approximated solution of a subproblem in (94) at every iteration, and thus single-loop algorithms, namely, VSmooth, RSmooth and RSub, are more computationally reliable than ManPGAda for large problem sizes. From this experimental result, VSmooth has a comparable numerical performance compared to the fairly standard Riemannian nonsmooth optimization algorithms [25, 59, 10] even for Problem 1.1 with a simple case, i.e., $g$ is convex and $\mathfrak{S}=\mathrm{Id}$, where the proposed VSmooth can be applied to Problem 1.1, via Problem 1.4, with more general setting of the nonlinearity of $\mathfrak{S}$ as seen in Section 5.1.
Figure 1: Convergence histories of all algorithms for the problem (88) regarding the value $f(\bm{U})$ at CPU time for each problem size. Markers are put at every 100 iterations.
Table 4: Averaged performance of all algorithms for the problem (88).
Problem size Algorithm fval feasi itr time sparsity
$(N,p)=(200,1)$ RSmooth 9.50229e-02 7.216e-16 3673.8 0.50 9.950e-01
RSub 9.50082e-02 0.000e+$0$ 17357 0.50 9.950e-01
ManPGAda 9.50082e-02 2.220e-17 3927.5 0.50 9.950e-01
VSmooth 9.50728e-02 2.220e-16 5493 0.50 9.950e-01
$(N,p)=(200,2)$ RSmooth 1.90027e-01 9.578e-16 2292.4 0.50 9.950e-01
RSub 1.89986e-01 9.687e-106 11810.9 0.50 9.950e-01
ManPGAda 1.89973e-01 2.116e-16 1956.2 0.50 9.950e-01
VSmooth 1.90124e-01 7.477e-16 2994.9 0.50 9.950e-01
$(N,p)=(200,20)$ RSmooth 1.90032e+$0$ 5.081e-15 1646.5 1.50 9.950e-01
RSub 1.89997e+$0$ 3.513e-15 8918.2 1.50 9.950e-01
ManPGAda 1.89987e+$0$ 6.250e-15 1377.4 1.50 9.950e-01
VSmooth 1.90149e+$0$ 3.050e-15 2552.4 1.50 9.950e-01
$(N,p)=(500,1)$ RSmooth 9.80240e-02 7.661e-16 1943 1.50 9.980e-01
RSub 9.80058e-02 0.000e+$0$ 7027.6 1.50 9.980e-01
ManPGAda 9.80058e-02 2.220e-17 1805.2 1.50 9.980e-01
VSmooth 9.80899e-02 1.554e-16 2513.7 1.50 9.980e-01
$(N,p)=(500,5)$ RSmooth 4.90112e-01 3.119e-15 1133.5 1.50 9.980e-01
RSub 4.89997e-01 9.807e-16 4875.8 1.50 9.980e-01
ManPGAda 4.90014e-01 2.370e-15 784.8 1.50 9.980e-01
VSmooth 4.90525e-01 2.025e-15 1299.5 1.50 9.980e-01
$(N,p)=(500,50)$ RSmooth 4.98856e+$0$ 1.732e-14 774.8 4.51 9.977e-01
RSub 4.90002e+$0$ 9.091e-15 2145.3 4.50 9.980e-01
ManPGAda 1.02775e+01 1.992e-14 5 5.57 8.736e-01
VSmooth 4.90486e+$0$ 6.629e-15 1150.1 4.50 9.980e-01
$(N,p)=(1000,1)$ RSmooth 9.90208e-02 8.882e-16 1076.8 3.00 9.990e-01
RSub 9.89987e-02 0.000e+$0$ 3238 3.00 9.990e-01
ManPGAda 9.89987e-02 4.441e-17 967.9 3.00 9.990e-01
VSmooth 9.91019e-02 3.775e-16 1338.7 3.00 9.990e-01
$(N,p)=(1000,10)$ RSmooth 1.02418e+$0$ 6.645e-15 570.5 3.01 9.988e-01
RSub 9.90010e-01 2.970e-14 1700.3 3.00 9.990e-01
ManPGAda 9.90017e-01 4.901e-15 328.8 3.00 9.990e-01
VSmooth 9.91335e-01 3.006e-15 612.5 3.00 9.990e-01
$(N,p)=(1000,100)$ RSmooth 1.12218e+01 3.053e-14 376.2 9.02 9.973e-01
RSub 1.02637e+01 3.397e-14 723.6 9.01 9.051e-01
ManPGAda 4.77660e+01 3.306e-14 3 18.35 3.618e-01
VSmooth 1.01238e+01 9.032e-15 502.4 9.01 9.989e-01
6 Conclusions
We addressed a manifold constrained nonsmooth optimization problem, where the cost function consists of a smooth function and the composite function of a weakly convex function and a smooth (nonlinear) mapping. For the target problem, we proposed a variable smoothing algorithm with a parametrization of a nonconvex constraint set. We also presented an asymptotic convergence analysis and a first-order oracle complexity $\mathcal{O}(\epsilon^{-3})$ of the proposed algorithm in terms of stationary points. The numerical experiments demonstrate that (i) the proposed algorithm improves the numerical performance of sparsity-aware application by using a weakly convex regularizer; (ii) the proposed algorithm has a comparable numerical performance compared with the existing Riemannian nonsmooth optimization algorithms based on a retraction.
[…]
Acknowledgements This work was supported partially by JSPS Grants-in-Aid (19H04134, 22KJ1270, 24K23885) and by JST SICORP (JPMJSC20C6). We thank the anonymous reviewers for their thoughtful comments, which greatly improved the clarity and quality of the paper.
Appendix A Related Work
A.1 Smoothing-type algorithms
We review smoothing-type algorithms for finding a stationary point of a nonsmooth function $J:\mathcal{X}\to\mathbb{R}$ because the proposed algorithm (Algorithm 1) can be viewed as a variant of smoothing-type algorithms. Smoothing methods have been extensively developed [87, 26, 44, 79, 53, 88, 63] with a smoothed surrogate function, say $J^{\langle\mu\rangle}:\mathcal{X}\to\mathbb{R}\ (\mu\in\mathbb{R}_{++})$, having the (sub-)gradient consistency property (see Remark 4.5) with $\lim_{\mathbb{R}_{++}\ni\mu\to 0}J^{\langle\mu\rangle}(\bm{x})=J(\bm{x})\ (\bm{x}\in\mathcal{X})$. The gradient consistency property has been utilized as a key tool to express the first-order optimality condition in terms of a gradient sequence of $\nabla J^{\langle\mu\rangle}$ with $\mu\searrow 0$. The basic update of a smoothing gradient method [26] is given by $\bm{x}_{n+1}\coloneqq\bm{x}_{n}-\gamma_{n}\nabla J^{\langle\mu_{n}\rangle}(\bm{x}_{n})\ (n\in\mathbb{N})$ with a properly chosen stepsize $\gamma_{n}\in\mathbb{R}_{++}$, where smoothing parameters $\mu_{n}$ are conditionally updated with $\sigma\in(0,1)$ and $\delta\in\mathbb{R}_{++}$ by setting $\mu_{n+1}\coloneqq\sigma\mu_{n}$ if $\left\lVert\nabla J^{\langle\mu_{n}\rangle}(\bm{x}_{n})\right\rVert<\delta\mu_{n}$; $\mu_{n+1}\coloneqq\mu_{n}$ otherwise. Thanks to the (sub-)gradient consistency property of $J^{\langle\mu\rangle}$, a cluster point of $(\bm{x}_{n})_{n=1}^{\infty}$ is guaranteed to be a (Clarke) stationary point of $J$ [26]. Based on the conditional update of smoothing parameters, many numerical algorithms, e.g., accelerated gradient methods [79, 63], projected gradient methods [87, 44] and manifold constrained smoothing methods [53, 88], for updating $(\bm{x}_{n})_{n=1}^{\infty}$ have been established with asymptotic convergence analyses in terms of a stationary point. However, these smoothing methods can be viewed as double-loop algorithms due to the conditional update of smoothing parameters.
Another class of smoothing-type methods has been proposed with iteratively decreasing smoothing parameters $(\mu_{n})_{n=1}^{\infty}$ [15, 74, 14, 13, 46, 81, 10, 66, 64], where these methods are called variable/adaptive/dynamic smoothing methods. Thanks to the use of iteratively decreasing smoothing parameters, these variable smoothing methods are single-loop algorithms. For convex optimization problems, the oracle complexities for finding the optimal value of the cost function have been shown for [15, 74, 14, 46, 81, 64], where an exceptional paper [81] guarantees that a generated sequence converges to a minimizer of the cost function. Recently, for nonconvex nonsmooth optimization problems, a variable smoothing algorithm [15] has been extended by [13] to a special version of Problem 1.1 where $\mathfrak{S}$ is a linear operator and $C\coloneqq\mathcal{X}$. The variable smoothing algorithm in [13] uses a smoothed surrogate function $h+{}^{\mu}g\circ\mathfrak{S}$ with the Moreau envelope $^{\mu}g$ of $g$ (see (21)), and its first-order oracle complexity $\mathcal{O}(\epsilon^{-3})$ for finding a certain $\epsilon$-approximate stationary point in the sense of (73) has been shown under the linearity of $\mathfrak{S}$. This variable smoothing algorithm [13] with an oracle complexity guarantee has also been extended by [10, 66] to Problem 1.1 where $C$ is a compact embedded submanifold of $\mathcal{X}$, but $\mathfrak{S}$ is still assumed to be linear. However, for variable smoothing algorithms [13, 66], except for [10], any asymptotic convergence analysis to a stationary point has not been reported yet, where [10] presents an asymptotic convergence analysis by exploiting the compactness of $C$ and the linearity of $\mathfrak{S}$ (see also Table 1 in Section 1).
A.2 Penalty-based strategy
In a case where $C$ in Problem 1.1 is defined by equality constraints, i.e., $C=H^{-1}(\bm{0})$ with some continuously differentiable mapping $H:\mathcal{X}\to\mathbb{R}^{m}$, penalty-based strategies have been proposed, e.g., [65, Ch. 12][82], as an alternative strategy to the parametrization strategy in Section 3 for translating Problem 1.1 into an unconstrained problem. Such penalty-based strategies aim to find a point satisfying the Karush-Kuhn-Tucker (KKT) condition by assuming the availability of $H$ with suitable regularities, e.g., Linearly Independent Constraint Qualification (LICQ), i.e., $\mathrm{rank}(\mathrm{D}H(\bm{x}))=m$ for all $\bm{x}\in C$, where $H$ with LICQ is called a (global) defining map [16, Cor. 8.76]. Indeed, LICQ of $H$ makes the KKT condition a necessary condition for the local optimality of Problem 1.1 with a smooth $F$ [65, Thm. 12.1], while the KKT condition without LICQ of $H$ is not guaranteed to be a necessary condition for the local optimality. However, the existence of a global defining map $H$ with LICQ is not guaranteed for general embedded submanifolds [16, p.50], e.g., non-orientable manifolds [54, Prop. 15.23, pp.106-107] (see [76] for applications of optimization over non-orientable manifolds). Moreover, penalty-based strategies [65, 82] may generate an infeasible sequence $(\bm{x}_{n})_{n=1}^{\infty}\subset\mathcal{X}$, i.e., $\bm{x}_{n}$ may not belong to $C$, where the feasibility is guaranteed at cluster points of $(\bm{x}_{n})_{n=1}^{\infty}$ in general.
In contrast, for an embedded submanifold $C$ in Problem 1.1, a parametrization $\bm{\varphi}$ of $C$ exists (see Example 1.3 (c)), and a stationary point $\bm{y}^{\star}\in\mathcal{Y}$ of Problem 1.4 satisfies a necessary condition for the local optimality of Problem 1.1 at $\bm{x}^{\star}\coloneqq\bm{\varphi}(\bm{y}^{\star})\in C$ (see Remark 3.6). Under an additional condition, e.g., submersion condition on $\bm{\varphi}$, $\bm{x}^{\star}$ is a stationary point of Problem 1.1 (see Theorem 3.2 and Corollary 3.4). Furthermore, the parametrization strategy described in Section 3 generates a sequence $(\bm{y}_{n})_{n=1}^{\infty}\subset\mathcal{Y}$ for finding a stationary point $\bm{y}^{\star}$; hence the feasibility of $(\bm{\varphi}(\bm{y}_{n}))_{n=1}^{\infty}\subset C$ is guaranteed via $\bm{\varphi}$ (see Algorithm 1 in Section 4.2; see also, e.g., [84, 36, 43, 41, 58, 47, 48, 56], for parametrization strategy with smooth cost function).
Finally, we remark that the submersion condition of a parametrization $\bm{\varphi}$ and LICQ of $H$ are closely related if both $\bm{\varphi}$ with the submersion condition and a global defining map $H$ with LICQ exist for an embedded submanifold $C$. Indeed, in this case, we have $T_{C}(\bm{x})=\mathrm{Ker}(\mathrm{D}H(\bm{x}))=\mathrm{Ran}(\mathrm{D}\bm{\varphi}(\bm{y}))\ (\bm{y}\in\mathcal{Y},\bm{x}\coloneqq\bm{\varphi}(\bm{y})\in C)$. However, the availabilities of $\bm{\varphi}$ with the submersion condition and a global defining map $H$ with LICQ are not comparable in general because we have an example [^21] The Klein bottle has no global defining map satisfying LICQ due to its non-orientability (see, e.g., [76, Sect. 3.3.1] and [54, Prop. 15.23, pp.106-107]). However, since $\mathbb{R}^{2}$ is the universal covering manifold [54, p.90] of the Klein bottle (see, e.g., [6]), there exists a smooth covering mapping [54, p.90] from $\mathbb{R}^{2}$ onto the Klein bottle. Such a smooth covering mapping is a parametrization $\bm{\varphi}$, defined over $\mathbb{R}^{2}$, of the Klein bottle such that the submersion condition is satisfied at every point in $\mathbb{R}^{2}$ by [54, Prop. 4.33 (a)]. of $C$ such that any global defining map $H$ of $C$ with LICQ does not exist, whereas a parametrization $\bm{\varphi}:\mathcal{Y}\to\mathcal{X}$ of $C$ with the submersion condition exists.
Appendix B Proof of Lemma 2.4
Although Lemma 2.4 can be checked by the subdifferential calculus rules [69, Thm. 10.6, Cor. 10.9], we present a short proof of Lemma 2.4 with some necessary facts in order to increase the readability for researchers in the area of applications. The subdifferential calculus rules heavily rely on the horizontal subdifferential, which gives constraint qualifications, and the subdifferential regularity of a function (see Definition 2.2 (b) and Example B.2).
Definition B.1 (Horizontal subdifferential [69, Def. 8.3 and Thm. 8.9]).
Let $J:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ be a proper lower semicontinuous function. Then, $\partial^{\infty}J(\widebar{\bm{x}})\coloneqq\{\bm{v}\in\mathcal{X}\mid(\bm{v},0)\in N_{\mathrm{epi}(J)}(\widebar{\bm{x}},J(\widebar{\bm{x}}))\}$ is called the horizontal subdifferential of $J$ at $\widebar{\bm{x}}\in\mathrm{dom}(J)$.
Example B.2 (Subdifferentially regular functions).
Subdifferentially regular functions include, e.g., continuously differentiable functions [69, Exm. 7.28], proper lower semicontinuous (weakly) convex functions [69, Exm. 7.27] [30, Prop. 4.4.15], and indicator functions of Clarke regular subsets [69, Exm. 7.28].
Fact B.3 presents useful properties, regarding $\partial^{\infty}J$, that will be used for our analyses.
Fact B.3 (Horizontal subdifferential being zero [69, Thm. 9.13]).
Let $J:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ be a proper lower semicontinuous function. Then, for $\widebar{\bm{x}}\in\mathrm{dom}(J)$, the following conditions are equivalent:
  1. (a)
    $\partial^{\infty}J(\widebar{\bm{x}})=\{\bm{0}\}$.
  2. (b)
    $J$ is locally Lipschitz continuous at $\widebar{\bm{x}}$.
  3. (c)
    $\partial_{\rm L}J$ is locally bounded at $\widebar{\bm{x}}$, i.e., there exists an open neighborhood $\mathcal{N}_{\widebar{\bm{x}}}\subset\mathcal{X}$ of $\widebar{\bm{x}}$ such that $\partial_{\rm L}J(\mathcal{N}_{\widebar{\bm{x}}})=\bigcup_{\bm{x}\in\mathcal{N}_{\widebar{\bm{x}}}}\partial_{\rm L}J(\bm{x})$ is bounded.
Moreover, if these conditions hold at $\widebar{\bm{x}}\in\mathrm{dom}(J)$, then $\partial_{\rm L}J(\widebar{\bm{x}})\neq\emptyset$.
Fact B.4 (Calculus rules for subdifferential).
Let $\mathcal{X}$ and $\mathcal{Y}$ be Euclidean spaces. Let $J_{1},J_{2}:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ be proper lower semicontinuous and subdifferentially regular functions, and $\mathcal{F}:\mathcal{Y}\to\mathcal{X}$ be a continuously differentiable mapping. Then, the following hold:
  1. (a)
    (Chain rule [69, Thm. 10.6]) If the constraint qualification $\partial^{\infty}J_{1}(\mathcal{F}(\widebar{\bm{y}}))\cap\mathrm{Ker}((\mathrm{D}\mathcal{F}(\widebar{\bm{y}}))^{*})=\{\bm{0}\}$ holds at $\widebar{\bm{y}}\in\mathrm{dom}(J_{1}\circ\mathcal{F})$, then $J_{1}\circ\mathcal{F}$ is subdifferentially regular at $\widebar{\bm{y}}$ and
    $\partial_{\rm L}(J_{1}\circ\mathcal{F})(\widebar{\bm{y}})$ $=(\mathrm{D}\mathcal{F}(\widebar{\bm{y}}))^{*}[\partial_{\rm L}J_{1}(\mathcal{F}(\widebar{\bm{y}}))];$ (95)
    $\partial^{\infty}(J_{1}\circ\mathcal{F})(\widebar{\bm{y}})$ $=(\mathrm{D}\mathcal{F}(\widebar{\bm{y}}))^{*}[\partial^{\infty}J_{1}(\mathcal{F}(\widebar{\bm{y}}))].$ (96)
  2. (b)
    (Sum rule [69, Cor. 10.9, Exe. 8.8]) If the constraint qualification $\partial^{\infty}J_{1}(\widebar{\bm{x}})\cap(-\partial^{\infty}J_{2}(\widebar{\bm{x}}))=\{\bm{0}\}$ holds at $\widebar{\bm{x}}\in\mathrm{dom}(J_{1}+J_{2})$, then $J_{1}+J_{2}$ is subdifferentially regular at $\widebar{\bm{x}}$, and
    $\partial_{\rm L}(J_{1}+J_{2})(\widebar{\bm{x}})$ $=\partial_{\rm L}J_{1}(\widebar{\bm{x}})+\partial_{\rm L}J_{2}(\widebar{\bm{x}}).$ (97)
    Moreover, if $J_{2}$ is continuously differentiable, then $J_{1}+J_{2}$ is subdifferentially regular at $\widebar{\bm{x}}\in\mathrm{dom}(J_{1})$, and
    $\partial_{\rm L}(J_{1}+J_{2})(\widebar{\bm{x}})$ $=\partial_{\rm L}J_{1}(\widebar{\bm{x}})+\nabla J_{2}(\widebar{\bm{x}});$ (98)
    $\partial^{\infty}(J_{1}+J_{2})(\widebar{\bm{x}})$ $=\partial^{\infty}J_{1}(\widebar{\bm{x}}).$ (99)
Proof of Lemma 2.4.
[…]
(Step 2) Recall that $\iota_{C}$ is subdifferentially regular (see Example B.2 and Example 2.3), and lower semicontinuous due to the closedness of $C$. Then, by letting $J_{1}\coloneqq F$, $J_{2}\coloneqq\iota_{C}$, and $\mathcal{F}\coloneqq\bm{\varphi}$, Fact B.4 yields that (i) the subdifferential regularities of $F\circ\bm{\varphi}$ and $F+\iota_{C}$; (ii) the equalities (17) and (18) from (97) and (95), where $J_{1}$ is subdifferentially regular and $\partial^{\infty}J_{1}(\widebar{\bm{x}})=\{\bm{0}\}\ (\widebar{\bm{x}}\in\mathcal{X})$ from (Step 1). ∎
Appendix C Proof of Proposition 4.2
To prove Proposition 4.2, we use the following lemma for two times.
Lemma C.1 (Lipschitz continuity of gradient of composite function).
Let $\mathcal{H},\mathcal{K}$ be Euclidean spaces. Let $\mathcal{F}:\mathcal{K}\to\mathcal{H}$ and $J:\mathcal{H}\to\mathbb{R}$ be continuously differentiable, and let $(\emptyset\neq)E\subset\mathcal{K}$ be a subset of $\mathcal{K}$. Assume that
  1. (i)
    The operator norm of $(\mathrm{D}\mathcal{F}(\cdot))^{*}$ is bounded above by $\kappa_{\mathcal{F}}>0$ over $\mathrm{Conv}(E)$, and thus $\mathcal{F}$ is Lipschitz continuous with a Lipschitz constant $\kappa_{\mathcal{F}}$ over $\mathrm{Conv}(E)$ by [69, Thm. 9.2 and Thm. 9.7].
  2. (ii)
    $(\mathrm{D}\mathcal{F}(\cdot))^{*}$ is Lipschitz continuous with a Lipschitz constant $L_{\mathrm{D}\mathcal{F}}>0$ over $E$.
  3. (iii)
    The norm of $\nabla J$ is bounded above by $\kappa_{J}>0$ over $\mathcal{F}(E)$.
  4. (iv)
    $\nabla J$ is Lipschitz continuous with a Lipschitz constant $L_{\nabla J}>0$ over $\mathcal{F}(E)$.
Then, $\nabla(J\circ\mathcal{F})$ is Lipschitz continuous with a Lipschitz constant $\kappa_{\mathcal{F}}^{2}L_{\nabla J}+\kappa_{J}L_{\mathrm{D}\mathcal{F}}$ over $E$.
Proof.
For $\bm{x}_{1},\bm{x}_{2}\in E$, we have
$\left\lVert\nabla(J\circ\mathcal{F})(\bm{x}_{1})-\nabla(J\circ\mathcal{F})(\bm{x}_{2})\right\rVert=\left\lVert(\mathrm{D}\mathcal{F}(\bm{x}_{1}))^{*}[\nabla J(\mathcal{F}(\bm{x}_{1}))]-(\mathrm{D}\mathcal{F}(\bm{x}_{2}))^{*}[\nabla J(\mathcal{F}(\bm{x}_{2}))]\right\rVert$ (100)
$\leq\left\lVert(\mathrm{D}\mathcal{F}(\bm{x}_{1}))^{*}[\nabla J(\mathcal{F}(\bm{x}_{1}))-\nabla J(\mathcal{F}(\bm{x}_{2}))]\right\rVert+\left\lVert(\mathrm{D}\mathcal{F}(\bm{x}_{1})-\mathrm{D}\mathcal{F}(\bm{x}_{2}))^{*}[\nabla J(\mathcal{F}(\bm{x}_{2}))]\right\rVert$ (101)
[…]
Proof of Proposition 4.2.
From Fact 2.7 (b), recall that $\nabla{}^{\mu}g$ is Lipschitz continuous with the Lipschitz constant $L_{\nabla{}^{\mu}g}\coloneqq\max\left\{\mu^{-1},\frac{\eta}{1-\eta\mu}\right\}=\mu^{-1}$ by $\mu\in(0,2^{-1}\eta^{-1}]$.
(a) Since $\nabla{}^{\mu}g$ is Lipschitz continuous with the Lipschitz constant $L_{\nabla{}^{\mu}g}$ and $\left\lVert\nabla{}^{\mu}g(\bm{z})\right\rVert\leq L_{g}\ (\bm{z}\in\mathcal{Z})$ by (22), Lemma C.1 with setting $(\mathcal{H},\mathcal{K},\mathcal{F},J,E)\coloneqq(\mathcal{Z},\mathcal{X},\mathfrak{S},{}^{\mu}g,C)$ ensures the Lipschitz continuity of $\nabla({}^{\mu}g\circ\mathfrak{S})$ with a Lipschitz constant $L_{\nabla({}^{\mu}g\circ\mathfrak{S})}\coloneqq\kappa_{\mathfrak{S}}^{2}L_{\nabla{}^{\mu}g}+L_{g}L_{\mathrm{D}\mathfrak{S}}=\kappa_{\mathfrak{S}}^{2}\mu^{-1}+L_{g}L_{\mathrm{D}\mathfrak{S}}$. Thus, $\nabla(h+{}^{\mu}g\circ\mathfrak{S})$ is Lipschitz continuous over $C$ with a Lipschitz constant $L_{\nabla(h+{}^{\mu}g\circ\mathfrak{S})}=L_{\nabla h}+L_{g}L_{\mathrm{D}\mathfrak{S}}+\kappa_{\mathfrak{S}}^{2}\mu^{-1}$, where $L_{\nabla h}$ is the Lipschitz constant of $\nabla h$ by the setting of Problem 1.4.
(b) To apply Lemma C.1, we derive an upper bound $\kappa_{h+{}^{\mu}g\circ\mathfrak{S}}\coloneqq\kappa_{h}+\kappa_{\mathfrak{S}}L_{g}>0$ of $\sup_{\bm{x}\in C}\left\lVert\nabla(h+{}^{\mu}g\circ\mathfrak{S})(\bm{x})\right\rVert$ as:
$\sup_{\bm{x}\in C}\left\lVert\nabla(h+{}^{\mu}g\circ\mathfrak{S})(\bm{x})\right\rVert\leq\sup_{\bm{x}\in C}\left\lVert\nabla h(\bm{x})\right\rVert+\sup_{\bm{x}\in C}\left\lVert\nabla({}^{\mu}g\circ\mathfrak{S})(\bm{x})\right\rVert\leq\kappa_{h}+\kappa_{\mathfrak{S}}L_{g},$ (105)
[…]
Appendix D Proof of Theorem 4.4
To prove Theorem 4.4, we need the following lemma.
Lemma D.1 (Subdifferential limit).
Let $\mathcal{Y}$ and $\mathcal{Z}$ be Euclidean spaces. Let $J:\mathcal{Z}\to\mathbb{R}$ be locally Lipschitz continuous at $\widebar{\bm{z}}\in\mathcal{Z}$, and $A:\mathcal{Z}\to\mathcal{Y}$ a linear operator. Suppose that (i) an arbitrarily given $(\widebar{\bm{z}}_{n})_{n=1}^{\infty}\subset\mathcal{Z}$ satisfies $\lim_{n\to\infty}\widebar{\bm{z}}_{n}=\widebar{\bm{z}}$; (ii) an arbitrarily given $(A_{n})_{n=1}^{\infty}$ with linear operators $A_{n}:\mathcal{Z}\to\mathcal{Y}\ (n\in\mathbb{N})$ satisfies $\lim_{n\to\infty}A_{n}=A$. Then,
$\mathop{\mathrm{Lim~sup}}_{n\to\infty}A_{n}(\partial_{\rm L}J(\widebar{\bm{z}}_{n}))\subset A\left(\mathop{\mathrm{Lim~sup}}_{n\to\infty}\partial_{\rm L}J(\widebar{\bm{z}}_{n})\right)\subset A(\partial_{\rm L}J(\widebar{\bm{z}})).$ (106)
Proof.
The last inclusion in (106) can be checked by
[…]
where the last equality follows by the outer semicontinuity of the limiting subdifferential [69, Prop. 8.7] with the continuity of $J$ at $\widebar{\bm{z}}$.
To show the first inclusion in (106), let $\bm{v}\in\mathop{\mathrm{Lim~sup}}_{n\to\infty}A_{n}(\partial_{\rm L}J(\widebar{\bm{z}}_{n}))$, i.e., there exists a sequence $(\bm{v}_{n})_{n=1}^{\infty}\subset\mathcal{Y}$ with $\bm{v}_{n}\in A_{n}(\partial_{\rm L}J(\widebar{\bm{z}}_{n}))$ such that its subsequence $(\bm{v}_{m(l)})_{l=1}^{\infty}$ converges to $\bm{v}$, where $m:\mathbb{N}\to\mathbb{N}$ is monotonically increasing. Then, there exists a sequence $(\bm{u}_{n})_{n=1}^{\infty}\subset\mathcal{Z}$ with $\bm{u}_{n}\in\partial_{\rm L}J(\widebar{\bm{z}}_{n})$ satisfying $\bm{v}_{n}=A_{n}(\bm{u}_{n})$. From Fact B.3, the local Lipschitz continuity of $J$ at $\widebar{\bm{z}}$ ensures the boundedness of $\bigcup_{n=n_{0}}^{\infty}\partial_{\rm L}J(\widebar{\bm{z}}_{n})$ with a sufficiently large $n_{0}\in\mathbb{N}$. This implies the boundedness of $(\bm{u}_{n})_{n=1}^{\infty}$, thereby its subsequence $(\bm{u}_{m(l)})_{l=1}^{\infty}$ converges to some $\bm{u}\in\mathcal{Z}$ (by passing to further subsequence if necessary), i.e., $\bm{u}\in\mathop{\mathrm{Lim~sup}}_{n\to\infty}\partial_{\rm L}J(\widebar{\bm{z}}_{n})$. We get $\bm{v}=\lim_{l\to\infty}A\bm{u}_{m(l)}$ by
$\left\lVert\bm{v}-A\bm{u}_{m(l)}\right\rVert$ $\leq\left\lVert\bm{v}-\bm{v}_{m(l)}\right\rVert+\left\lVert\bm{v}_{m(l)}-A_{m(l)}\bm{u}_{m(l)}\right\rVert+\left\lVert A_{m(l)}\bm{u}_{m(l)}-A\bm{u}_{m(l)}\right\rVert$ (108)
$=\left\lVert\bm{v}-\bm{v}_{m(l)}\right\rVert+\left\lVert A_{m(l)}\bm{u}_{m(l)}-A\bm{u}_{m(l)}\right\rVert\quad(\because\bm{v}_{m(l)}=A_{m(l)}\bm{u}_{m(l)})$ (109)
$\leq\left\lVert\bm{v}-\bm{v}_{m(l)}\right\rVert+\left\lVert A_{m(l)}-A\right\rVert_{\rm op}\left\lVert\bm{u}_{m(l)}\right\rVert\overset{l\to\infty}{\to}0$ (110)
$\hskip 100.00015pt(\because\bm{v}_{m(l)}\overset{l\to\infty}{\to}\bm{v},\ A_{m(l)}\overset{l\to\infty}{\to}A,\ \bm{u}_{m(l)}\overset{l\to\infty}{\to}\bm{u}).$ (111)
[…]
Proof of Theorem 4.4.
(a) Let $\bm{v}\in\mathop{\mathrm{Lim~sup}}\limits_{\mathbb{R}_{++}\ni\mu\searrow 0,\ \bm{y}\to\widebar{\bm{y}}}\nabla({}^{\mu}J\circ\mathcal{F})(\bm{y})$. Then, there exist (i) $(\bm{y}_{n})_{n=1}^{\infty}\subset\mathcal{Y}$ with $\lim_{n\to\infty}\bm{y}_{n}=\widebar{\bm{y}}$; (ii) $(\mu_{n})_{n=1}^{\infty}\subset(0,\eta^{-1})$ with $\mu_{n}\searrow 0$; and (iii) $\bm{v}_{n}\coloneqq\nabla({}^{\mu_{n}}J\circ\mathcal{F})(\bm{y}_{n})=(\mathrm{D}\mathcal{F}(\bm{y}_{n}))^{*}[\nabla{}^{\mu_{n}}J(\mathcal{F}(\bm{y}_{n}))]$ with $\lim_{n\to\infty}\bm{v}_{n}=\bm{v}$. From Fact 2.7 (c), we have $\nabla{}^{\mu_{n}}J(\mathcal{F}(\bm{y}_{n}))\in\partial_{\rm L}J(\mathrm{prox}_{\mu_{n}J}(\mathcal{F}(\bm{y}_{n})))\ (n\in\mathbb{N})$. This implies $\bm{v}_{n}\in(\mathrm{D}\mathcal{F}(\bm{y}_{n}))^{*}[\partial_{\rm L}J(\mathrm{prox}_{\mu_{n}J}(\mathcal{F}(\bm{y}_{n})))]\ (n\in\mathbb{N})$, and thus
$\bm{v}\in\mathop{\mathrm{Lim~sup}}_{n\to\infty}(\mathrm{D}\mathcal{F}(\bm{y}_{n}))^{*}[\partial_{\rm L}J(\mathrm{prox}_{\mu_{n}J}(\mathcal{F}(\bm{y}_{n})))].$ (112)
For $n\in\mathbb{N}$, let $\widebar{\bm{z}}_{n}\coloneqq\mathrm{prox}_{\mu_{n}J}(\mathcal{F}(\bm{y}_{n}))$ and $A_{n}\coloneqq(\mathrm{D}\mathcal{F}(\bm{y}_{n}))^{*}$. Then, we have $\lim_{n\to\infty}A_{n}=(\mathrm{D}\mathcal{F}(\widebar{\bm{y}}))^{*}\eqqcolon A$ from the continuous differentiability of $\mathcal{F}$. Recall that $J$ is locally Lipschitz continuous [^22] Since every finite convex function is locally Lipschitz continuous [69, Exm. 9.14], $J+\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2}$ and $\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2}$ are locally Lipschitz continuous due to the weak convexity of $J$. Then, $J=(J+\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2})+(-\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2})$ is locally Lipschitz continuous because the sum of locally Lipschitz continuous functions is locally Lipschitz continuous [69, Exe. 9.8]. , and thus $\partial_{\rm L}J(\bm{z})\neq\emptyset$ for all $\bm{z}\in\mathcal{Z}$, by Fact B.3, i.e., $\mathrm{dom}(\partial_{\rm L}J)=\mathcal{Z}$. From $\mathrm{dom}(\partial_{\rm L}J)=\mathcal{Z}$, $\lim_{n\to\infty}\mathcal{F}(\bm{y}_{n})=\mathcal{F}(\widebar{\bm{y}})$, and $\mu_{n}\searrow 0$, Fact 2.7 (a) implies $\lim_{n\to\infty}\widebar{\bm{z}}_{n}=\lim_{n\to\infty}\mathrm{prox}_{\mu_{n}J}(\mathcal{F}(\bm{y}_{n}))=P_{\overline{\mathrm{dom}(\partial_{\rm L}J)}}(\mathcal{F}(\widebar{\bm{y}}))=\mathcal{F}(\widebar{\bm{y}})\eqqcolon\widebar{\bm{z}}$. From the local Lipschitz continuity of $J$, Lemma D.1 with $\lim_{n\to\infty}\widebar{\bm{z}}_{n}=\widebar{\bm{z}}$ and $\lim_{n\to\infty}A_{n}=A$ yields
[…]
from which $\bm{v}\in\partial_{\rm L}(J\circ\mathcal{F})(\widebar{\bm{y}})$ holds by (112).
(b) Theorem 4.4 (a) with $J\coloneqq g$ and $\mathcal{F}\coloneqq\mathfrak{S}\circ\bm{\varphi}$ yields $\partial_{\rm L}(g\circ\mathfrak{S}\circ\bm{\varphi})(\widebar{\bm{y}})\supset\mathop{\mathrm{Lim~sup}}\limits_{\mathbb{R}_{++}\ni\mu\searrow 0,\ \mathcal{Y}\ni\bm{y}\to\widebar{\bm{y}}}\nabla({}^{\mu}g\circ\mathfrak{S}\circ\bm{\varphi})(\bm{y})$. The inclusion in (37) is obtained by
[…]
$\supset\nabla(h\circ\bm{\varphi})(\widebar{\bm{y}})+\mathop{\mathrm{Lim~sup}}_{\mathbb{R}_{++}\ni\mu\searrow 0,\ \mathcal{Y}\ni\bm{y}\to\widebar{\bm{y}}}\nabla({}^{\mu}g\circ\mathfrak{S}\circ\bm{\varphi})(\bm{y})=\mathop{\mathrm{Lim~sup}}_{\mathbb{R}_{++}\ni\mu\searrow 0,\ \mathcal{Y}\ni\bm{y}\to\widebar{\bm{y}}}\nabla((h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi})(\bm{y}),\hskip-60.00009pt$ (116)
where the last equality follows by the continuity [^23] Indeed, the continuity of $\nabla(h\circ\bm{\varphi})$ yields the following relation: $\bm{v}\in\nabla(h\circ\bm{\varphi})(\widebar{\bm{y}})+\mathop{\mathrm{Lim~sup}}_{\mathbb{R}_{++}\ni\mu\searrow 0,\ \mathcal{Y}\ni\bm{y}\to\widebar{\bm{y}}}\nabla({}^{\mu}g\circ\mathfrak{S}\circ\bm{\varphi})(\bm{y})$ (117) $\Leftrightarrow\exists(\bm{y}_{n})_{n=1}^{\infty}\subset\mathcal{Y},\ \exists(\mu_{n})_{n=1}^{\infty}\subset(0,\eta^{-1})\ \mathrm{such\ that}\ \widebar{\bm{y}}=\lim_{n\to\infty}\bm{y}_{n},\ \mu_{n}\searrow 0,$ (118) $\hskip 20.00003pt\mathrm{and}\ \bm{v}=\nabla(h\circ\bm{\varphi})(\widebar{\bm{y}})+\lim_{n\to\infty}\nabla({}^{\mu_{n}}g\circ\mathfrak{S}\circ\bm{\varphi})(\bm{y}_{n})=\lim_{n\to\infty}\left(\nabla(h\circ\bm{\varphi})(\bm{y}_{n})+\nabla({}^{\mu_{n}}g\circ\mathfrak{S}\circ\bm{\varphi})(\bm{y}_{n})\right)$ (119) $\Leftrightarrow\bm{v}\in\mathop{\mathrm{Lim~sup}}_{\mathbb{R}_{++}\ni\mu\searrow 0,\ \mathcal{Y}\ni\bm{y}\to\widebar{\bm{y}}}(\nabla(h\circ\bm{\varphi})(\bm{y})+\nabla({}^{\mu}g\circ\mathfrak{S}\circ\bm{\varphi})(\bm{y}))=\mathop{\mathrm{Lim~sup}}_{\mathbb{R}_{++}\ni\mu\searrow 0,\ \mathcal{Y}\ni\bm{y}\to\widebar{\bm{y}}}\nabla((h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi})(\bm{y}).$ (120) of $\nabla(h\circ\bm{\varphi})$. Moreover, we have
[…]
Then, the inequality in (38) is obtained by
[…]
$=\liminf_{n\to\infty}d(\bm{0},\nabla((h+{}^{\mu_{n}}g\circ\mathfrak{S})\circ\bm{\varphi})(\bm{y}_{n}))=\liminf_{n\to\infty}\left\lVert\nabla((h+{}^{\mu_{n}}g\circ\mathfrak{S})\circ\bm{\varphi})(\bm{y}_{n})\right\rVert,$ (123)
where the first equality follows by the fact $\liminf_{n\to\infty}d(\bm{v},E_{n})=d\left(\bm{v},\mathop{\mathrm{Lim~sup}}_{n\to\infty}E_{n}\right)$ for $\bm{v}\in\mathcal{Y}$ and a given sequence $(E_{n})_{n=1}^{\infty}$ of subsets $E_{n}\subset\mathcal{Y}$ (see [69, Exe. 4.8]). ∎
Appendix E Proof of Corollary 5.1
To prove Corollary 5.1, we need the following fact and lemma.
Fact E.1 ([48, Lemma A.4 and proof in Proposition 2.9]).
  1. (a)
    For $\bm{A}\in\mathbb{R}^{l\times m}$ and $\bm{B}\in\mathbb{R}^{m\times n}$, $\|\bm{A}\bm{B}\|_{F}\leq\|\bm{A}\|_{2}\|\bm{B}\|_{F}$ and $\|\bm{A}\bm{B}\|_{F}\leq\|\bm{A}\|_{F}\|\bm{B}\|_{2}$ hold, where $\left\lVert\cdot\right\rVert_{F}$ and $\left\lVert\cdot\right\rVert_{2}$ denote the Frobenius norm and the spectral norm.
  2. (b)
    For $\bm{V}\in Q_{N,p}$, we have $\|(\bm{I}+\bm{V})^{-1}\|_{2}\leq 1$.
  3. (c)
    For $\bm{V}_{1},\bm{V}_{2}\in Q_{N,p}$, $\left\|(\bm{I}+\bm{V}_{1})^{-1}-(\bm{I}+\bm{V}_{2})^{-1}\right\|_{F}\leq\|\bm{V}_{1}-\bm{V}_{2}\|_{F}$.
  4. (d)
    For $\bm{V},\bm{D}\in Q_{N,p}$, $\mathrm{D}\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V})[\bm{D}]=-2\bm{S}(\bm{I}+\bm{V})^{-1}\bm{D}(\bm{I}+\bm{V})^{-1}\bm{I}_{N\times p}$
Lemma E.2 (Boundedness and Lipschitz continuity of $\mathrm{D}\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V})$).
For $\bm{\varphi}_{\bm{S}}^{\rm Cay}$ in (77) with $\bm{S}\in{\rm O}(N)$, we have
$(\bm{V}\in Q_{N,p})\quad$ $\left\lVert\mathrm{D}\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V})\right\rVert_{\rm op}\leq 2;$ (124)
$(\bm{V}_{1},\bm{V}_{2}\in Q_{N,p})\quad$ $\left\lVert\mathrm{D}\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}_{1})-\mathrm{D}\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}_{2})\right\rVert_{\rm op}\leq 4\left\lVert\bm{V}_{1}-\bm{V}_{2}\right\rVert_{F}.$ (125)
Proof.
(Proof of (124)) Let $\bm{V},\bm{D}\in Q_{N,p}$ such that $\left\lVert\bm{D}\right\rVert_{F}\leq 1$. By Fact E.1 (d), we have
$\left\lVert\mathrm{D}\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V})[\bm{D}]\right\rVert_{F}=2\left\lVert\bm{S}(\bm{I}+\bm{V})^{-1}\bm{D}(\bm{I}+\bm{V})^{-1}\bm{I}_{N\times p}\right\rVert_{F}$ (126)
[…]
implying thus (124), where we used $\left\lVert\bm{S}\right\rVert_{2}=\left\lVert\bm{I}_{N\times p}\right\rVert_{2}=1$.
(Proof of (125)) Let $\bm{V}_{1},\bm{V}_{2},\bm{D}\in Q_{N,p}$ such that $\left\lVert\bm{D}\right\rVert_{F}\leq 1$. By Fact E.1 (d), we have
$\left\lVert\mathrm{D}\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}_{1})[\bm{D}]-\mathrm{D}\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}_{2})[\bm{D}]\right\rVert_{F}$ (128)
$=2\left\lVert\bm{S}\left((\bm{I}+\bm{V}_{1})^{-1}\bm{D}(\bm{I}+\bm{V}_{1})^{-1}-(\bm{I}+\bm{V}_{2})^{-1}\bm{D}(\bm{I}+\bm{V}_{2})^{-1}\right)\bm{I}_{N\times p}\right\rVert_{F}$ (129)
[…]
$\leq 2\left\lVert(\bm{I}+\bm{V}_{1})^{-1}\bm{D}\left((\bm{I}+\bm{V}_{1})^{-1}-(\bm{I}+\bm{V}_{2})^{-1}\right)\right\rVert_{F}$ (131)
$\quad+2\left\lVert\left((\bm{I}+\bm{V}_{1})^{-1}-(\bm{I}+\bm{V}_{2})^{-1}\right)\bm{D}(\bm{I}+\bm{V}_{2})^{-1}\right\rVert_{F}.\quad(\because\left\lVert\bm{S}\right\rVert_{2}=\left\lVert\bm{I}_{N\times p}\right\rVert_{2}=1)$ (132)
The first term in (132) can be evaluated further as
$\left\lVert(\bm{I}+\bm{V}_{1})^{-1}\bm{D}\left((\bm{I}+\bm{V}_{1})^{-1}-(\bm{I}+\bm{V}_{2})^{-1}\right)\right\rVert_{F}$ (133)
[…]
Since the second term in (132) can be evaluated similarly to (136), we have $\left\lVert\mathrm{D}\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}_{1})[\bm{D}]-\mathrm{D}\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}_{2})[\bm{D}]\right\rVert_{F}\leq 4\left\lVert\bm{V}_{1}-\bm{V}_{2}\right\rVert_{F}$, implying thus (125). ∎
Proof of Corollary 5.1.
(a) By noting that $\bm{\varphi}_{\bm{S}}^{\rm Cay}:Q_{N,p}\to{\rm St}(p,N)\setminus E_{N,p}(\bm{S})$ is a diffeomorphism [48, Prop. 2.2], Corollary 3.4 with $C\coloneqq{\rm St}(p,N)$ and $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$ completes the proof [^24] Although $\bm{\varphi}_{\bm{S}}^{\rm Cay}$ is not surjective onto ${\rm St}(p,N)$, the same statement in Corollary 3.4 follows (see Remark 3.5). (see also the footnote7 in Corollary 3.4).
(b) To invoke Proposition 4.2, we check below the conditions (A1)-(A5) in Proposition 4.2 with $C\coloneqq{\rm St}(p,N)$ and $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$. Since ${\rm St}(p,N)$ is compact and $\mathfrak{S}$ is twice continuously differentiable, (A1)-(A3) are automatically satisfied (see Remark 4.3 (a)). For (A4)-(A5), see Lemma E.2. Therefore, by Proposition 4.2 (b), Assumption 4.7 (a) is satisfied with $C\coloneqq{\rm St}(p,N)$ and $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$. ∎